### v2

# Quantum Sensing Using Cold Atoms and Molecules for Nuclear and Particle Physics

Abstract
Technologies for manipulating single atoms and molecules have advanced drastically in the past decades. Due to their excellent controllability of internal states, atoms and molecules serves as one of the ideal platforms as quantum systems. One major research direction in atomic systems is the precise determination of energy differences between two energy levels in atoms. This field, known as precision measurement, is symbolized by the remarkable fractional uncertainty of or lower achieved in the state-of-the-art atomic clocks. Two-level systems in atoms are sensitive to various external fields and can therefore be function as quantum sensors. The effect of these fields manifests as energy shifts in the two-level system. Traditionally, such shifts are induced by electric or magnetic fields, as recognized even before the advent of the precision spectroscopy with lasers. However, with high-precision measurements, energy shifts can also be caused by hypothetical fields weakly coupled to ordinary matter, or by small effects mediated by massive particles, which are conventionally dealt with in the field of nuclear and particle physics. In most cases, the atomic systems as quantum sensors have not been sensitive enough to detect such effects. Instead, experiments searching for these interactions have placed constraints on coupling constants, except in a few cases where effects are predicted by the Standard Model of particle physics. Nonetheless, measurements and searches for these effects in atomic systems have led to the emergence of a new field of physics. In some cases, they open new parameter spaces to explore in conventionally investigated topics, e.g., dark matter, fifth force, CP-violation, and other physics beyond the Standard Model. In other cases, these measurements provide alternative experimental approaches to established topics, e.g., variations of fundamental constants searched astronomically, parity violation and nuclear structure studied in high-energy scattering experiments, and supersymmetric particles investigated with high-energy colliders. The use of cold atoms and molecules as quantum sensors for phenomena originating from nuclear and particle physics evolved significantly in the past decades. This paper highlights the recent developments in the field.
I Introduction to quantum sensing for nuclear and particle physics
Quantum sensing is defined in the following three ways, as carefully described in Ref. [1].
(i) a measurement of a physical quantity using quantized energy levels as a detector.
(ii) a measurement of a physical quantity based on the superposition of quantum states.
(iii) a measurement of a physical quantity employing quantum entanglement to enhance its sensitivity.
There are many physical realizations of quantized energy levels defined in (i). Examples are superconducting qubits [2, 3], mechanical harmonic oscillators including micro- or nanoparticles and suspended mirrors [4, 5], spin systems in solid state materials such as nitrogen-vacancy centers [6, 7, 8], cold atoms [9] and molecules [10] that are often trapped, and trapped ions [11, 12]. Among them, this paper focuses on cold atoms, molecules, and ions. Laser cooling and trapping allows them to be kept in the temperature of mK or even lower without any cryogenic systems. Coherence can be maintained for seconds thanks to the excellent isolation from the environment. Rapidly developing trapping and manipulation techniques enable independent and arbitrary control of the quantum state of a single atom, ion, or molecule. This capability makes these systems suitable for quantum sensors.
When an atomic or molecular system is selected as a quantum system, it generally has many energy levels. For the quantum sensing purposes, specific energy levels are typically selected, and control of the states are performed by applying oscillating electromagnetic fields on or near resonance. In such cases, the system can be well approximated as a two-level system consisting of the ground state and the excited state . The relevant quantities are summarized in Fig. 1. The energy difference between the two energy levels are , where is the Planck constant. The excited state has a natural decay linewidth denoted by . The oscillating electromagnetic field has a frequency of detuned by from .
An external field can induce a frequency shift of from in the two-level system. can be converted to the magnitude of the external field, with a known relation between the field strength and . One example is the frequency shift caused by a magnetic field, which is known as the Zeeman effect. When a small external magnetic field is applied, assuming that the ground state is insensitive to the magnetic field, , where is the Bohr magneton, is Landé g-factor for the excited state, and is the magnetic quantum number for the total angular momentum . Frequency shifts can be some other fields. Another well-known effect is the Stark effect, which is induced by an external electric field.
The external field is not necessarily limited to electromagnetic or other well-established fields in conventional physics. Consider a hypothetical bosonic field that induces , whose constituent particle has a mass and thus can be a candidate for dark matter (such fields are discussed in Section VII). If a two-level system is sufficiently sensitive, the field can be detected using this system. Searching for such a hypothetical field effectively serves as a search for new particles. Even if no frequency shift is observed, the measurement sets an upper bound on the interaction strength between the field, meaning that the interaction is weaker than the sensitivity of the two-level system. This falls within the realm of physics beyond Standard Model (BSM), which is one of the major topic in the contemporary particle physics. Another example is the sensitivity of atomic spectroscopy to nuclear charge distribution. The finite size of the nucleus causes a deviation in the energy levels of the electronic states in an atom compared to the case of a point charge. As a result, slight differences in nuclear charge radii between isotopes can be detected as shifts in transition frequencies between isotopes for the same electronic transition, which is called the isotope shift. Here, a structure of fm, which corresponds to an energy scale of MeV, is probed by spectroscopy of electronic excitations at an energy scale of eV.
These two examples illustrate how the quantum sensors, particularly those with precision spectroscopy, are sensitive to the nuclear and particle physics. One approach is to try to detect hypothetical interactions with small couplings to the Standard Model (SM) particles through high-precision measurement, such as state-of-the-art atomic clocks. Examples covered in this paper include searches for time variation of fundamental constants, ultralight dark matter, and new force between an electron and a neutron. Although axions and axion-like particles fall within this type of interaction, they are not discussed in this paper, as their detection schemes do not significantly overlap with the field of cold atoms: instead, vapor cells are commonly used. Some other review papers[13, 14, 15, 16] deal with them in details. Theoretical proposals for these interactions have been tested experimentally, and so far, the experiments give null results, setting constraints on such hypothetical interactions. Another approach involves phenomena conventionally studied through high-energy collisions. In this case, some precision measurement experiments detect high-energy phenomena in atomic systems. The topics covered in this paper are atomic parity violation (APV), electric dipole moment (EDM) searches, and nuclear structure determination via isotope shifts. APV addresses the physics of GeV, corresponding to the mass of the W and Z bosons, whereas EDM searches aim to probe physics of TeV, searching for supersymmetric particles. The effects of massive particles can manifest through loop-diagrams that are higher-order processes than tree-level Feynman diagrams, unless the heavy particle itself mediates the interaction. Note that throughout this paper, the term "higher-order perturbation" specifically refers to such fundamental interactions in particle physics, while the higher-order perturbations of electromagnetic fields are referred to as "higher-order effects."
The effects of nuclear and particle physics phenomena are typically extremely small, compared to the conventional shifts caused by external electromagnetic fields. Distinguishing these tiny shifts from those induced by conventional electromagnetic fields is crucial. In this paper, the term "signal" refers to the quantity aimed to be detected: namely, the effects originating from nuclear and particle physics, e.g., the effect of the field. All other effects that behave similarly to the targeted signal are regarded as "backgrounds". These backgrounds include shifts caused by conventional external fields such as electric and magnetic fields. Even among effects induced by nuclear and particle physics, one can be considered as the signal, while others act as the backgrounds. For example, in searches for BSM physics, other effects from the SM are backgrounds. To discriminate signal from backgrounds properly, multiple methods are employed. In addition, a proper theoretical framework is essential to extract nuclear and particle physics effects from slight energy shifts. The following discussions focus on how to select an appropriate two-level system to extract a desired nuclear and particle physics phenomenon, and how to cancel all backgrounds that can potentially mimic the signal.
Definition (ii) is also broad. In a two-level system, a superposition state is frequently used to detect a phase drift between the two-level system and the driving oscillating electromagnetic field (see Section II). Other methods of utilizing superposition to enhance the sensitivity of quantum sensors are discussed, e.g., in Section IX. Among the three definitions of quantum sensing, Definition (iii) is the most stringent. Entangled states, such as squeezed states and Schrödinger’s cat states, are typically difficult to generate and maintain. As a result, examples of measurements with entangled states are limited. This paper discusses quantum sensing with entanglement when relevant reports are available. Experimental results are covered in Section III, while theoretical proposals are mentioned in Section X.
This paper first briefly summarizes how atomic and molecular systems are prepared as quantum sensors. This section is intended for readers outside of the field of atomic, molecular, and optical physics (AMO) to offer a glimpse of how AMO experiments operate. As a realization of the two-level system, optical atomic clocks are first discussed. This is partly because precision spectroscopy of atomic two-level systems forms the foundation of precision measurements and because optical atomic clocks are the state-of-the-art in this field. Additionally, atomic clocks serve as platforms for exploring nuclear and particle physics through precision measurements. The systematic shifts and uncertainties discussed here provide typical factors to consider in other types of precision spectroscopy. This part includes details of AMO physics. The discussion is followed by searches for variation of fundamental constants. This shows the sensitivity of atomic clocks as quantum sensors, also introducing the importance of two clock architectures: highly charged ion (HCI) clocks and nuclear clocks. These two systems are not only promising candidate for next-generation optical clocks [17, 18], but also serve as quantum sensors for variations of fundamental constants and other nuclear and particle physics phenomena. A related topic is searches for ultralight dark matter, where periodic modulation of fundamental constants are interpreted as potential signals of ultralight dark matter. After this, searches for a new force between an electron and a neutron through isotope shifts are discussed, which also function as precise measurements of the nuclear structure. Then searches for atomic parity violation is mentioned, which involves weak-force interactions between the nucleus and an electron. Finally, the paper discusses searches for EDMs in electrons and nuclei. The section for the variation of fundamental constants, ultralight dark matter, fifth force search with isotope shifts, APV, and EDM searches are structured with separate discussions on the theoretical frameworks and experimental results. These subsections provide overviews of nuclear and particle physics theories and AMO experiments, respectively.
How are these topics related to quantum sensors? As introduced, the basis of quantum sensors discussed in this paper is the two-level system. In an atomic clock, a two-level system is selected as the source of stability. Various systematic shifts, mainly induced by external electromagnetic fields, are suppressed as a clock. This opens the possibility of detecting small signal originating from nuclear and particle physics phenomena. Examples of such signals are the variation of fundamental constants, ultralight dark matter, a new force between an electron and a neutron, and nuclear charge distribution. In the searches for APV, a two-level system with the same parity is chosen as the quantum sensor. The magnitude of parity-violating transition rate between the two-level system quantifies APV as the signal. In EDM searches, a spin 1/2 state serves as the quantum sensor for the EDM, and the deviation of the precession from purely spin system due to the EDM is the signal.
II Typical procedure to utilize atoms and molecules for quantum sensors
In this section, typical experimental procedures for preparing atoms as quantum sensors are briefly summarized. In some cases, for example molecular beam experiments or ion trap experiments, certain cooling and trapping procedures are different from those described below for atoms.
Trapping atoms begins with generating an atomic beam, a process that depends on the vapor pressure of the atoms. For low-vapor-pressure atoms such as Yb or Sr, an atomic oven heated to a few hundred degrees Celsius is used. For alkali atoms, which have relatively high vapor pressure, atomic dispensers are commonly used. The resulting atomic beam has a center velocity of a few hundred m/s, requiring deceleration. A Zeeman slower [19] is sometimes employed for this purpose, where a circularly polarized laser beam slightly off-resonant from a broad atomic transition propagates to the opposite direction of the atomic beam and also a spatially varying magnetic field is applied. This combination slows down atoms within a specific velocity range to m/s.
Sufficiently slow atoms can be trapped by a magneto-optical trap (MOT) [20]. A MOT is formed by a combination of circularly polarized light and a quadrupole magnetic field. In the standard configuration, counter-propagating laser beams with polarization, red-detuned from a broad-linewidth transition, provide a dissipation term. The quadrupole magnetic field induces a position-dependent scattering rate, generating an approximately harmonic potential. As a result, the atoms are simultaneously slowed, cooled, and confined within the trap. For a transition with MHz, the Doppler-limited temperature is mK, where is the Boltzmann constant. Using a narrower-linewidth transition, e.g., kHz, atoms can be cooled down to K. In addition, for atoms with Zeeman sublevels in the ground state, polarization gradient cooling [21] can further reduce the temperature below the Doppler limit.
After being trapped in a MOT, atoms are typically transferred to an optical lattice. An optical lattice is formed by pairs of counter-propagating laser beams that create a standing wave. Due to the ac Stark shift, atoms are attracted to the most (least) intense part of the light field when the trapping laser is red- (blue-) detuned from broad-linewidth transitions [22]. Depending on the number and configuration of the pairs of the counterpropagating laser beams, one-, two-, or three-dimensional lattice can be formed.
For precision measurements, the key role of an optical lattice is to put atoms in a recoil-free environment. Thermal atoms that are not trapped change their motional states upon absorbing or emitting a photon. The energy changes due to this recoil is called recoil energy , where is the mass of the atom, is the speed of light and is the recoil frequency. When the trapping frequency of atoms along the direction of the probe light resonant with the two-level system is , and , the atoms remain in their motional states after absorbing or emitting a photon, making the spectroscopy recoil-free. Intuitively, the momentum of the photon is absorbed by the entire trap, similar to the Mössbauer effect. is called Lamb-Dicke parameter, and when the condition is satisfied, atoms are in Lamb-Dicke regime, where the atoms do not change their motional state when they absorb or emit photons.
In the optical lattice, atoms can be further cooled using various techniques. Demonstrated methods include evaporative cooling [23], sideband-resolved Raman cooling [24, 25], cooling with a narrow-linewidth transition [26], and Sisyphus cooling [27]. The final temperature is very low, sometimes 100 nK or lower. Colder atoms with lower average energy are essential for achieving trapping in a shallow optical lattice typically used in the state-of-the-art optical lattice clocks.
The final stage of state preparation involves initializing all atoms into a single internal state. The ground state of an atom generally consists of multiple sublevels, such as Zeeman sublevels and hyperfine structures. In general, atoms are distributed over these states, but to utilize them as quantum sensors, they need to be initialized in a single state. Optical pumping [28, 29] is the simplest method for this purpose. The exact procedure depends on the system. For example, if the ground state has several Zeeman sublevels with different but no hyperfine structure, irradiating the atoms with -polarized light resonant with a broad-linewidth transition can transfer all atoms into the highest state. More sophisticated techniques such as stimulated Raman adiabatic passage[30, 31], rapid adiabatic passage [32, 33], and coherent population trapping [34] are also widely used for state preparation.
After state preparation is completed, atoms undergo the measurement sequence. While the measurement sequence depends on experiments, there are two basic procedures for a sequence. The sequence is performed on a two-level system as illustrated in Fig. 1. The laser intensity is characterized by the Rabi frequency . A more detailed formalism can be found in standard textbooks on atomic physics [35].
The first sequence is the Rabi sequence, where resonant light drives the two-level system. If the initial state is , the final state after a pulse of duration satisfies . When and , . Such a pulse to drive the system from to is called a pulse. If the pulse duration is half of the pulse, , and this pulse is referred to as a pulse. The relative phase between and in can be controlled by adjusting the phase offset of the driving light field. If is perturbed by external fields, the nominal pulse duration may not result in the desired final state. To mitigate this, composite pulses, which are extensively studied in the field of nuclear magnetic resonance (NMR) [36, 37], can be applied in laser spectroscopy.
The second basic sequence is the Ramsey sequence. In this sequence, atoms are initially prepared in . First, a pulse is applied, placing the atoms in a superposition state . Atoms are then left for a duration of , which is referred to as the interrogation time. During this period, if the laser frequency generally dependent on time is different from , the accumulated phase is recorded as a phase shift between the atomic system and the laser. is read out by another pulse and subsequent proper state detection of and . Note that when is approximately constant, , meaning that the Ramsey sequence can be used to measure . If and is shifted by due to an external field, serves as a sensor for . Both the Rabi and Ramsey sequences are used to measure the frequency offset between the laser system and the atomic system to lock the laser to the atomic resonance.
To function as a sensor, the final state needs to be detected. The most common detection method involves driving a broad-linewidth transition, e.g., MHz, from using a light field and measuring the response of the atoms. One method is fluorescence imaging, where scattered photons from the atoms are collected by a photodetector or a camera. Alternatively, absorption imaging can be used, in which a camera records the probe light transmitted through the atomic cloud. The presence of atoms causes attenuation of the probe beam due to light scattering in other directions, resulting in a shadow in the image. To obtain the fraction of atoms in over the total atom number, population in also needs to be detected. This is achieved by transferring atoms from back to and imaging them separately, after removing all atoms originally in . This process of transferring atoms from to through another excited state from which atoms decay to is known as repumping.
III Atomic clock
III.1 Introduction
III.1.1 Components of clocks
Modern stable clocks, including atomic clocks, consist of three main components: an oscillator, a counter, and a reference. The periodic motion of the oscillator is counted by the counter, which displays the time. For the stability purpose, the oscillator is locked to the reference. Nowadays, the most stable and accurate clocks are realized by optical atomic clocks, i.e., an atomic clock using an optical transition for clock operation. In an optical atomic clock, these three components are physically realized as a stable laser, a frequency comb, and an atomic transition in or near optical frequencies, respectively. The frequency of the stable laser is tuned to the resonant frequency of the atomic transition, so that the response of atoms as a result of interrogation by the laser can serve as a signal for feedback to stabilize the laser frequency. Since the atomic transition provides the ultimate source of stability, a deep understanding of the behavior of atoms is crucial for improving the accuracy of these clocks.
When a system is regarded as stable, measurements performed in the system at different times yields the same value. The variation of the measured value is characterized by uncertainty, which is categorized into two types: statistical uncertainty and systematic uncertainty. Statistical uncertainty reflects the spread of measured values when repeated measurements are performed under the same conditions. Systematic uncertainties accounts for uncertainties other than the statistical uncertainty. In atomic clocks, systematic uncertainties typically include the uncertainties on the estimated frequency shifts induced by external fields and other factors. More stable systems enable higher precision measurements, because of their smaller statistical uncertainty. Precision is defined by a small fractional uncertainty, i.e., the uncertainty divided by the measured value. Accuracy refers to how close the measurement value is to the true value.
For a stable reference, it is desirable to have a high sensitivity to small frequency shifts. This requirement can be achieved by using a transition with a large quality factor . Such transitions are forbidden transitions, where a forbidden transition is defined as a transition that cannot be driven by the electric dipole (E1) interaction. The absence of transition probability due to the strongest interaction between atoms and light results in a long-lived excited state, if the excited state is without alternative decay path to the ground state allowed by E1 transitions. In addition, for atomic clocks, it is crucial to select an atomic resonance with minimal sensitivity to external fields.
Historically, microwave transitions between states of two hyperfine levels in the ground state of alkali atoms, such as Cs and Rb, were used as atomic transitions. Over the decades, the accuracy of the Cs atomic clock has improved by six orders of magnitude, as shown in Fig. 2. With the advent of laser spectroscopy, optical atomic clocks have advanced much more rapidly than microwave atomic clocks. One advantage of optical atomic clocks is the transition frequency , which is five orders of magnitude larger than that for microwave atomic clocks and thus helps increase .
Optical clocks are sorted into two types: optical lattice clocks and ion clocks. An optical lattice clock interrogates thousands of atoms trapped in an optical lattice, whereas an ion clock conventionally operates with a single ion trapped in a Paul trap. Initially, the ion clocks were the leading optical clocks in terms of performance. The optical lattice clock was proposed in 2001 [92, 93], and first demonstrated in 2005 [67]. Around this time, the accuracy of optical atomic clocks started to surpass that of microwave clocks, and since then, their accuracies improved by three orders of magnitude. Currently, the smallest fractional uncertainties achieved by optical lattice clocks and ion clocks are comparable, with each type offering distinct advantages and disadvantages. The early development of optical clocks has been summarized well in previous review papers [94, 95].
III.1.2 Allan deviation and the stability of atomic clocks
The stability of atomic clocks is typically described by the Allan deviation [96]. It is a widely used measure of the stability of time-varying signals and is defined as follows:
where denotes the expectation operator, , and is the th data point taken at time intervals of . The Allan deviation follows a power-law behavior depending on the dominant noise source. It scales as , , , , and for the phase noise, flicker frequency noise, white frequency noise, random walk frequency noise, and frequency drift, respectively.
The stability of state-of-the-art atomic clocks described by the Allan deviation is limited by quantum mechanical randomness when a superposition state between and collapses into either or upon measurement. This effect is known as quantum projection noise (QPN) [97], and the corresponding stability limit is called the standard quantum limit (SQL). The SQL is given by
where is the length of a single measurement cycle, is the total measurement time, is the Wineland parameter for spin squeezing [98], and is the number of atoms in the system [99]. For a coherent spin state [100], , while when the spin state is squeezed in the phase direction during the Ramsey sequence. The optimal is the Ramsay coherence time , which for alkaline-earth-atom-like atoms (see Section III.2.2) can be calculated as [101]
where () is the atom loss rate from the trap for the excited (ground) state, is the Raman scattering rate of the trap laser for the excited state, is the rate at which black body radiation (BBR) induces the transition from the excited state to the state via the state, and is the branching ratio from the to the states.
III.2 Popular atomic species for atomic clocks
For ion clocks and optical lattice clocks, atoms with certain kinds of transitions are preferred. A narrow-linewidth transition is essential for clock operation. Also, a broad-linewidth transition is important for efficient laser cooling. Atoms with both types of transitions can be sorted into two main groups.
III.2.1 Alkali-atom-like species
Alkali-atom-like species have a single electron outside a closed shell and have an energy structure shown in Fig. 3(a). A key advantage of these species is the strong transition from the ground state to a excited state suitable for laser cooling. The alkali atoms themselves do not have a narrow-linewidth transition suitable for clock transition, where a clock transition is defined as a transition selected for clock operation. However, some singly-ionized ions of alkaline-earth-atom-like atoms have such narrow-linewidth transitions. In these ions, narrow-linewidth transition used as the clock transition is typically from the ground state to a excited state, corresponding to an electric quadrupole (E2) transition. Examples of ion clocks using this transition are Ca+ [102], Sr+ [103], Ba+ [104], Ra+ [105], and Hg+ [61]. The linewidth of the E2 transition is typically Hz, as listed in Table 1. An exception is Yb+, which also has an electric octupole (E3) transition to an state, in addition to the two E2 transitions to states. Because the E3 transition is a higher-order effect than an E2 transition, it has an exceptionally narrow linewidth of 3.20(16) nHz [106].
III.2.2 Alkaline-earth-atom-like species
Alkaline-earth-atom-like atoms have two valence electrons outside of a closed shell. In the ground state, all electronic spins pair up, resulting in zero electronic angular momentum. Compared to alkali atoms, these species offer a wider range of electronic states, as shown in Fig. 3(b), some of which are advantageous for atomic clocks. The ground state is a singlet state that can be excited to the singlet state via a broad-linewidth E1 transition. This transition is used for the first stage of laser cooling. Among the triplet states, the state has a total angular momentum of , allowing transitions from the ground state that satisfy the selection rule , except that the transition is between a singlet and a triplet state. Such intercombination transitions are weakly allowed due to spin-orbit (L-S) coupling, which results in a significantly narrower linewidth than the transition to the state. The transitions in some atoms, such as Sr and Yb have narrow-linewidth convenient for cooling atoms down to K or lower, leading to the use in the second stage of leaser cooling.
The transition from the ground state to the state is strictly forbidden because both states have . However, if the nucleus has a nonzero magnetic moment, it induces slight mixing between the state and the state, making the transition probability finite. The resulting linewidth is as narrow as mHz, as summarized in Table 1. Also, thanks to the absence of electronic angular momentum, the transition is insensitive to magnetic fields, which is advantageous for stability under ambient fields. Because of these features, the transition is widely used as the clock transition. Note that without a nuclear magnetic moment, a strong magnetic field must be applied to induce mixing at a level comparable to the case with a nuclear magnetic moment. Experimentally, Rabi frequencies of Hz are reported for Sr[107] and Yb[108] with a magnetic field of a few mT and a clock laser intensity of mW/cm2.
Atom | Transition | |||
27Al+ | 1121 | 7.73(53) mHz [109] | ||
40Ca+ | 411 | 0.186 Hz [110] | ||
88Sr+ | 445 | 0.4 Hz [111] | ||
171Yb+ | 688 | 2.58(27) Hz [112] | ||
171Yb+ | 642 | 3.20(16) nHz [106] | ||
199Hg+ | 1065 | 1.83(8) Hz [113] | ||
87Sr | 429 | 1.35(3) mHz [114] | 813 | |
88Sr | 429 | 813 | ||
171Yb | 518 | 7.03(6) mHz [115] | 759 | |
199Hg | 1129 | 0.1 Hz [95] | 363 |
The state also has a sufficiently narrow linewidth to serve as a clock transition, because the transition from the ground state to the state is the magnetic quadrupole transition. However, it has not been used as intensively as the state for clock operations. Instead, this state is used as a probe for many-body physics in quantum degenerate gases [116]. Yb has one more metastable state: the state. This state is theoretically predicted to be beneficial for various fundamental physics searches [117, 118]. Recently the direct transitions to this state from the ground state [119, 120] and the state [121] have been observed for the first time. Although the uncertainty in its absolute frequency measurement still needs improvement before it can be used as a clock transition, an initial analysis of its potential application in nuclear and particle physics has already been reported [122].
Many atomic species with alkaline-earth-atom-like structure are used in atomic clocks. Among neutral atoms, most alkaline-earth-atom-like atoms have been experimentally investigated: Mg [123], Ca [91], Sr [78], Cd [124], Yb [125], and Hg [126]. In the case of Ba the states lie at lower energies than the states, preventing the state from serving as a clock transition. So far, only a theoretical proposal is available for Be [127]. The states in Ca have exceptionally a narrow linewidth, and in the past, the state was investigated as a clock transition. Sr and Yb are the most common atomic species for optical lattice clocks. Recently, the narrow-linewidth transitions in these atoms have also been applied to atom interferometry [128] and quantum computation [129, 130]. Some ion clocks also have alkali-earth-atom-like electronic structure. Examples are Al+ [65], In+ [79], and Lu+ [131].
III.3 Optical lattice clock
An optical lattice clock is an atomic clock where thousands of neutral atoms are trapped in an optical lattice. The trapping laser is tuned to a special frequency called the magic frequency (often referred to as the magic wavelength when describing the corresponding wavelength) [92, 93]. At this frequency, the ac Stark shift caused by the trapping laser is cancelled out between the ground and excited states of the clock transition. Most atoms suitable for optical lattice clocks have alkaline-earth-atom-like structure, as they have a narrow-linewidth transition in neutral atoms.
The stability of the state-of-the-art optical lattice clock reached in 2013 [132]. In this report, the stability at 1 s was , and by integrating over s, the stability reached the low range. This high stability is achieved by a combination of low QPN of with and ms. to further improve the SQL, increasing is crucial, which requires improving the stability of the clock laser. Efforts to develop more stable clock lasers are ongoing using cryogenic cavities [133, 134], achieving a fractional stability of from 0.8 s to 80 s, or 16 mHz linewidth for 1542 nm laser.
A recently developed technique for quickly estimating systematic shifts is the differential measurement of multiple atomic ensembles with a common clock laser. The advantage of this method is that the differential response of atoms is immune to laser fluctuation and noise, including the Dick effect [135, 136], enabling precise comparisons even with a less stable laser. This method can be implemented various ways, such as two atomic ensembles in separate vacuum chambers[137], multiple atomic ensembles in the same vacuum chamber[138], and a single atomic ensemble analyzed in two subregions using camera imaging [139]. The highest reported stability at 1 s is , while the record long-term stability of an atomic ensemble is [139]. Higher stability allows for the estimation of systematic uncertainties in a shorter time, contributing to improving clock accuracy. An optical lattice clock with a fractional uncertainty of was reported in 2015 [75], and now the best performing clock has a fractional uncertainty of [78].
III.3.1 Major systematic uncertainties
For an optical clock to be highly accurate, it is essential to accurately estimate systematic shifts. Optical lattice clocks are typically stable enough that statistical uncertainty can be minimized by increasing the overall measurement time. As a result, their overall uncertainties are primarily limited by systematic uncertainties. Historically, improvements in the accuracy of optical lattice clocks have been enabled by a deeper understanding of systematic shifts. The pursuit of better uncertainty estimates have motivated the development of techniques to further mitigate these shifts. A notable example is the dc Stark shift. When optical lattice clocks were reported for the first time, the dc Stark shift caused by stray electric fields in a standard vacuum chamber were not a major contributor to total systematic uncertainty. However, as the overall uncertainty of optical lattice clocks improved, the impact of ambient electric fields became increasingly significant[140]. To precisely quantify this effect, the dc Stark shift was carefully measured with a chamber with electrodes inside [141]. In addition, to prevent charge accumulation on insulator surfaces, the inner surface of the viewports has conductive coating, typically indium tin oxide, which helped suppress stray electric fields.
One of the major sources of uncertainty in optical lattice clock is the BBR shift. This uncertainty arises from two factors: the amount of BBR experienced by the atoms and the uncertainty in the polarizability of the ground and particularly the excited state. To estimate the polarizability precisely, the transition matrix for the transition, which has the smallest transition frequency among the E1 transitions from the state, was carefully measured for Sr [78] and Yb [142]. To control the amount of BBR, a metallic shield with stabilized temperature is implemented inside the vacuum chamber. The temperature is calibrated to the 1-mK level for Sr[75] and Yb[125], confirmed by a Pt thermometer located near the atoms. In addition, the inner surfaces of both the vacuum chamber and the shield are coated with carbon nanotube to keep the emissivity of the chamber wall as close to 1 as possible [143]. Surrounding atoms with a cold shield inside a vacuum chamber reduces the amount of BBR at the atom location, as dictated by the Stefan-Boltzmann law. Such a cryogenic optical lattice clock at 95 K achieved a small BBR uncertainty [74].
A smaller uncertainty from the BBR shift can also be achieved by using atoms less sensitive to BBR. Hg, Cd, Mg, and Tm are attractive for this reason [144]. Clocks based on Hg[126], Cd[124], and Tm[145] have already been realized, while a Mg clock is under development, with the clock transition frequency and the magic wavelength already measured [123].
Another major source of uncertainty originates from the ac Stark shift caused by the trapping laser (Ref.[78] for Sr, [125] for Yb). When the total uncertainty was on the order of , the shift was mitigated by carefully selecting magic wavelength for Sr[70] and Yb[72]. However, with an orders of magnitude smaller uncertainty, higher order effects must also be taken into account. One such effect arises from vector and tensor shifts, which depends on the polarizations of the trapping lasers[146]. The shift proportional to the trap depth is formulated as
where , with , , and being the lattice polarization vector, lattice propagation vector, and lattice polarization ellipticity and being the direction of the quantization axis. , , and are the differential scalar, vector, and tensor shift coefficients. For a one-dimensional lattice, properly selecting the relative orientation between the polarization and the magnetic field can cancel some of these shifts. However, in a three-dimensional lattice, such cancellation is not possible in all three dimensions. This limitation is one of the reasons why state-of-the-art optical lattice clocks use one-dimensional optical lattice.
The ac Stark shift including nonlinear terms for a one-dimensional optical lattice is characterized as follows [147]:
where with being recoil energy for lattice light, is the quantum number for the longitudinal vibrational state, is the differential E1 polarizability between the ground and excited states of the clock transition, is the lattice laser frequency, is the detuning of the lattice laser frequency from the E1 magic frequency and and are the differential multipolarizability and hyperpolarizability, respectively. This formulation accounts for the effect of the longitudinal temperature of atoms in the lattice via and incorporates the shifts in Eq. 4. The radial temperature can be included by modifying as , where is the exponent on for each term in Eq. III.3.1. Using this formalism, the concept of operational magic intensity was itroduced [147]. However, despite these characterizations, the lattice ac Stark shift remains one of the major sources of uncertainty in the best optical lattice clocks. A careful analysis on the order of is reported for Sr [148], leading to a fractional uncertainty of [78]. Similar analysis is also going on for Yb [149]. An alternative approach to suppress the uncertainty from the ac Stark shift is to use atoms with small higher-order effects. Particularly, hyperpolarizability for Hg, Cd, and Zn is calculated to be smaller than for Sr and Yb [150]. These clocks can be potentially useful to suppress uncertainties related to the lattice ac Stark shift, in addition to the BBR shift.
The ac Stark effect also shifts the transition frequency due to the clock laser itself. Because the clock laser must be applied to drive the clock transition, minimizing this shift requires long irradiation times with low and uniform intensity. For such operations, inducing Rabi flopping is more effective than using a Ramsey sequence, which requires a separate interrogation time in addition to pulses. As a result, in state-of-the-art optical lattice clocks, a single pulse is commonly used to measure the frequency offset of the clock laser from the atomic resonance.
The density shift, which arises from atom-atom interactions, is another source of uncertainty. Since the shift is proportional to the density of atoms in the lattice, reducing the density helps mitigate its effects. Another way to control the density shift is by trapping atoms in a three-dimensional optical lattice and removing excess atoms by lowering the Fermi temperature of a degenerate Fermi gas[151]. However, both approaches have drawbacks. Reducing the density of atoms increases the QPN due to the small number of atoms, while the three-dimensional optical lattice induces a large uncertainty in the ac Stark shift due to the tensor shift. Additionally, the edge effect [152] needs to be carefully managed. Recent studies utilize low-density fermionic gas in large vertical one-dimensional optical lattices. First, a large optical lattice, e.g., a beam waist size of 260 m [139], is used to reduce the density of atoms. To achieve such a lattice, trap laser intensity is enhanced with an intermediate finesse cavity ([153], [139]). Second, fermionic isotopes naturally suppress the collisional shift by s-wave collisions. For Sr, the remaining shift due to p-wave collisions is canceled by inter-layer s-wave collisions, which is induced by the delocalized atoms over a few layers due to the low trap depth () [139]. This approach has enabled the clock to reach fractional uncertainty [78]. For Yb, delocalization of atomic cloud to reduce the density in this large one-dimensional optical lattice is performed. [154] Other systematic uncertainties, such as background gas collision shift, second order Zeeman shift, and servo error are well-controlled and significantly suppressed compared to the major uncertainties discussed here.
III.3.2 Varieties of optical lattice clocks
As optical lattice clock technology advances, systems designed for various applications are also developed. These systems extend beyond the conventional setup, where a single atomic ensemble is loaded into a single location in an optical lattice.
Transportable optical lattice clocks have been developed for applications such as geodesy testing and clock comparisons at remote locations. One example is a laboratory built on a trailer [155]. More compact versions have been used to test the validity of general relativity at an observation tower [156], and efforts to make it even smaller are ongoing. Clock comparisons across continents using such transportable clocks were recently reported [157, 158].
Clock comparisons are also performed using fiber links. Metrology laboratories around the world are linked via optical fiber networks, with the longest network located in Europe [159, 160]. These links enable precise clock comparisons between remote national metrology institutes[160], and facilitate comparisons between different types of clocks with unprecedented accuracy [161]. They are also used to measure the height difference based on time delay [162].
For differential comparisons between atomic ensembles, multiple ensembles can be trapped in a single one-dimensional optical lattice. One approach involves using a moving optical lattice[138], allowing atoms to be trapped at up to six spatially distinct locations. These relative measurements have been used to test general relativity [163] and enhance clock stability [164]. In a fixed optical lattice, two ensembles can be trapped by shifting the position of the MOT [165].
Continuous interrogation helps suppress the Dick effect [135, 136], where pulsed interrogation of the atomic system introduces instability at frequencies corresponding to the period of the experiment cycle. In addition, continuous interrogation enables feedback on the phase, causing the Allan deviation to scale rather than . This allows for a faster estimation of systematic shifts. The continuous generation of an ultracold atomic beam has been reported as a step toward developing a continuous optical lattice clock [166].
The rapidly advancing technology of optical tweezer arrays is also applied to Sr optical atomic clocks [167, 168]. A stability study using 88Sr atoms has been reported[101], though characterization of the uncertainties has not been documented. The ability to control individual atoms is advantageous for generating entangled states, which can further enhance stability, as discussed in the next section.
III.3.3 Enhancement of stability by entanglements
The atomic clock system is one of the few systems where entanglement has been demonstrated to enhance their performances. The first proof-of-principle experiments were performed with microwave clocks using 87Rb [169] . Later, larger stability enhancement was achieved with a system where atoms were uniformly coupled to a cavity mode [170]. These studies utilized spin-squeezed states as the source of entanglement. The first spin-squeezed optical clock was demonstrated with Yb[99]. Spin squeezing was performed on the spin-1/2 system in the ground state of 171Yb [171] and then transferred to the excited state of the clock transition to improve clock performance. The study reported a 4.4 dB improvement below the SQL in the stability of the atomic system with a stability of , limited by the frequency fluctuations of the clock laser. An improvement in the clock stability was demonstrated at the stability of [172] with a Sr optical lattice clock. Note that the spin squeezing described here was performed in a cavity quantum electrodynamics (QED) system, where atoms were located in a high-finesse cavity that enhances their coupling to light.
The enhanced stability of optical clocks using Greenberger-Horne-Zeilinger (GHZ) states has been demonstrated with a tweezer array clock.[173] The GHZ state is generated through a strong atom-atom interaction via a Rydberg state. These strongly entangled atoms exhibit high phase sensitivity, leading to improved clock performance when the Ramsey sequence is used. However, because the GHZ and other entangled states are more sensitive to decoherence, atomic coherence time decreases, resulting in shorter interrogation time compared to state-of-the-art optical lattice clocks. This negates the advantage of the reduced QPN. The overall stability of the clock does not necessarily improve with entangled states. Nevertheless, the shorter interrogation time decreases the duration of a single clock cycle, increasing the bandwidth of clock measurements. This increase can be beneficial when the interrogation time is constrained, such as in fountain clocks, or in searches for ultralight dark matter with relatively large masses (see Section VII).
III.4 Ion clocks
III.4.1 State-of-the-Art Systems
In a conventional ion clock, a single ion is trapped in a Paul trap, which uses an oscillating quadrupole electric field to confine charged particles[174, 175, 176]. Since Paul traps have been in use longer than magic-wavelength optical lattices, ion clocks were developed as the frequency standard before optical lattice clocks, as shown in Fig. 2. Although the 1-s stability for the ion clock is not as high as that of optical lattice clocks, their ability to operate over long periods allows their accuracy to be comparable to that of optical lattice clocks.
The ionic species with the smallest fractional uncertainties have traditionally been Al+ [65] and Yb+ [64]. Lu+ [131], and In+ [79] recently reached the fractional uncertainties of similar level. Al+, In+ and Lu+ have alkaline-earth-atom-like electronic structures, while Yb+ has an alkali-atom-like structure with an E3 transition, where an electron in the orbital is excited. Alkali-atom-like structure with E2 transitions, such as Sr+ [103], Ca+ [102], Ba+ [104], and Ra+ [105] have also been investigated. Some of these clocks are summarized in Table 1. For certain ions, the broad-linewidth transitions suitable for laser cooling have wavelengths too short for readily available laser systems. In such cases, laser cooling is achieved through sympathetic cooling by trapping another ion in the same Paul trap. The preparation and the detection of the quantum state are also carried out using this second ion through the Coulomb coupling of their motional states, which is known as quantum logic spectroscopy [177]. For example, in an Al+ clock, the Al+ ion is sympathetically cooled by a Mg+ ion, and quantum logic spectroscopy is performed via the Mg+ ion [65].
III.4.2 Uncertainties
State-of-the-art ion clocks have fractional uncertainties around , with their accuracy limited by several factors. One of the major sources of systematic uncertainty is the micromotion-related Doppler shift, which arises from the ion’s micromotion in a Paul trap. In some clocks, this uncertainty is the dominant contributor to the overall uncertainties [65, 64, 103]. Suppressing this effect can be achieved by optimizing the Paul trap’s configuration. The high performance of the best Al+ ion clock was made possible by designing a new electrode configuration that minimizes micromotions [65], offering a significant improvement over previous-generation setups.
The BBR shift is another dominant source of uncertainty [63, 103, 102, 65, 64, 79] and has been extensively studied. For Lu+[178] and Ca+[179], dedicated studies evaluating the BBR shift have been published. A cryogenic clock is realized for a Ca+ ion clock [102].
With respect to the ac Stark shift, the probe-induced ac Stark shift can also be a significant source of uncertainty[64]. This is particularly relevant for ion clocks that use E2 or E3 transitions as the clock transition, as these require much higher laser intensity compared to the , which is weakly mixed with an E1 transition by the nuclear magnetic moment. To mitigate the ac Stark shift induced by the clock laser, the hyper-Ramsey sequence, where pulse lengths are adjusted to cancel the shift, is advantageous[180]. This sequence is particularly beneficial for the spectroscopy requiring high intensity, such as transitions without any mixing with an E1 transition and two-photon transitions. The hyper-Ramsey sequence was experimentally investigated in an Yb+ ion clock using the E3 transition [181]. Even with its application in Ref. [64] the probe-induced ac Stark shift remains a major source of the uncertainty.
Depending on the atomic species and system architecture, other effects can also contribute significantly to uncertainty. The quadratic Zeeman shift is sometimes a major factor[65, 79]. In the case of In+ [79], the uncertainty arises from the relatively large uncertainty in the -factor difference. Al+ has a coefficient for the second-order Zeeman shift more than 10 times larger than that of In+ [182]. The In+ clock identifies thermal time dilation as its major source of uncertainty [79], which can be mitigated by using a narrow-linewidth transition for laser cooling. Other factors, such as background gas collision and servo error, which are common sources of uncertainty in the optical lattice clocks, also contribute to the overall uncertainty.
The main disadvantage of ion clocks is their large QPN due to the small number of ions in the trap, unlike optical lattice clocks which contain a much larger number of neutral atoms. The ultimate solution to reducing QPN is to increase the number of trapped ions. Multi-ion clocks have been investigated both theoretically and experimentally. Theoretically proposal exist for both Paul traps [183] and Penning traps [184]. In Paul traps, a path towards uncertainty below is demonstrated, and the scaling of QPN is reported [79]. A key source of uncertainty in multi-ion clocks is the quadrupole shift, which arises because ions are not located at the center of the Paul trap. As a result, each ion experiences a different quadrupole electric field, leading to inhomogeneous broadening of the transition. This effect is particularly problematic for ions with large quadrupole moments, such as Ca+, Sr+, and Ba+ [185]. However, Al+ and In+ are much less affected due to their small quadrupole moments [186]. Suppression of this quadrupole shift has been demonstrated using dynamic decoupling [187]. Also, quantum logic spectroscopy has been successfully applied to multiple ions[188].
When comparing an ion clock with an optical lattice clock, carefully designed sequences can enhance stability. By coherently linking the zero-dead-time interrogation of Yb optical lattice clocks to an Al+ ion clock, the stability of the frequency comparison was improved by an order of magnitude [189]. This enhancement was achieved by extending the interrogation time of the Al+ ion clock, thereby improving the SQL. The longer interrogation time was made possible by the long coherence time of the clock laser, which was maintained through frequent phase corrections without deadtime by the interrogation of the Yb optical lattice clocks.
IV Variation of fundamental constants
IV.1 Theoretical basis
The constancy of fundamental constants is one of the fundamental assumptions in modern physics. If a fundamental constant were to vary over time or space, it would violate the local position invariance, which states that the result of any local non-gravitational experiment is independent of the time and location of the experiment. Local position invariance is part of Einstein’s equivalence principle, along with the weak equivalence principle and local Lorentz invariance. The weak equivalence principle asserts the equivalence of gravitational and inertial mass, while local Lorentz invariance states that the outcome of any local non-gravitational experiment is independent of the velocity of the free-falling reference frame where it is performed. Some review papers [190, 191] discuss Einstein’s equivalence principle in detail.
To observe any variation of fundamental constants, they must be dimensionless, because a reference is always required for such measurements. For example, the time variation of the proton mass or the electron mass cannot be measured independently, but changes in the proton-to-electron mass ratio can be tested. Other important dimensionless fundamental constants include coupling constants for fundamental interactions, such as the fine structure constant .
Historically, the possibility of the variation of fundamental constants was first proposed by Dirac[192]. Nowadays, several mechanisms in BSM physics can induce such variations in the coupling constants of fundamental interactions. In theories with extra dimensions, such as Kaluza-Klein models [193], superstring theories [194], and brane-world models [195], the fundamental constants remain uniform across all dimensions but appear effectively constant in the observable -dimensional world. In addition, a field introduced to explain the inflation theory can couple to fundamental constants. This field is typically a scalar field, due to its simplicity and its ability to acquire a vacuum expectation value without violating local Lorentz invariance. [196]
Searches for variations of fundamental constants have a long history. Early investigations were performed through astronomical observations [197] and studies of the natural nuclear reactor at Oklo[198]. Since than, extensive astronomical searches for variations of fundamental constants have been long ongoing[196]. Particularly important were observations based on atomic absorption lines which suggested that may have had a different value in the region of universe with large redshift . A study in 2001 [199] reported over , where is the variation of , wit no identified systematic effects that could account for the results. Subsequent analyses incorporating additional data have supported spatial variation of with significance (see Ref.[200] and references therein). However, recent observations at large themselves are consistent with no variation [200, 201, 202]. If this spatial variation extends to the scale of the solar system, a variation of yr-1 is expected due to the motion of the solar system in the universe [203]. Comparisons using atomic clocks provide a valuable cross-check against astronomical searches, as they allow for better control of environment of atoms than astronomical searches.
The search for time variations of fundamental constants on Earth is performed by comparing the transition frequencies of two different clock transitions over a long period. Because the dependence of atomic transitions on fundamental constants varies between transitions, the ratio of their transition frequencies has a nonzero dependence on fundamental constants. Consider as an example. The binding energies of energy levels in hydrogen with a principal quantum number scale as . The fine structure, which arises from L-S coupling, is proportional to . As a result, the frequency ratio of these two energy scales is proportional to . For precise frequency measurements, narrow-linewidth transitions used in atomic clocks are more suitable than broad-linewidth transitions in hydrogen. In multi-elecron atoms, the dependence of a transition is estimated by theoretical calculations of the electronic structure and is characterized by a constant in the following equation:
where is the sensitivity parameter to variations of .
Table 2 summarizes for major transitions used in atomic clocks, along with some transitions proposed for future measurements that have particularly large K. For neutral atoms and singly-ionized ions, relative to the ground state is at most . HCIs (see Sec. V) are necessary to obtain . The nuclear clock transition in 229Th is predicted to be highly sensitive to the variation of [204]. (see Sec. VI)
Atom | Transition | (nm) | Ref. | |
Rb | ground state hyperfine | RF | [205] | |
Cs | ground state hyperfine | RF | [205] | |
Al+ | 267 | [206] | ||
Sr | 698 | [206] | ||
Sr+ | 674 | [207] | ||
Yb | 578 | [206] | ||
Yb | 507 | [117] | ||
Yb | 431 | [117] | ||
Yb | 1695 | [118] | ||
Yb | 2875 | [208] | ||
Yb+ | 411 | [207, 209] | ||
Yb+ | 436 | [207, 209] | ||
Yb+ | 467 | [210] | ||
Hg+ | 282 | [210] | ||
Pr10+ | 2650(49) | [211] | ||
Pr10+ | 1409(14) | [211] | ||
Sm14+ | 2614(470) | [211] | ||
Ir17+ | 1978 | [212] | ||
Cf15+ | 812 | [213] | ||
Cf16+ | 1899 | [214] | ||
Cf16+ | 1030 | [214] | ||
Es16+ | 1430 | [213] | ||
229Th | nuclear clock transition | 148 | [204] |
IV.2 Status of experimental searches in laboratories
The historical advancement in the constraint on the time variation of by clock comparisons is summarized in Fig. 4. The constraint on the time variation of is described as the upper bound on . Before the data points shown in Fig. 4, an early report in 1976 set a constraint of yr-1 [215]. Another report combined a clock comparison and astrophysical observations and set the constraint of yr-1 in 1993 [216]. Since 1995, the constraint has rapidly improved with the development of atomic clocks. At an early stage, the Hg+ clock played a significant role [217, 218, 90, 219, 61], and recent best constraints have been obtained using Yb+ ion clocks [220, 221, 222, 223], both of which have large (see Table 2). Because the constraints are estimated for a linear drift, frequency ratio measurements over a long period enhance the sensitivity. The recent reports using Yb+ ion clocks utilize the frequency ratio data spanning more than 10 years[223]. The best constraint reported so far, /yr [223], is on the same order of magnitude as the spatial variation reported in astronomical observations, when converted according to the motion of the Earth [203].
To increase the sensitivity to the time variation of in a fixed amount of time, either the stability of an atomic clock or has to be enhanced. Notably, the state in neutral Yb can increase sensitivity [117, 118, 208]. Although for this transition is not the largest in Table 2, optical lattice clocks have higher short-term stability than ion clocks, which makes this state beneficial for enhancing sensitivity, especially for short-term variations. Note that the large ’s in the and transitions are mainly due to their small transition energy, and the excited state is the same as the transition. In this case, the overall sensitivity for the variation of needs to be carefully estimated for unstable clocks. Ref. [117] argues that the enhancement in is cancelled by the clock instability, if the instability is mainly due to noise on the . To take advantage of transitions with large , transitions in highly charged ions and 229Th are promising.
Time variation of is also widely investigated. Similar to , the dependence on varies between transitions and is calculated theoretically. Refs. [220, 221] performed a combined analysis of the constraints on and , resulting in yr-1 and yr-1, respectively, both consistent with zero. More stringent constraints of yr-1 and yr-1 have been reported from frequency comparisons between Yb, Sr, and Cs clocks [228] and between the E3 transitions in Yb+ ion and a Cs clock [222], respectively. In addition, molecules are used to investigate time variation of because their internal degrees of freedom exhibit different dependence on . The energy differences for purely electronic, vibrational, and rotational transitions , , and scale with as follows [229]:
Laboratory-based studies employed SF6 molecules [230] and ultracold KRb molecules [231]. The most stringent constraint based on molecular measurements is /yr [231], which is less precise than the best constraint from the Yb+ ion clock. However, these reports provide a valuable cross-check for atomic clock comparisons.
V Spectroscopy of highly charged ions
V.1 Properties of highly charged ions
A major source of systematic uncertainty in optical clocks arises from the ac Stark shift, which is caused by the deformation of the electronic cloud around the nucleus due to an external electric field. This effect can be mitigated by reducing the deformation, which is achieved by holding electrons more tightly within the atom. Such strong confinement of electrons can be achieved in a highly charged ion (HCI) [232]. Qualitatively, by stripping away loosely bound the outer-shell electrons, only those with strong binding remain, leading to reduced susceptibility to external electric fields. Optical clocks based on HCIs have been proposed as a next-generation clock capable of surpassing the accuracy limit of both optical lattice clocks and ion clocks [17].
The qualitative behavior of energy levels in HCIs can be understood through the dependence of various energy scales, where is the atomic number, as summarized in Table 3. For hydrogen-like ions, the scaling of physical quantities can be derived from the Schrödinger equation and other exact solutions. This calculation shows that the energy difference for the transition increases proportional to , while the ac Stark shift decreases with scaling. In HCIs, the dependence arises from three different charges, , , and , where is the effective charge that maintains the energy level scaling of an external electron in the HCI at and is the charge of the ion. The scaling summarized in Table 3 is consistent with hydrogen-like ions when .
A key tradeoff of HCI clocks compared to optical lattice clocks is the number of ions that can be interrogated simultaneously. HCI clocks typically trap a single HCI,resulting in lower clock stability due to QPN. A HCI is more susceptible to deionization from collisions with residual gas than a singly charged ion, which potentially can limit the duration of long-term operation.
hydrogen like ion | highly charged ion | |
binding energy | ||
fine structure | ||
hyperfine structure | ||
Lamb shift | ||
ac Stark shift | ||
BBR shift | ||
electric quadrupole shift |
Beyond their role in next-generation atomic clocks, HCIs are also useful for various fundamental physics searches. One example is testing QED in strong fields. The intense electric field experienced by electrons in U90+ is used to probe QED in extreme conditions[234]. In addition, HCIs exhibit higher sensitivity to the variation of than neutral atoms. This is because high-velocity electrons in HCIs are affected more by relativistic effects, which scale strongly with . Following a pioneering work [235],various ionic species have been proposed as suitable candidates for searching for the variation of . Several transitions with large are listed in Table 2, with additional transitions provided in Table 5. Transitions with longer wavelengths tend to have large , as explained by Eq. 6. Another important application is the search for fifth force between an electron and a neutron (See Sec. VIII for further details). Due to their significantly different electronic structures compared to neutral atoms, HCIs offer high sensitivity for detecting new forces through isotope shift measurements [236].
ion | wavelength (nm) | (s-1) |
171Yb69+ | 2160 | 0.43 |
195Pt77+ | 1080 | 3.4 |
199Hg79+ | 1150 | 2.8 |
203Tl80+ | 338 | 111.2 |
205Tl80+ | 335 | 114.2 |
207Pb81+ | 886 | 6.2 |
To perform laser spectroscopy on HCIs, both the atomics species and must be properly selected. The scaling of binding energy suggests that electronic transitions between different states, which typically occur in the visible-light range for neutral atoms, shift into the X-ray region for HCIs. Instead, energy differences between hyperfine or fine structures increase to a range suitable for laser spectroscopy. For hyperfine structures, hydrogen-like ions of heavy atoms exhibit eV energy differences. Table. 4 summarizes ions with nuclear spin and transition frequencies accessible by lasers. For fine structures, eV energy differences are realized at . An example is the energy-level structure of Pr10+, as shown in Fig. 5, where the 255.6 nm transition is between two fine-structure levels. The two states are another pair of fine structures. The transition between J=3 and J=2 states in the state of Cf16+, listed in Table 5, is also a transition between fine structures.
Naively, energy differences between different orbitals exceed the energy for visible light. However, an exceptional case where visible light can drive electronic transitions in HCIs is known as level crossing [17]. Fig. 6 provides an intuitive illustration of this phenomenon. In a hydrogen atom, different states for the same level are degenerate, where is the azimuthal quantum number. However, in multielectron atoms, the states are ordered from the low energy as . Consider, for example, the energy of the and orbitals. In a hydrogen-like ion, the orbital has a higher energy than the orbital, whereas in a neutral atom, the orbital has higher energy. At a certain in HCIs, the energy difference between these two states reverses sign. Around this , the energy of the two orbitals is nearly the same, as shown in Fig. 6, sometimes falling into the visible light range.
In the case of Pr, level crossing occurs around . As shown in Fig. 5, Pr10+ exhibits two low-energy transitions between the ground state and the states, in addition to a transition between fine-structure levels. The two states have long lifetimes, making them suitable for clock applications. Transitions from these states to the ground state have high sensitivity to the variation of (see Table 2 for a comparison between neutral atoms and singly-ionized ions).
Atom | configuration 1 | J | configuration 2 | J | (nm) | (s) | Ref. | |
Pr10+ | 1/2 | 5/2 | 2650(49) | [238] | ||||
Pr10+ | 1/2 | 7/2 | 1409(14) | [238] | ||||
Nd10+ | 2200(430) | [211] | ||||||
Nd10+ | 1480(240) | [211] | ||||||
Pm14+ | 1/2 | 7/2 | 966 | [235] | ||||
Sm13+ | 494(22) | 0.367 | [211] | |||||
Sm13+ | 444(18) | 0.133 | [211] | |||||
Sm13+ | 386(14) | 0.141 | [211] | |||||
Sm14+ | 2614(470) | 8.515 | [211] | |||||
Sm14+ | 1182(100) | 0.556 | [211] | |||||
Eu14+ | 7/2 | 11/2 | 1657(330) | [239] | ||||
Eu14+ | 7/2 | 13/2 | 971(130) | [239] | ||||
Ho14+ | 420 | 37 | [240] | |||||
W8+ | 1646 | [212] | ||||||
W8+ | 1573 | [212] | ||||||
W8+ | 899 | [212] | ||||||
Ir16+ | 267 | [212] | ||||||
Ir17+ | 283 | [212] | ||||||
Ir17+ | 1978 | [212] | ||||||
Pt17+ | 402.2 | [241] | ||||||
Pu8+ | 0 | 1 | 654 | [242] | ||||
Pu9+ | 3/2 | 5/2 | 1172 | [242] | ||||
Pu9+ | 3/2 | 7/2 | 872 | [242] | ||||
Pu9+ | 3/2 | 3/2 | 862 | [242] | ||||
Pu10+ | 2 | 3 | 1989 | [242] | ||||
Pu10+ | 2 | 2 | 866 | [242] | ||||
Pu10+ | 2 | 4 | 817 | [242] | ||||
Pu11+ | 5/2 | 3/2 | 1461 | [242] | ||||
Am9+ | 1 | 0 | 974 | [242] | ||||
Am9+ | 1 | 4 | 905 | [242] | ||||
Bk15+ | 0 | 3 | 483 | [242] | ||||
Bk15+ | 0 | 2 | 340 | [242] | ||||
Cf15+ | 812 | 6900 | [213] | |||||
Cf16+ | 3 | 0 | 1899 | [214] | ||||
Cf16+ | 3 | 2 | 1638 | 0.178 | [214] | |||
Cf17+ | 5/2 | 1/2 | 535 | [214] | ||||
Es16+ | 1430 | 16000 | [213] | |||||
Es17+ | 1343 | 11000 | [213] |
V.2 Experimental methods for precision spectroscopy of HCI
Early HCI spectroscopy was performed by classifying the astronomical observations and detecting radiation in fusion researches, including tokamak experiments [243]. These spectroscopic studies primarily focused on the X-ray region and involved mixtures of multiple ionic species. To enable spectroscopy of specific ions, the development of ion sources was essential. The earliest sources of HCIs were based on electron-cyclotron resonance ion sources [244], followed by the invention of electron beam ion sources[245]. Early optical spectroscopy of HCIs was performed in 1987 for fine structure transitions [246], and in 1994 for a hyperfine structure transition [247].
Nowadays, HCIs are generated by electron beam ion traps (EBITs) [248]. An EBIT is an electromagnetic ion trap formed by a strong electron beam and electrodes. The negative charge of the electron beam confines the ions radially, while axial confinement is achieved by electrodes. The energy of the electron beam determines the maximum number of electrons that can be stripped from neutral atoms.
The generated HCIs is extracted from an EBIT and then transferred to a Paul trap. At this stage, the HCIs are hot and must be cooled toward its motional ground state for precision spectroscopy. Due to the absence of broad-linewidth transitions suitable for laser cooling in HCIs, cooling is typically achieved through sympathetic cooling. This method, along with high-precision spectroscopy, has been demonstrated in Ar13+ [249]. In this experiment, an Ar13+ ion is stopped by a Coulomb crystal of Be+, and subsequently laser-cooled using a broad transition in Be+. The experiment is conducted in a cryogenic trap to suppress the BBR shift [250]. The transition at 441 nm serves as the clock transition, with its transition frequency measured to a fractional uncertainty of , referenced to the state-of-the-art 171Yb+ clock. The dominant source of uncertainty in this measurement is the statistical uncertainty, while the total systematic uncertainty for the Ar13+ system is . The primary contributor to systematic uncertainty arises from micromotions, which can be suppressed in the future work [251].
Other ionic species are also investigated, motivated by theoretical proposals. One of the extensively investigated species is Pr ions. Pr9+ was first experimentally examined in Ref.[252]. Although for the state is smaller than ’s listed in Table 5, energies of the long-lived states have been determined with a fractional uncertainty. Pr10+ also has an early spectroscopy report [238]. Combined with an improved theoretical calculation of the energies of the metastable excited states [211], the energy of the state was determined with a fractional uncertainty. Transition frequency measurements were also performed for a narrow-linewidth transition in Ni+12 [253] and many charge states for Xe [254] and Ca [236] ions. The investigations of Xe and Ca ions are motivated by the search for the fifth force through isotope shift measurements (see Sec. VIII). In particular, Ca14+ was recently utilized to set a new constraint on the fifth force[255].
VI Nuclear clock
An even more rigid object against external electric fields than the inner-shell electronic cloud is the nucleus. However, using a typical nuclear transition at keV for precision spectroscopy with a laser is highly challenging. Exceptionally, a low-lying nuclear isomer state in 229Th is barely accessible by laser excitation from the ground state. The existence of this extremely low-energy nuclear excited state was first proposed by Kroger and Reich [256]. The idea of using this isomer state as a clock was initially proposed in 2003 [18]. At the time, the energy of this transition was known with only one-digit precision and was estimated to be lower than its currently known value. However, in recent years, substantial advancements have been made. The transition frequency is now determined to be 2 020 407 384 335(2) kHz, corresponding to an energy of 148.38 nm and 8.3557 eV, with a fractional uncertainty of . Additional details of the transition are summarized in Fig. 7. Note that the next lowest-energy nuclear excited state known today is in 235mU, which is 76.7 eV above the ground state. With this state in 229Th, a clock can be operated based on a nuclear transition. Because the nucleus is times smaller than the atom, it is nearly unaffected by the ac Stark shift.
Historical development of the energy estimates for the first excited state of 229Th is summarized in Table 6. Initially, the energy of the excited state was measured by low-energy gamma rays emitted during the decay of 233U to 229Th [257, 258, 259, 260]. More recent, precise measurements utilized internal conversion and X-ray excitation of the nucleus to a higher excited state than the one involved in the clock transition [261]. In addition, spectroscopy of photons emitted during the relaxation of the clock state further reduced the uncertainty in the transition energy[262], improving the precision of the transition frequency to 4 digit precision. A breakthrough occurred when laser excitation of the transition and subsequent decay of the excited state was observed in 229Th4+ ions doped in a CaF2 crystal [263]. Soon afterthat a similar observation was reported with Th-doped LiSrAlF6 [264]. These studies measured the wavelength with 7 digit precision, confirming that the observation was not crystal specific but resulted from 229Th interacting with the light field. A few months later, spectroscopy using the seventh harmonic generation of a frequency comb determined the transition energy with 2 kHz uncertainty [265]. The observed lifetime of the excited state is 1770(11) s. While this long lifetime provides excellent coherence for long interrogation time, it also makes quick manipulation challenging. To address this, methods to shorten the lifetime were explored [266, 267]. Further studies have also investigated the fine structure [204] and thermal effects [268]. In the field of optical lattice clocks and ion clocks, well-established methodologies can reduce the uncertainty of resonant frequency measurements from 1 kHz to Hz. However, a key difference between these optical clock systems and the 229Th nuclear clock is the lack of a stable continuous-wave laser operating at the resonant frequency of the clock transition.
The nuclear clock can be implemented in two different ways. To prevent the 229Th nucleus from rapidly decaying via internal conversion, the 229Th atoms must be ionized into 229Th3+ ions. These ions can be trapped using an ion trap similar to those used in HCI clocks. This approach has the advantage of leveraging well-established spectroscopy techniques from ion clocks, making it easier to estimate systematic uncertainties. Recently, Laser spectroscopy of 229Th3+ ions in a Paul trap was demonstrated [269]. However, due to the small number of ions in the system, the QPN is large. An alternative method, unique to nuclear clocks, is doping a crystal with 229Th3+ or 229Th4+ ions. Because nucleus is highly isolated from the surrounding electrons, the effects being in a crystal are minimal compared to those on atoms. This approach allows for a large number of ions, [264], to be incorporated, which helps suppress QPN. However, a recent study showed a strong temperature dependence of the transition frequency [268], attributed to changes in the electric field induced by the thermal expansion of the lattice constant.
Year | Energy (eV, otherwise noted) | Ref. |
1976 | <100 | [256] |
1990 | [257] | |
1990 | <5 | [270] |
1994 | 3.5(1.0) | [258] |
2005 | 5.5(1.0) | [271] |
2007 | 7.6(5) | [272] |
2019 | 8.28(17) | [261] |
2019 | 8.30(92) | [259] |
2020 | 8.10(17) | [260] |
2023 | 8.338(24) | [262] |
2024 | 8.367(24) | [273] |
2024 | 8.355 74(3) | [263] |
2024 | 8.355 733(10) | [264] |
2024 | 2 020 407 384 335(2) kHz | [265] |
For fundamental physics applications, the nuclear clock is expected to be highly sensitive to variation of fundamental constants. Because nuclear structure depends on the coupling constant of the strong force , it should be sensitive to variation of [274]. In addition, theoretically predictions suggest that it also exhibits high sensitivity to the variation of , with [204]. Intuitively, this large arises because the small transition energy results from a cancellation between a large repulsive Coulomb energy and an equally large attractive energy from the strong force. The large also makes the nuclear clock a promising tool for searching for ultralight dark matter (see Sec. VII). Another potential research direction involves developing a system that operates without vacuum chamber [275]. Such a system could be advantageous for compact portable optical clock.
VII Ultralight dark matter
VII.1 Theoretical basis
Dark matter is massive objects in the universe that do not interact with light. It was first proposed to explain the rotation curves of galaxies [276]. Today, multiple observations strongly support its existence, such as gravitational lensing [277] and fluctuations in the cosmic microwave background [278]. Dark matter is estimated to be five times more abundant than ordinary matter.
Various dark matter candidates have been examined, e.g., massive compact halo objects [279] with astronomical mass scales, weakly interacting massive particle typically searched for in the mass range from 0.1 GeV to 1000 GeV [280], and axions and axion-like particles, which are generally investigated in the sub-1 eV mass region [280]. Among these candidates, particles with masses of 1 eV or smaller have to be bosonic due to their expected number density, the energy density of dark matter, and the Pauli exclusion principle. This class of low-mass dark matter is referred to as ultralight dark matter [281, 282]. Searches for ultralight dark matter are conducted using various methods including astronomical observation [283, 284, 285, 286, 287], microwave technology [288, 289, 290, 291], laser-based techniques [292, 293], accelerometers [294, 295, 296], and nuclear EDM searches [297], among others. Notably, atomic clocks are particularly sensitive to topological dark matter, which consists of spin-0 particles forming topological defects in the universe and coupling to non-derivative terms in the SM Lagrangian[298]. Note that axions and axion-like particles are also important candidates for ultralight dark matter. However, because these particles are extensively reviewed elsewhere [13, 14, 15, 16], the discussion here is limited to the spinless particles.
Searches for ultralight dark matter using atomic clocks are sensitive to mass ranges whose the corresponding energy scale is the same as or smaller than the single cycle of the clock operation, which is s. This corresponds to a dark matter mass of eV/c2 or lower. The lower bound for this mass is determined by the overall duration of clock operation. From a cosmological perspective, dark matter is generally expected to have a mass greater than eV in order to account for all observed phenomena with a single candidate. This is because, at this mass, the de Broglie wavelength of dark matter exceeds the size of dwarf galaxies. However, if dark matter consists of multiple components, mass ranges below eV may still be relevant for investigation.
The first proposal for searching for ultralight dark matter using atomic clocks focused on detecting topological dark matter with quadratic coupling to SM particles [299]. This concept initially relied on GPS satellites [300] and required collaboration between multiple research groups across different geographic locations for experimental realization [301]. Later, a new approach was proposed to detect ultralight dilaton dark matter, putting emphasis on the linear coupling to fundamental constants [298]. This proposal included clock comparisons within a single location. After this proposal, various experimental efforts and theoretical analyses of existing data were conducted to put constraints on the ultralight dark matter, starting with a constraint derived from Dy spectroscopy [302].
For dark matter that couples to fundamental constants, a scalar field is typically assumed. The interaction between this field and ordinary matter is described by the following Lagrangian density [298, 303, 304]:
with G being the Newtonian constant of gravitation and being Planck mass, is the electromagnetic field tensor, is the intensity tensor for the gluon field, is the coupling constant for quantum chromodynamics (QCD), is the function for the running coupling constant for , are the mass of an electron, an up quark and a down quark, respectively, is the anomalous dimension characterizing , and is the spinor for fermions. is integer, and correspond to linear and quadratic coupling, respectively. Some studies parametrize the coupling constant as , in which case is associated with the energy scale for new physics [299, 304]. Each term in Eq. VII.1 describes a tree-level interaction, where a SM particle emits an ultralight dark matter particle, with the coupling constant for this interaction being . Well-known particles that satisfy these properties are the dilaton and moduli. Similar interactions can be assumed for axions when CP symmetry is violated [305].
When couples to ordinary matter, fundamental constants vary according to the coupling constants in the Lagrangian in the following way.
Here, for , and is the energy scale for QCD, which dominates the mass of a proton. The dark matter field oscillates at a frequency (: the mass of the ultralight dark matter) with an amplitude . The corresponding dark matter density is given by . In general, atomic transitions depend on fundamental constants in different ways. These dependences can be numerically calculated for multi-electron systems to evaluate the sensitivity of various transitions to the variation of fundamental constants.
Searching for ultralight dark matter through this interaction is relatively easy to implement in a laboratory because atomic transitions and the size of an atom depend differently on fundamental constants [306, 307]. Take as an example. The wavelength of a narrow-linewidth laser used for precision spectroscopy is stabilized by locking it to a stable reference cavity, where the mirror separation is determined by the spacer material. Since the laser wavelength is proportional to , the frequency scales as , where is the Bohr radius. However, an atomic transition does not generally follow the same -dependence, as discussed in Sec. IV. This difference in -scaling leads to a periodic variation in the atomic resonance relative to the cavity resonance when couples to the ultralight dark matter field .
Three other fundamental constants can also influence the atomic transition frequency. In general, when the frequencies of two atomic transitions A and B are monitored, their ratio is depicted as
where denotes the change in and ( are the sensitivity coefficient for . Because four fundamental constants can vary while only a single frequency ratio is measured experimentally, setting a constraint on one fundamental constant requires assuming that the coupling constants for the other three are zero.
VII.2 Status of experimental searches
At the end of an experimental analysis, the amplitudes of a sinusoidal oscillation in the frequency ratio are obtained. The frequency assumed for the sinusoidal fit can be arbitrarily chosen within the range up to the Nyquist frequency. As a result, a single dataset for measured over time can provide constraints on dilaton ultralight dark matter across a wide mass range, as shown in Fig. 8. So far, no experimental results detected sinusoidal oscillations with amplitudes significantly larger than their statistical uncertainties. The resulting constraints exclude coupling constants stronger than a certain threshold for a given mass. In Fig. 8, these excluded regions are represented as shaded area. In this kind of searches for new particles, it is common practice to plot mass on the horizontal axis and a coupling constant for a specific interaction on the vertical axis.
Figure 8(a) shows the constraints on . Because most electronic transitions exhibit finite sensitivity to the variation of , state-of-the-art atomic clock comparisons have significantly contributed to improving these constraint [161, 311, 223]. Notably, the result from PTB [223] surpasses previous reports by an order of magnitude by combining two pairs of clock comparisons.
Comparisons between optical clocks have no sensitivity to , primarily because all optical transitions are more or less proportional to , resulting in a constant when frequency ratio is taken. Instead, comparisons between an electronic transition or a reference cavity and a microwave transition between hyperfine levels exhibit a high sensitivity. The comparison between an Si cavity and a hydrogen maser [309] reported constraints down to eV. The lower bound on mass is limited by the measurement duration. A long-term comparison between an Yb optical lattice clock and a Cs fountain clock, spanning 298 days, provided constraints for smaller masses [308].
For and , the contribution of clock comparisons is relatively small compared to satellite [315, 295] and torsion balance [316, 295] experiments, as these macroscopic experiments involve much large masses. Although Fig. 8(c) and (d) show slight excess in the performance of the clock comparisons at low masses, the final result of the MICROSCOPE experiment [314], which improved by a factor of 4.6 compared to the initial result shown in the plot, is not yet included. With this updated result, most clock comparison constraints become marginal. Note that the constraint by Yb does not follow the formalism discussed in the previous subsection but arises from modulation in nuclear charge radii [312].
Constraints on quadratic coupling have been also established through both experimental and theoretical analyses. Early constraints were provided by theoretical studies, where analyses based on Big Bang Nucleosynthesis (BBN) and cosmic microwave background imposed limits several orders of magnitude more stringent than fifth-force searches [304]. Experimental constraints from atomic spectroscopy data include measurements from Dy [304], Rb/Cs clock comparison [303], Sr clock comparisons between the ground and an observation deck [317], and Yb+/Cs clock comparisons [311]. While constraints from BBN are quite stringent, clock comparisons contribute significantly in the low-mass region for .
VII.3 Future proposals
To increase the search region, three factors are important. A longer measurement time allows for the detection of smaller masses. In principle, ultralight dark matter with arbitrarily small masses can be searched given a sufficiently long measurement period. However, for masses below eV, careful interpretation is required, as discussed in Sec. VII.1. Stability of clocks is another crucial factor. To investigate smaller coupling constants, the clocks need to be more stable, assuming the same atomic clocks are used. The sensitivity of a transition to the fundamental constant also plays a significant role. Even with the same clock stability, transitions highly sensitive to variations of fundamental constants enhance the overall sensitivity to ultralight dark matter. For example, in the case of , transitions with large , as listed in Table 2, are particularly suitable for high-sensitivity measurements.
Various systems are proposed to be sensitive to searches for ultralight dark matter. One approach is molecular spectroscopy. Microwave spectroscopy of a rovibrational transition in SrOH has a strong sensitivity to and [318]. Spectroscopy of Ca+ ion and I ion is expected to provide a comparable constraint on [319]. Another proposal involves placing an atomic clock on a satellite orbiting close to the sun [320]. The higher density of dark matter bound to the sun enhances sensitivity, particularly improving constraints in the higher mass range, from to eV, beyond existing limits. Other molecular and atomic spectroscopy studies have also placed constraints in the higher mass region [321, 322]. Atom interferometry can contribute within the same mass range as atomic clock comparisons [323]. The 229Th nuclear clock is particularly sensitive to ultralight dark matter, with Ref. [324] discussing its sensitivity . In addition, the large shown in Table 2 results in high sensitivity to [325].
VIII Fifth force search with isotope shifts
VIII.1 Theoretical basis
Theoretical proposal for searching for new forces using isotope shift measurements emerged around 2018 [326, 327]. Around the same time, studies on Higgs coupling to the electron and the up and down quarks [328], as well as other new physics searches [329] were also proposed. Since then, various experimental reports have been published, ranging from simple isotope shift measurements to comprehensive analyses incorporating new isotope shift data, motivated by these theoretical proposals.
Isotope shifts refer to the frequency difference of an atomic transition between isotopes. This phenomenon was already known in the first half of the 20th century [330, 331]. For optical transitions with THz, isotope shifts typically range from 10 MHz to 1 GHz. These shifts arise mainly from two sources and can be phenomenologically described as:
The first term, referred to as the mass shift, results from differences in nuclear mass. The second term, known as the field shift, is caused by variation in the nuclear charge radius due to differences in the number of neutrons. In Eq. 10 and the following discussion, the subscript labels a transition, and , unless otherwise specified. For example, is the transition frequency for the transition of the isotope with mass number . is the inverse mass of the isotope with mass number , is the th moment of the charge radius for the isotopes with mass number , and and are constants characterizing the mass shift and field shift, respectively. can be eliminated two isotope shift equations for transitions and .
Here, . The equation shows that is a linear function of . The two-dimensional plot of versus is called the King plot [332].
Suppose an additional term in Eq. 10. This term introduces on the right-hand side of Eq. 11, causing the King plot to become nonlinear. One possible source of this additional term is a hypothetical force between an electron and a neutron, generating the following Yukawa-type potential energy for an electron:
where is the spin for the new boson mediating the force, and are the coupling constants of the boson to an electron and a neutron, respectively, and is the mass of the new boson. Nonlinearity in a King plot can be attributed to this new force. Conversely, if the plot remains linear, it provides evidence against the existence of such a force.
Higher-order effects of the nuclear charge radius, which is within the SM, also induce nonlinearity. The following equation is typically assumed in the state-of-the-art analysis.
The third and fourth terms correspond to the 4th-moment shift and the quadratic field shift, respectively, where is the second order perturbative effect due to . The equation analogous to Eq. 11 is
In this equation, the third, fourth, and fifth terms all contribute to nonlinearity. As a result, properly subtracting background effects from the SM contributions are crucial when searching for new forces.
Even with effects that induce nonlinearity, the magnitude of this nonlinearity remained smaller than the uncertainty of the isotope shift measurements when the broad-linewidth transitions were used. This was experimentally tested decades ago[333]. Moreover, Eq. 10, which predicts linearity in the King plot, is widely used to determine differences in nuclear root-mean-square (RMS) charge radii [334, 335], for heavy atoms by renormalizing all field shifts into a single term by a proportionality coefficient common to all isotopes. One exception is the nonlinearity observed in Sm [336], which was explained by the field shift [337]. With the advancement in precision spectroscopy, as shown in Fig. 2, isotope shift measurements are now sensitive enough to detect nonlinearity beyond the measurement uncertainty caused by the natural linewidth of transitions and other systematic broadening.
Several conditions are necessary for high-sensitivity measurements. Searching for new forces through King plot nonlinearity requires an element with at least four stable isotopes. The more isotopes and transitions available, the more information can be extracted. Narrow-linewidth transitions are preferred for high-precision measurements. In addition, spinless nuclei are favored to avoid spin-spin interactions that are difficult to precisely estimate. In fact, significant nonlinearity is observed when isotopes with nonzero spin are included in the King plot for Sr [338] and Yb [122], with the second-order hyperfine interaction fully explaining the observed nonlinearity in Sr [339]. For higher sensitivity, transitions involving excited states with large difference in radial wave functions are preferable. For example, a combination of and transitions are expected to yield higher sensitivity in the King plot compared to a combination of and the transitions. To suppress unwanted backgrounds arising from SM effects, light nuclei are preferred, because of their small charge radii and reduced relativistic effects. One key source of uncertainty is , which can be eliminated by using three narrow-linewidth transitions [340]. Another approach to analyzing isotope shift data is to sort them by isotopes rather than transitions[341], which claimed to provide constraints on the new bosons comparable to those obtained through conventional King plot analysis.
This new force coupling to an electron and a neutron is introduced phenomenologically, without necessarily being supported by strong theoretical motivations. However, one benefit of this approach is its sensitivity to the X17 particle [342]. The force discussed in this section is spin-independent, while spin-dependent fifth force searches are covered in a recent review paper [13]. Also, the formalism and recent progresses in the King plot analysis is also summarized in another recent review paper[343].
VIII.2 Experimental results
Since the publications of theory papers around 2018, a large number of experimental results has been reported on this topic, ranging from isotope shift measurements alone to direct constraints on the existence of new forces. For certain atomic species, such as Xe [254], Cd [358], and Nd [359], isotope shifts were reported in the context of the search for new bosons. The constraints on the existence of new bosons are summarized in Fig. 9. The first experimental constraints on the new bosons were established through isotope shift measurements for Yb+ [348] and Ca+ [344]. In these studies, isotope shift measurements of Ca+ with a 20-Hz uncertainty excluded the existence of new bosons based on the observed linearity of the King plot, whereas nonlinearity is observed in the isotope shift measurements of Yb+ with a 300-Hz uncertainty. The initial Yb+ study focused on two E2 transitions, and later measurements included isotope shifts for the E3 transition with an accuracy of 500 Hz[346]. Additionally, multiple transitions in neutral Yb were carefully analyzed [347, 349, 122]. The most recent results for the E2 and E3 transitions of Yb+ achieved uncertainties of 5 and 16 Hz, respectively [350]. While Ca+ and some Yb measurements did not observe nonlinearity, consistent with constraint based on the measurement for electrons and neutron scattering, certain Yb dataset exhibit nonlinearity , leading to allowed regions for the new bosons. These allowed and excluded regions contradict each other, suggesting that the observed nonlinearity may be induced by some background effects within the SM. This calls for further investigation. A recent observation of nonlinearity in Ca [255] was based on the combination of the precision spectroscopy of Ca+ and Ca14+. This study attributed the observed nonlinearity to the second order mass shift and nuclear polarization, setting most stringent constraint on the new bosons.
If the effect of nuclear charge radii does not have to be eliminated using isotope shift data from another transition, the effect of the new boson can be probed through the spectroscopy of a single transition. Such a constraint was obtained from spectroscopy of transition in hydrogen and deuterium[357], as shown in Fig. 9. In this case, nuclear charge radii information is derived from spectroscopy of muonic atoms.
When nonlinearity is observed in the King plot, its source must be carefully analyzed. For this purpose, nonlinearity is decomposed into the following two components: , where is the residual of four data points from a linear fit [348] (Note that in Ref. [346] is equivalent to ). On the plane, a single source of nonlinearity with varying strength forms a line passing through the origin. When residuals of King plots for different combination of transitions in Yb are plotted on this plane, the data points approximately align along a single line. This suggests that the observed nonlinearity primarily originates from a single source, which can be explained by nuclear density functional theory calculations. These calculations incorporate several factors, including the surface thickness of nuclear density, relativistic corrections, nuclear deformation, and pairing effects [346]. The isotope shift measurements are so precise that they can select a nuclear model that accurately describes the experimental results. [346, 350].
After accounting for the influence of the nuclear charge distribution, any remaining nonlinearity still can be carefully examined. If this residual nonlinearity is attributed to new bosons, constraints on the existence of new bosons mediating forces between a neutron and an electron can be derived from isotope shift measurements of Yb. Some analyses of currently available data show statistically significant nonlinearity, while others show linear behavior. As a result, different combinations of transitions produce both allowed and excluded regions, as shown in Fig. 9.
The rapid advancement of King plot analysis has driven the need for precise mass measurements [360, 361]. Because the inverse mass difference plays an important role, highly accurate mass measurements of different isotopes are essential to ensure that mass uncertainties do not limit the search for new bosons. Following the first report on Yb [348], the mass uncertainty of 168Yb was improved [362], and recent report on Yb includes even more precise mass measurements [350]. For isotope shift measurements with even smaller uncertainties, further improvements in mass measurements are necessary, unless generalized King plot is implemented [340]. To enhance the sensitivity of these searches, the use of radioactive isotopes [338, 340] are proposed.
IX Atomic parity violation
IX.1 Theoretical basis
Parity (P) transformation refers to flipping the sign of one spatial coordinate. In three dimensions, this is equivalent to the transformation . Under this transformation, fundamental equations of motion, such as the Newton’s equation of motion, the Schrödinger equation, and the Dirac equation, remain unchanged. This invariance under P transformation is known as P symmetry. However, among the four fundamental forces, P symmetry is completely violated in the weak force. Historically, possibility of P violation was proposed by T.D. Lee and C. N. Yang [363]. Shortly thereafter, C. S. Wu experimentally confirmed it [364]. The discovery was later incorporated into the SM as a part of the weak interaction theory.
In atoms, where the electromagnetic force dominates, P symmetry is naively conserved. If the P symmetry is violated in interactions in atoms, it is induced by the weak force or BSM physics. Examples of such interactions are summarized in the Feynman diagrams in Fig. 10. Fig.10(a) shows an interaction between an electron and a nucleus mediated by a Z boson instead of a photon. Although this interaction occurs at tree-level, its effect is very small due to the large mass of the Z boson. Fig.10(b) and (c) depict higher-order perturbative processes, where an additional Z boson mediating the force, or a virtual W or Z boson forms a loop within the nucleus. If a hypothetical boson appears in a loop or mediates the force between the nucleus and the electron, this interaction would indicate the presence of BSM physics[365, 366]. The effective Hamiltonian for spin-independent APV is given by
where is the Fermi constant for the weak force, with being the Dirac matrices, and () is the creation (annihilation) operator for an electron. The effective weak charge consists of the SM contribution and the BSM contribution .
with and being number of protons and neutrons, respectively, being the Weinberg angle, and being the BSM contribution to the vector current, where are the coupling of the BSM physics to up and down quarks, respectively. This equation indicates that APV experiments effectively measure . The energy scale of this interaction corresponds to the size of the hydrogen-like 1S orbital for a nucleus with charge , i.e., the innermost electronic orbital. This leads to a momentum transfer of [365]. Intuitively, this occurs because electrons experience the weak force primarily in regions where the Coulomb potential from the nucleus is unscreened. This momentum transfer is significantly lower than that in other measurements of performed in scattering experiments. Searches for deviations of the from serves as a probe for P-violating BSM physics, e.g., Z’ boson and leptoquarks [365, 367, 368].
Spin-dependent APV originates from all three Feynman diagrams shown in Fig. 10. Diagram (c) illustrates the contribution of the nuclear anapole moment, which is analogous to a solenoidal current bent into a toroidal shape. The anapole moment is proportional to nuclear spin , and serves as a probe for nuclear parity violation. Some constraints on the anapole moment were set through the Cs experiment discussed in the following section.
The effect of P violation can be observed as a nonzero transition amplitude for an electronic transition between states of the same parity. P violation causes mixing between states of opposite parity, which slightly allows transitions between the same parity states, something normally forbidden in E1 transitions. In practice, measuring APV involves evaluating the contribution of the P-violating term in the total matrix element of the otherwise forbidden transition between electronic orbitals of the same parity.
Because the effect of APV is extremely small, various enhancement techniques are used to amplify the signal to an observable level. The mixing parameter between an state and a state due to APV satisfies the following equation.
where is the transition energy for the transition of interest. Due to the scaling [369, 370], heavy atoms are preferred for APV measurements. The measured quantity is the interference term between the P-conserving amplitude and the P-violating amplitude :
where is the dc electric field applied to mix parity-odd and parity-even states, is the tensor transition polarizability, is the handedness of the polarization of the oscillating field, is the magnitude of the oscillating field perpendicular to , and are Clebsch-Gordan coefficients, and is the E1 transition matrix element due to APV [371]. is extracted from experimental measurements and converted to and .
IX.2 Experimental reports
The two-level system suitable for APV consists of the ground state and an excited state with the same parity as the ground state. Due to the enhancement, heavy atoms are preferred, and historically measurements were mainly performed using Cs atoms, as shown in Table 7. In these experiments, the 6S ground state and the 7S excited state are regarded as the two-level system. The E1 transition between these two state is forbidden, but they can be coupled via an M1 transition. In the experiment, a standing wave generated by a build-up cavity helps suppress the M1 transition. To drive the 6S-7S transition in a controlled manner, the 7P state is mixed with the 7S state by a static electric field ( in Eq. IX.1). By flipping the sign of and , additional discrimination between the backgrounds and the signal is achieved. After accounting for all systematic uncertainties, most of which are related to the M1 amplitude , is obtained[372], which leads to around 2.4 MeV.
Another attempt to observe APV is performed using Yb. In this case, the two-level system consists of the ground state and the state. A significant parity-violating effect is observed [373] for this transition. One advantage of Yb over Cs is the number of stable isotopes. While Cs has only one, Yb has five spinless isotopes. This allows for the estimation of neutron contribution to by measuring the P-violating effect across multiple isotopes [374].
The next generation experiment aims to measure the APV in Fr. Because all Fr isotopes are unstable, these experiments are conducted at accelerator facilities. Fr atoms generated by nuclear reactions were successfully loaded into a magneto-optical trap in the 1990s at Stony Brook [375]. Since then, relevant transitions have been characterized in preparation for APV measurement[376, 377]. An effort to improve the accuracy of APV measurements is Cs is also ongoing [378].
Another method to enhance sensitivity is to use a small . To achieve this, the use of diatomic molecules [379] and dysprosium [380] is proposed. In these species, closely spaced states with the opposite parity enhances the mixing shown in Eq. IX.1. Experimental sensitivity has been reported for a BaF molecule experiment [381]. In the case of BaF, the measured quantity is the nuclear spin dependent APV. A MOT for BaF is recently realized [382].
X Electric dipole moment search
X.1 Theoretical basis
Electric dipole moments (EDMs) of elementary particles serve as a probe for high-energy physics through the violation of time-reversal (T) symmetry. If a spin-1/2 elementary particle has a nonzero EDM, it violates T symmetry because the EDM remains unchanged while the magnetic dipole moment flips under the T operation. Assuming the CPT theorem [386, 387] (C stands for the charge conjugation), this implies a violation of the symmetry under the CP operation. Such a CP-violating moment is predicted in the SM of particle physics. For the electron, four-loop perturbations predict an EDM of approximately cm [388]. denotes the EDM of , where is for electron or some nuclei such as Yb, Hg, and Ra. CP-violating interactions between electrons and nucleus mimics an electron EDM. This effect is estimated to range from cm [389] to cm [388]. In supersymmetric theories, CP-violating moments can arise from one-loop diagrams, resulting in significantly larger EDM values. This argument extends to other elementary particles and nuclei, making EDM searches an important avenue to test supersymmetric theories. Also, CP-violation is closely linked to the matter-antimatter asymmetry in the universe. Moreover, if an axion-like particle contributes via loop diagrams at a vertex, EDM searches can put constraints on such particles [390, 391].
The search for EDM began with a theoretical proposal in 1950 [392]. Since then, experiments have been conducted for various particles, including the neutron, electron, muon, and certain nuclei, as summarized in a separate review paper[393]. Here, since AMO physics is focused on, only searches for electron EDM and some nuclear EDM are discussed.
Schiff’s theorem [394] plays an important role in EDM searches. It states that the overall EDM of an atom vanishes if it consists of point-like nonrelativistic charged particles interacting with a static electric field. However, a measurable EDM can arise due to relativistic effects, the finite size of the nucleus, and the presence of an unpaired electron. As a result, searches for the electron EDM are always performed using atoms or molecules with an unpaired electron. The significance of relativistic effects and enhancement of the electron EDM in an atom, as discussed in Ref. [395], forms the foundation for EDM searches using atoms and molecules.
In experiments for EDM searches, the small effect of the EDM is investigated using a Ramsey sequence under both a magnetic field and an electric field . If a particle only has a magnetic moment due to its spin, it precesses according to the Hamiltonian . If the particle possesses a nonzero EDM , the amount of precession changes according to the modified Hamiltonian .
X.2 Electron EDM
The electron is one of the elementary particles for which EDM searches have long been performed, as shown in Fig. 11. Because it is a lepton, BSM physics contributions at single-loop level arises from new particles with leptonic quantum number, e.g., the supersymmetric partner of the electron [396]. The electron EDM was first investigated indirectly by examining its effect on the Lamb shift in hydrogen atoms [397]. Although the first direct search was performed with bare electrons [398], modern experiments primarily use neutral atoms and molecules. This approach allows for the application of stronger effective electric fields, enhancing the EDM-induced precession through internal electric fields.
For an atom with atomic number , the enhancement factor of the electron EDM is [423]. Because of this dependence, heavy atoms are typically used in EDM searches. Most of the measurements with atoms shown in Fig. 11 were performed with Cs and Tl.
To achieve even stronger internal electric fields, polar molecules are utilized. Molecules composed of a soft, heavy atom and a highly electronegative atom can generate an internal electric field on the order of 10 GV/cm [424, 425, 426]. Molecules are selected such that they have one additional electron beyond fully closed inner shells and valence electrons involved in chemical bonding. This extra electron serves as the probe for EDM measurements. The YbF experiment [416] pioneered the use of molecules in EDM searches, and recent progresses are made with ThO [421] and HfF+ molecules [422].
In these molecules, a set of states known as -doublet states is used for EDM searches. is the projection of the electron’s total angular momentum on the internuclear axis of the diatomic molecule. For the case of , states are nearly degenerated, forming two closely spaced states with parity and [427]. The -doublet can be fully polarized by applying a small external electric field. When polarized, a pair of states at the same energy exhibits the same orientation but opposite spin orientation. The Ramsey sequence is performed under external magnetic and electric fields. The measured signal corresponds to the precession of the state relative to the state. To cancel systematic shifts and extract the EDM signal, the experiment involves flipping external electric and magnetic fields, as well as the orientation of the molecule. The difference in precession across all combinations of these flips are systematically processed to extract the desired EDM signal.
The figure of merit for the sensitivity follows the same SQL as atomic clocks [428]:
where is the number of molecules including detection efficiency. The choice of molecule is based on and the coherence time of the -doublet states, which limits . For ThO, the coherence time is 4.2(5) ms [429]. Given this timescale, a beam experiment is suitable, and measurements are conducted using buffer gas-cooled cryogenic beam[430]. It is relatively easy to increase the number of molecules to enhance the sensitivity by reduced QPN. In contrast, the -doublet states in HfF+ ion have significantly longer lifetime, allowing an interrogation time of up to 3 s [422]. Although the number of trapped ions is much smaller compared to molecules in a beam, this extended interrogation time compensates for this limitation. As a result, the sensitivity of these two state-of-the-art experiments are different by only a small factor.
Once a sufficient level of sensitivity is achieved, all systematic shifts must be accurately estimated and compensated to extract the electron EDM with small uncertainty. Two major sources of uncertainty are nonuniform magnetic fields and non-inverting components of the applied electric and magnetic fields. Another source of uncertainty arises from laser power stability, including variations in the overlap between the light and the molecules. Additionally, in the HfF+ experiment, where ions undergo circular motion, uncertainties specific to this motion also needs to be carefully considered.
The most stringent constraint on the electron EDM comes from the HfF+ experiment, which sets an upper limit of cm [422]. The ThO experiment provides the constraint of cm [421]. When properly interpreted within the framework of supersymmetric theories, these constraints correspond to TeV-scale physics[396]. Previous experiments have already excluded the naive supersymmetric theory, and the latest results constrain alternative theoretical frameworks. The highest attainable energy scale accessible through these constraints is 40 TeV [422].
Upgrades to the ThO experiment [431] and investigation into a successor experiment for the HfF+ experiment[432], as well as for improvements for the YbF experiment [433] are currently underway. At the proposal stage, enhancing sensitivity using entangled states is discussed. One such proposal suggests integrating a cavity QED system with an electron EDM measurement using Fr to generate a squeezed-spin state for Fr [434]. Experimental efforts are also focused on future EDM searches with radioactive isotopes, particularly in molecules. RaF has gained a significant interest not only for electron EDM searches but also for APV studies[435]. Experimental investigations of molecular spectroscopy of RaF are ongoing [436]. In general, Ra-containing molecules attract considerable attention, with both theoretical[437, 438] and experimental[439] studies being conducted on various species. In addition, polyatomic molecules are being explored as promising next-generation candidates for electron EDM searches. A theoretical proposal [440] pointed out that certain polyatomic molecules have states easy to polarize, analogous to doublets, yet independent of electronic spins. This property prevents spin-related complications in laser cooling and sensitivity to the electron EDM. This proposal triggered investigations on YbOH [441] and YbOCH3 [442]. Recently, a MOT for SrOH was demonstrated [443] in the context of electron EDM searches. Furthermore, a theoretical proposal suggests that the sensitivity of EDM searches with polyatomic molecules could be enhanced through quantum entanglement [444].
X.3 Nuclear EDM
Within the technology of AMO physics, nuclear EDMs can also be measured, with current constraints reaching the 10-TeV scale physics. Nuclear EDM measurements are performed using atomic or molecular systems where electrons are fully paired to eliminate any effects from the electron EDM. Nuclei with spin 1/2 are preferred due to the simple structure and the absence of deformation effects. As a result, popular species for nuclear EDM searches are 129Xe, 171Yb, 199Hg, 225Ra, and the neutron. To measure a nuclear EDM, spin-polarized nuclei are used. One method to prepare and detect such a spin-polarized system is through optical pumping and optical detection with lasers. Among the species listed above, experiments with 171Yb, 199Hg, and 225Ra rely on such atomic physics techniques and are the focus here. For other systems, a detailed summary can be found in another review paper [393].
In the measurement sequence, nuclei are placed under magnetic and electric fields, and a Ramsey sequence is performed to detect the tiny phase difference when the direction of the electric field is reversed. The relevant energy level corresponds to the nuclear spin 1/2 in a strong magnetic field, meaning the energy scale for the Ramsey sequence is in the range of NMR spectroscopy, not laser spectroscopy. Because the internal electric field of the nucleus is zero, the applied electric field originates from electrodes, and is significantly smaller than in electron EDM searches. However, due to the exceptionally long coherence time of nuclear spin, the interrogation time can reach s, resulting in constraints of cm for 199Hg [445], cm [446] for 171Yb, and cm for 226Ra [447], which are comparable to the constraint on the electron EDM.
The Hg experiment uses a vapor cell, making the major systematic uncertainty related to atomic motion. Yb atoms are laser-cooled and trapped in an optical dipole trap, eliminating motion-related uncertainty. However, the optical dipole trap introduces other sources of uncertainty. The 225Ra experiment [447] has a similar configuration to the Yb experiment, with its source of uncertainty arising from a phase shift proportional to the square of the electric field and the magnetic field correlated to the electric field.
XI Summary and Outlook
High precision measurements with quantum states in atoms can detect small energy shifts induced by phenomena originating from nuclear and particle physics. For future advancements, several key aspects are important. One is improvement in measurement precision. Even within the same experimental system, increased measurement precision enhances sensitivity to nuclear and particle physics phenomena in most of the topics covered in this paper. In the past, improvements in clock uncertainty contributed to searches for ultralight dark matter and time variation of fundamental constants. Secondly, the use of highly sensitive quantum sensors can significantly improve the sensitivity. A notable example is searches for variation of the fine structure constant. Switching from conventional ion clocks to HCI clocks or nuclear clocks would enhance the sensitivity by orders of magnitude. In addition, a solid theoretical understanding of the measurements is crucial. Theoretical calculations of atomic or molecular structures determine the sensitivity of a quantum sensor to the quantity of interest. Examples include the sensitivity coefficient for the variation of the fine structure constant and the internal electric field for the electron EDM search. Further improvements in the accuracy of these theoretical calculations can refine the determination of parameters in nuclear and particle physics. Moreover, new theoretical proposals can inspire extensive experimental searches for previously unexplored parameter spaces. Historically, searches for the fifth force using isotope shifts and ultralight dark matter expanded following initial theoretical proposals. With the combination of these factors, further exploration of nuclear and particle physics using atoms and molecules as quantum sensors are expected.
Acknowledgements.
The author thanks Takashi Higuchi, Joonseok Hur, Yevgeny Stadnik, and Takumi Kobayashi for helpful discussions and comments on the manuscript. A.K. acknowledges the support of Japan Society for the Promotion of Science KAKENHI Grants No. JP22H01161, and Japan Science and Technology Agency FOREST Grant No. JPMJFR212S. Data sharing is not applicable to this article as no new data were created or analyzed in this study.References
- Degen, Reinhard, and Cappellaro [2017] C. L. Degen, F. Reinhard, and P. Cappellaro, “Quantum sensing,” Rev. Mod. Phys. 89, 035002 (2017).
- Kjaergaard et al. [2020] M. Kjaergaard, M. E. Schwartz, J. Braumüller, P. Krantz, J. I.-J. Wang, S. Gustavsson, and W. D. Oliver, “Superconducting qubits: Current state of play,” Annu. Rev. Condens. Matter Phys. 11, 369–395 (2020).
- Krantz et al. [2019] P. Krantz, M. Kjaergaard, F. Yan, T. P. Orlando, S. Gustavsson, and W. D. Oliver, “A quantum engineer’s guide to superconducting qubits,” Appl. Phys. Rev. 6, 021318 (2019).
- Barzanjeh et al. [2022] S. Barzanjeh, A. Xuereb, S. Gröblacher, M. Paternostro, C. A. Regal, and E. M. Weig, “Optomechanics for quantum technologies,” Nat. Phys. 18, 15–24 (2022).
- Millen et al. [2020] J. Millen, T. S. Monteiro, R. Pettit, and A. N. Vamivakas, “Optomechanics with levitated particles,” Rep. Prog. Phys. 83, 026401 (2020).
- Schirhagl et al. [2014] R. Schirhagl, K. Chang, M. Loretz, and C. L. Degen, “Nitrogen-vacancy centers in diamond: Nanoscale sensors for physics and biology,” Annu. Rev. Phys. Chem. 65, 83–105 (2014).
- Doherty et al. [2013] M. W. Doherty, N. B. Manson, P. Delaney, F. Jelezko, J. Wrachtrup, and L. C. Hollenberg, “The nitrogen-vacancy colour centre in diamond,” Phys. Rep. 528, 1–45 (2013).
- Barry et al. [2020] J. F. Barry, J. M. Schloss, E. Bauch, M. J. Turner, C. A. Hart, L. M. Pham, and R. L. Walsworth, “Sensitivity optimization for NV-diamond magnetometry,” Rev. Mod. Phys. 92, 015004 (2020).
- Gross and Bloch [2017] C. Gross and I. Bloch, “Quantum simulations with ultracold atoms in optical lattices,” Science 357, 995–1001 (2017).
- Bohn, Rey, and Ye [2017] J. L. Bohn, A. M. Rey, and J. Ye, “Cold molecules: Progress in quantum engineering of chemistry and quantum matter,” Science 357, 1002–1010 (2017).
- Bruzewicz et al. [2019] C. D. Bruzewicz, J. Chiaverini, R. McConnell, and J. M. Sage, “Trapped-ion quantum computing: Progress and challenges,” Appl. Phys. Rev. 6, 021314 (2019).
- Monroe and Kim [2013] C. Monroe and J. Kim, “Scaling the ion trap quantum processor,” Science 339, 1164–1169 (2013).
- Cong et al. [2024] L. Cong, W. Ji, P. Fadeev, F. Ficek, M. Jiang, V. V. Flambaum, H. Guan, D. F. J. Kimball, M. G. Kozlov, Y. V. Stadnik, et al., “Spin-dependent exotic interactions,” (2024), arXiv:2408.15691 [hep-ph] .
- Graham et al. [2015] P. W. Graham, I. G. Irastorza, S. K. Lamoreaux, A. Lindner, and K. A. van Bibber, “Experimental searches for the axion and axion-like particles,” Ann. Rev. of Nucl. Part. Sci. 65, 485–514 (2015).
- Choi, Im, and Shin [2021] K. Choi, S. H. Im, and C. S. Shin, “Recent progress in the physics of axions and axion-like particles,” Ann. Rev. of Nucl. Part. Sci. 71, 225–252 (2021).
- Di Luzio et al. [2020] L. Di Luzio, M. Giannotti, E. Nardi, and L. Visinelli, “The landscape of qcd axion models,” Phys. Rep. 870, 1–117 (2020), the landscape of QCD axion models.
- Berengut et al. [2012a] J. C. Berengut, V. A. Dzuba, V. V. Flambaum, and A. Ong, “Highly charged ions with E1, M1, and E2 transitions within laser range,” Phys. Rev. A 86, 022517 (2012a).
- Peik and Tamm [2003] E. Peik and C. Tamm, “Nuclear laser spectroscopy of the 3.5 eV transition in Th-229,” Europhys. Lett. 61, 181 (2003).
- Phillips and Metcalf [1982] W. D. Phillips and H. Metcalf, “Laser deceleration of an atomic beam,” Phys. Rev. Lett. 48, 596–599 (1982).
- Raab et al. [1987] E. L. Raab, M. Prentiss, A. Cable, S. Chu, and D. E. Pritchard, “Trapping of neutral sodium atoms with radiation pressure,” Phys. Rev. Lett. 59, 2631–2634 (1987).
- Dalibard and Cohen-Tannoudji [1989] J. Dalibard and C. Cohen-Tannoudji, “Laser cooling below the doppler limit by polarization gradients: simple theoretical models,” J. Opt. Soc. Am. B 6, 2023–2045 (1989).
- Grimm, Weidemüller, and Ovchinnikov [2000] R. Grimm, M. Weidemüller, and Y. B. Ovchinnikov, “Optical dipole traps for neutral atoms,” (Academic Press, 2000) pp. 95–170.
- Ketterle and Druten [1996] W. Ketterle and N. V. Druten, “Evaporative cooling of trapped atoms,” (Academic Press, 1996) pp. 181–236.
- Monroe et al. [1995] C. Monroe, D. M. Meekhof, B. E. King, S. R. Jefferts, W. M. Itano, D. J. Wineland, and P. Gould, “Resolved-sideband Raman cooling of a bound atom to the 3D zero-point energy,” Phys. Rev. Lett. 75, 4011–4014 (1995).
- Hamann et al. [1998] S. E. Hamann, D. L. Haycock, G. Klose, P. H. Pax, I. H. Deutsch, and P. S. Jessen, “Resolved-sideband Raman cooling to the ground state of an optical lattice,” Phys. Rev. Lett. 80, 4149–4152 (1998).
- Zhang et al. [2022] X. Zhang, K. Beloy, Y. S. Hassan, W. F. McGrew, C.-C. Chen, J. L. Siegel, T. Grogan, and A. D. Ludlow, “Subrecoil clock-transition laser cooling enabling shallow optical lattice clocks,” Phys. Rev. Lett. 129, 113202 (2022).
- Chen et al. [2024a] C.-C. Chen, J. L. Siegel, B. D. Hunt, T. Grogan, Y. S. Hassan, K. Beloy, K. Gibble, R. C. Brown, and A. D. Ludlow, “Clock-line-mediated Sisyphus cooling,” Phys. Rev. Lett. 133, 053401 (2024a).
- Hawkins and Dicke [1953] W. B. Hawkins and R. H. Dicke, “The polarization of sodium atoms,” Phys. Rev. 91, 1008–1009 (1953).
- Happer [1972] W. Happer, “Optical pumping,” Rev. Mod. Phys. 44, 169–249 (1972).
- Vitanov et al. [2017] N. V. Vitanov, A. A. Rangelov, B. W. Shore, and K. Bergmann, “Stimulated Raman adiabatic passage in physics, chemistry, and beyond,” Rev. Mod. Phys. 89, 015006 (2017).
- Gaubatz et al. [1990] U. Gaubatz, P. Rudecki, S. Schiemann, and K. Bergmann, “Population transfer between molecular vibrational levels by stimulated Raman scattering with partially overlapping laser fields. A new concept and experimental results,” J. Chem. Phys. 92, 5363–5376 (1990).
- Treacy and Demaria [1969] E. Treacy and A. Demaria, “Adiabatic inversion in the infrared,” Phys. Lett. A 29, 369–370 (1969).
- Loy [1974] M. M. T. Loy, “Observation of population inversion by optical adiabatic rapid passage,” Phys. Rev. Lett. 32, 814–817 (1974).
- B.J. Dalton and Knight [1985] R. M. B.J. Dalton and P. Knight, “Coherent population trapping,” Optica Acta: Int. J. Optics 32, 61–70 (1985).
- Cohen-Tannoudji, Dupont-Roc, and Grynberg [1998] C. Cohen-Tannoudji, J. Dupont-Roc, and G. Grynberg, Atom—Photon Interactions (Wiley, 1998).
- Levitt [1982] M. H. Levitt, “Symmetrical composite pulse sequences for nmr population inversion. i. compensation of radiofrequency field inhomogeneity,” J. Magn. Reson. 48, 234–264 (1982).
- Vandersypen and Chuang [2005] L. M. K. Vandersypen and I. L. Chuang, “Nmr techniques for quantum control and computation,” Rev. Mod. Phys. 76, 1037–1069 (2005).
- Essen and Parry [1957] L. Essen and J. V. L. Parry, “The caesium resonator as a standard of frequency and time,” Phil. Trans. R. Soc. A 250, 45–69 (1957).
- Beehle et al. [1962] R. E. Beehle, W. R. Atkinson, L. E. Heim, and C. S. Snider, “A comparison of direct and servo methods for utilizing cesium beam resonators as frequency standards,” IRE Trans. Instrum. I-11, 231–238 (1962).
- Beehler and Glaze [1966] R. E. Beehler and D. J. Glaze, “The performance and capability of cesium beam frequency standards at the National Bureau of Standards,” IEEE Trans. Instrum. Meas. 15, 48–55 (1966).
- Glaze [1970] D. J. Glaze, “Improvements in atomic cesium beam frequency standards at the National Bureau of Standards,” IEEE Trans. Instrum. Meas. 19, 156–160 (1970).
- Glaze et al. [1974] D. J. Glaze, H. Hellwig, D. W. Allan, S. Jarvis, and A. E. Wainwright, “Accuracy evaluation and stability of the NBS primary frequency standards,” IEEE Trans. Instrum. Meas. 23, 489–501 (1974).
- Taylor and Phillips [1984] B. N. Taylor and W. D. Phillips, “Precision measurement and fundamental constants II,” (1984).
- Drullinger et al. [1993] R. Drullinger, J. Lowe, D. Glaze, and J. Shirley, “NIST-7, the new US primary frequency standard,” in 1993 IEEE International Frequency Control Symposium (1993) pp. 71–74.
- Drullinger, Shirley, and Lee [1996] R. E. Drullinger, J. H. Shirley, and W. D. Lee, “NIST-7, the U.S. primary frequency standard: New evaluation techniques,” Proceedings of the 28th Annual Precise Time and Time Interval Systems and Applications Meeting , 255 (1996).
- Shirley, Lee, and Drullinger [2001] J. H. Shirley, W. D. Lee, and R. E. Drullinger, “Accuracy evaluation of the primary frequency standard NIST-7,” Metrologia 38, 427 (2001).
- Jefferts et al. [2002] S. R. Jefferts, J. Shirley, T. E. Parker, T. P. Heavner, D. M. Meekhof, C. Nelson, F. Levi, G. Costanzo, A. D. Marchi, R. Drullinger, et al., “Accuracy evaluation of NIST-F1,” Metrologia 39, 321 (2002).
- Heavner et al. [2005] T. P. Heavner, S. R. Jefferts, E. A. Donley, J. H. Shirley, and T. E. Parker, “NIST-F1: recent improvements and accuracy evaluations,” Metrologia 42, 411 (2005).
- Heavner et al. [2014] T. P. Heavner, E. A. Donley, F. Levi, G. Costanzo, T. E. Parker, J. H. Shirley, N. Ashby, S. Barlow, and S. R. Jefferts, “First accuracy evaluation of NIST-F2,” Metrologia 51, 174–182 (2014).
- Weyers et al. [2018] S. Weyers, V. Gerginov, M. Kazda, J. Rahm, B. Lipphardt, G. Dobrev, and K. Gibble, “Advances in the accuracy, stability, and reliability of the PTB primary fountain clocks,” Metrologia 55, 789 (2018).
- Weyers et al. [2001] S. Weyers, U. Hübner, R. Schröder, C. Tamm, and A. Bauch, “Uncertainty evaluation of the atomic caesium fountain CSF1 of the PTB,” Metrologia 38, 343 (2001).
- Bize et al. [2005] S. Bize, P. Laurent, M. Abgrall, H. Marion, I. Maksimovic, L. Cacciapuoti, J. Grünert, C. Vian, F. P. d. Santos, P. Rosenbusch, et al., “Cold atom clocks and applications,” J. Phys. B: At. Mol. Opt. Phys. 38, S449 (2005).
- Santarelli et al. [1999] G. Santarelli, P. Laurent, P. Lemonde, A. Clairon, A. G. Mann, S. Chang, A. N. Luiten, and C. Salomon, “Quantum projection noise in an atomic fountain: A high stability cesium frequency standard,” Phys. Rev. Lett. 82, 4619–4622 (1999).
- Bernard et al. [1999] J. E. Bernard, A. A. Madej, L. Marmet, B. G. Whitford, K. J. Siemsen, and S. Cundy, “Cs-based frequency measurement of a single, trapped ion transition in the visible region of the spectrum,” Phys. Rev. Lett. 82, 3228–3231 (1999).
- von Zanthier et al. [2000] J. von Zanthier, T. Becker, M. Eichenseer, A. Y. Nevsky, C. Schwedes, E. Peik, H. Walther, R. Holzwarth, J. Reichert, T. Udem, et al., “Absolute frequency measurement of the In+ clock transition with a mode-locked laser,” Opt. Lett. 25, 1729–1731 (2000).
- Udem et al. [2001] T. Udem, S. A. Diddams, K. R. Vogel, C. W. Oates, E. A. Curtis, W. D. Lee, W. M. Itano, R. E. Drullinger, J. C. Bergquist, and L. Hollberg, “Absolute frequency measurements of the Hg+ and Ca optical clock transitions with a femtosecond laser,” Phys. Rev. Lett. 86, 4996–4999 (2001).
- Stenger et al. [2001] J. Stenger, C. Tamm, N. Haverkamp, S. Weyers, and H. R. Telle, “Absolute frequency measurement of the 435.5-nm 171Yb+-clock transition with a kerr-lens mode-locked femtosecond laser,” Opt. Lett. 26, 1589–1591 (2001).
- Margolis et al. [2004] H. S. Margolis, G. P. Barwood, G. Huang, H. A. Klein, S. N. Lea, K. Szymaniec, and P. Gill, “Hertz-level measurement of the optical clock frequency in a single 88Sr+ ion,” Science 306, 1355–1358 (2004).
- Schneider, Peik, and Tamm [2005] T. Schneider, E. Peik, and C. Tamm, “Sub-hertz optical frequency comparisons between two trapped ions,” Phys. Rev. Lett. 94, 230801 (2005).
- Oskay et al. [2006] W. H. Oskay, S. A. Diddams, E. A. Donley, T. M. Fortier, T. P. Heavner, L. Hollberg, W. M. Itano, S. R. Jefferts, M. J. Delaney, K. Kim, et al., “Single-atom optical clock with high accuracy,” Phys. Rev. Lett. 97, 020801 (2006).
- Rosenband et al. [2008] T. Rosenband, D. B. Hume, P. O. Schmidt, C. W. Chou, A. Brusch, L. Lorini, W. H. Oskay, R. E. Drullinger, T. M. Fortier, J. E. Stalnaker, and ohters, “Frequency ratio of Al+ and Hg+ single-ion optical clocks; metrology at the 17th decimal place,” Science 319, 1808–1812 (2008).
- Chou et al. [2010] C. W. Chou, D. B. Hume, J. C. J. Koelemeij, D. J. Wineland, and T. Rosenband, “Frequency comparison of two high-accuracy Al+ optical clocks,” Phys. Rev. Lett. 104, 070802 (2010).
- Madej et al. [2012] A. A. Madej, P. Dubé, Z. Zhou, J. E. Bernard, and M. Gertsvolf, “ 445-THz single-ion reference at the level via control and cancellation of systematic uncertainties and its measurement against the si second,” Phys. Rev. Lett. 109, 203002 (2012).
- Huntemann et al. [2016] N. Huntemann, C. Sanner, B. Lipphardt, C. Tamm, and E. Peik, “Single-ion atomic clock with systematic uncertainty,” Phys. Rev. Lett. 116, 063001 (2016).
- Brewer et al. [2019a] S. M. Brewer, J.-S. Chen, A. M. Hankin, E. R. Clements, C. W. Chou, D. J. Wineland, D. B. Hume, and D. R. Leibrandt, “27Al+ quantum-logic clock with a systematic uncertainty below ,” Phys. Rev. Lett. 123, 033201 (2019a).
- Tofful et al. [2024] A. Tofful, C. F. A. Baynham, E. A. Curtis, A. O. Parsons, B. I. Robertson, M. Schioppo, J. Tunesi, H. S. Margolis, R. J. Hendricks, J. Whale, et al., “171Yb+ optical clock with systematic uncertainty and absolute frequency measurements,” Metrologia 61, 045001 (2024).
- Takamoto et al. [2005] M. Takamoto, F.-L. Hong, R. Higashi, and H. Katori, “An optical lattice clock,” Nature 435, 321–324 (2005).
- Le Targat et al. [2006] R. Le Targat, X. Baillard, M. Fouché, A. Brusch, O. Tcherbakoff, G. D. Rovera, and P. Lemonde, “Accurate optical lattice clock with atoms,” Phys. Rev. Lett. 97, 130801 (2006).
- Takamoto et al. [2006] M. Takamoto, F.-L. Hong, R. Higashi, Y. Fujii, M. Imae, and H. Katori, “Improved frequency measurement of a one-dimensional optical lattice clock with a spin-polarized fermionic 87Sr isotope,” J. Phys. Soc. Jpn. 75, 104302 (2006).
- Boyd et al. [2007] M. M. Boyd, A. D. Ludlow, S. Blatt, S. M. Foreman, T. Ido, T. Zelevinsky, and J. Ye, “87Sr lattice clock with inaccuracy below ,” Phys. Rev. Lett. 98, 083002 (2007).
- Ludlow et al. [2008] A. D. Ludlow, T. Zelevinsky, G. K. Campbell, S. Blatt, M. M. Boyd, M. H. G. de Miranda, M. J. Martin, J. W. Thomsen, S. M. Foreman, J. Ye, et al., “Sr lattice clock at fractional uncertainty by remote optical evaluation with a Ca clock,” Science 319, 1805–1808 (2008).
- Lemke et al. [2009] N. D. Lemke, A. D. Ludlow, Z. W. Barber, T. M. Fortier, S. A. Diddams, Y. Jiang, S. R. Jefferts, T. P. Heavner, T. E. Parker, and C. W. Oates, “Spin- optical lattice clock,” Phys. Rev. Lett. 103, 063001 (2009).
- Bloom et al. [2014] B. J. Bloom, T. L. Nicholson, J. R. Williams, S. L. Campbell, M. Bishof, X. Zhang, W. Zhang, S. L. Bromley, and J. Ye, “An optical lattice clock with accuracy and stability at the level,” Nature 506, 71–75 (2014).
- Ushijima et al. [2015] I. Ushijima, M. Takamoto, M. Das, T. Ohkubo, and H. Katori, “Cryogenic optical lattice clocks,” Nat. Photon. 9, 185–189 (2015).
- Nicholson et al. [2015] T. L. Nicholson, S. L. Campbell, R. B. Hutson, G. E. Marti, B. J. Bloom, R. L. McNally, W. Zhang, M. D. Barrett, M. S. Safronova, G. F. Strouse, et al., “Systematic evaluation of an atomic clock at total uncertainty,” Nat. Commun. 6, 6896 (2015).
- McGrew et al. [2018a] W. F. McGrew, X. Zhang, R. J. Fasano, S. A. Schäffer, K. Beloy, D. Nicolodi, R. C. Brown, N. Hinkley, G. Milani, M. Schioppo, et al., “Atomic clock performance enabling geodesy below the centimetre level,” Nature 564, 87–90 (2018a).
- Bothwell et al. [2019] T. Bothwell, D. Kedar, E. Oelker, J. M. Robinson, S. L. Bromley, W. L. Tew, J. Ye, and C. J. Kennedy, “JILA SrI optical lattice clock with uncertainty of ,” Metrologia 56, 065004 (2019).
- Aeppli et al. [2024] A. Aeppli, K. Kim, W. Warfield, M. S. Safronova, and J. Ye, “Clock with systematic uncertainty,” Phys. Rev. Lett. 133, 023401 (2024).
- Hausser et al. [2025] H. N. Hausser, J. Keller, T. Nordmann, N. M. Bhatt, J. Kiethe, H. Liu, I. M. Richter, M. von Boehn, J. Rahm, S. Weyers, et al., “115In+-172Yb+ Coulomb crystal clock with systematic uncertainty,” Phys. Rev. Lett. 134, 023201 (2025).
- Layer, Deslattes, and Schweitzer [1976] H. P. Layer, R. D. Deslattes, and W. G. Schweitzer, “Laser wavelength comparison by high resolution interferometry,” Appl. Opt. 15, 734–743 (1976).
- Jennings et al. [1983] D. A. Jennings, C. R. Pollock, F. R. Petersen, R. E. Drullinger, K. M. Evenson, J. S. Wells, J. L. Hall, and H. P. Layer, “Direct frequency measurement of the I2-stabilized He–Ne 473-THz (633-nm) laser,” Opt. Lett. 8, 136–138 (1983).
- Beausoleil et al. [1987] R. G. Beausoleil, D. H. McIntyre, C. J. Foot, E. A. Hildum, B. Couillaud, and T. W. Hänsch, “Continuous-wave measurement of the 1S Lamb shift in atomic hydrogen,” Phys. Rev. A 35, 4878–4881 (1987).
- McIntyre et al. [1989] D. H. McIntyre, R. G. Beausoleil, C. J. Foot, E. A. Hildum, B. Couillaud, and T. W. Hänsch, “Continuous-wave measurement of the hydrogen 1S-2S transition frequency,” Phys. Rev. A 39, 4591–4598 (1989).
- Andreae et al. [1992] T. Andreae, W. König, R. Wynands, D. Leibfried, F. Schmidt-Kaler, C. Zimmermann, D. Meschede, and T. W. Hänsch, “Absolute frequency measurement of the hydrogen 1S-2S transition and a new value of the rydberg constant,” Phys. Rev. Lett. 69, 1923–1926 (1992).
- Acef et al. [1993] O. Acef, J. Zondy, M. Abed, D. Rovera, A. Gérard, A. Clairon, P. Laurent, Y. Millerioux, and P. Juncar, “A CO2 to visible optical frequency synthesis chain: accurate measurement of the 473 THz HeNe/I2 laser,” Opt. Commun. 97, 29–34 (1993).
- Schnatz et al. [1996] H. Schnatz, B. Lipphardt, J. Helmcke, F. Riehle, and G. Zinner, “First phase-coherent frequency measurement of visible radiation,” Phys. Rev. Lett. 76, 18–21 (1996).
- Udem et al. [1997] T. Udem, A. Huber, B. Gross, J. Reichert, M. Prevedelli, M. Weitz, and T. W. Hänsch, “Phase-coherent measurement of the hydrogen transition frequency with an optical frequency interval divider chain,” Phys. Rev. Lett. 79, 2646–2649 (1997).
- Riehle et al. [1999] F. Riehle, H. Schnatz, B. Lipphardt, G. Zinner, T. Trebst, and J. Helmcke, “The optical calcium frequency standard,” IEEE Trans. Instrum. Meas. 48, 613–617 (1999).
- Niering et al. [2000] M. Niering, R. Holzwarth, J. Reichert, P. Pokasov, T. Udem, M. Weitz, T. W. Hänsch, P. Lemonde, G. Santarelli, M. Abgrall, et al., “Measurement of the hydrogen - transition frequency by phase coherent comparison with a microwave cesium fountain clock,” Phys. Rev. Lett. 84, 5496–5499 (2000).
- Fischer et al. [2004] M. Fischer, N. Kolachevsky, M. Zimmermann, R. Holzwarth, T. Udem, T. W. Hänsch, M. Abgrall, J. Grünert, I. Maksimovic, S. Bize, et al., “New limits on the drift of fundamental constants from laboratory measurements,” Phys. Rev. Lett. 92, 230802 (2004).
- Wilpers et al. [2007] G. Wilpers, C. W. Oates, S. A. Diddams, A. Bartels, T. M. Fortier, W. H. Oskay, J. C. Bergquist, S. R. Jefferts, T. P. Heavner, T. E. Parker, et al., “Absolute frequency measurement of the neutral 40Ca optical frequency standard at 657 nm based on microkelvin atoms,” Metrologia 44, 146 (2007).
- Katori [2002] H. Katori, “Spectroscopy of strontium atoms in the Lamb-Dicke confinement, in Proc. 6th Symp. on Frequency Standards and Metrology (ed. Gill, P.),” in Frequency Standards and Metrology (World Scientific, 2002) pp. 323–330.
- Katori et al. [2003] H. Katori, M. Takamoto, V. G. Pal’chikov, and V. D. Ovsiannikov, “Ultrastable optical clock with neutral atoms in an engineered light shift trap,” Phys. Rev. Lett. 91, 173005 (2003).
- Poli et al. [2013] N. Poli, C. W. Oates, P. Gill, and G. M. Tino, “Optical atomic clocks,” La Rivista del Nuovo Cimento 036, 555–624 (2013).
- Ludlow et al. [2015] A. D. Ludlow, M. M. Boyd, J. Ye, E. Peik, and P. O. Schmidt, “Optical atomic clocks,” Rev. Mod. Phys. 87, 637–701 (2015).
- Allan [1966] D. Allan, “Statistics of atomic frequency standards,” Proc. IEEE 54, 221–230 (1966).
- Itano et al. [1993] W. M. Itano, J. C. Bergquist, J. J. Bollinger, J. M. Gilligan, D. J. Heinzen, F. L. Moore, M. G. Raizen, and D. J. Wineland, “Quantum projection noise: Population fluctuations in two-level systems,” Phys. Rev. A 47, 3554–3570 (1993).
- Wineland et al. [1994] D. J. Wineland, J. J. Bollinger, W. M. Itano, and D. J. Heinzen, “Squeezed atomic states and projection noise in spectroscopy,” Phys. Rev. A 50, 67–88 (1994).
- Pedrozo-Peñafiel et al. [2020] E. Pedrozo-Peñafiel, S. Colombo, C. Shu, A. F. Adiyatullin, Z. Li, E. Mendez, B. Braverman, A. Kawasaki, D. Akamatsu, Y. Xiao, et al., “Entanglement on an optical atomic-clock transition,” Nature 588, 414–418 (2020).
- Arecchi et al. [1972] F. T. Arecchi, E. Courtens, R. Gilmore, and H. Thomas, “Atomic coherent states in quantum optics,” Phys. Rev. A 6, 2211–2237 (1972).
- Young et al. [2020] A. W. Young, W. J. Eckner, W. R. Milner, D. Kedar, M. A. Norcia, E. Oelker, N. Schine, J. Ye, and A. M. Kaufman, “Half-minute-scale atomic coherence and high relative stability in a tweezer clock,” Nature 588, 408–413 (2020).
- Huang et al. [2022] Y. Huang, B. Zhang, M. Zeng, Y. Hao, Z. Ma, H. Zhang, H. Guan, Z. Chen, M. Wang, and K. Gao, “Liquid-nitrogen-cooled + optical clock with systematic uncertainty of ,” Phys. Rev. Appl. 17, 034041 (2022).
- Steinel et al. [2023] M. Steinel, H. Shao, M. Filzinger, B. Lipphardt, M. Brinkmann, A. Didier, T. E. Mehlstäubler, T. Lindvall, E. Peik, and N. Huntemann, “Evaluation of a 88Sr+ optical clock with a direct measurement of the blackbody radiation shift and determination of the clock frequency,” Phys. Rev. Lett. 131, 083002 (2023).
- Arnold et al. [2020] K. J. Arnold, R. Kaewuam, S. R. Chanu, T. R. Tan, Z. Zhang, and M. D. Barrett, “Precision measurements of the 138Ba+ clock transition,” Phys. Rev. Lett. 124, 193001 (2020).
- Holliman et al. [2022] C. A. Holliman, M. Fan, A. Contractor, S. M. Brewer, and A. M. Jayich, “Radium ion optical clock,” Phys. Rev. Lett. 128, 033202 (2022).
- Lange et al. [2021a] R. Lange, A. A. Peshkov, N. Huntemann, C. Tamm, A. Surzhykov, and E. Peik, “Lifetime of the level in Yb+ for spontaneous emission of electric octupole radiation,” Phys. Rev. Lett. 127, 213001 (2021a).
- Akatsuka, Takamoto, and Katori [2010] T. Akatsuka, M. Takamoto, and H. Katori, “Three-dimensional optical lattice clock with bosonic atoms,” Phys. Rev. A 81, 023402 (2010).
- Poli et al. [2008] N. Poli, Z. W. Barber, N. D. Lemke, C. W. Oates, L. S. Ma, J. E. Stalnaker, T. M. Fortier, S. A. Diddams, L. Hollberg, J. C. Bergquist, et al., “Frequency evaluation of the doubly forbidden transition in bosonic ,” Phys. Rev. A 77, 050501 (2008).
- Rosenband et al. [2007] T. Rosenband, P. O. Schmidt, D. B. Hume, W. M. Itano, T. M. Fortier, J. E. Stalnaker, K. Kim, S. A. Diddams, J. C. J. Koelemeij, J. C. Bergquist, et al., “Observation of the 1SP0 clock transition in ,” Phys. Rev. Lett. 98, 220801 (2007).
- Pelzer et al. [2021] L. Pelzer, K. Dietze, J. Kramer, F. Dawel, L. Krinner, N. Spethmann, V. Martinez, N. Aharon, A. Retzker, K. Hammerer, et al., “Tailored optical clock transition in 40Ca+,” Measurement: Sensors 18, 100326 (2021).
- Dubé, Madej, and Jian [2016] P. Dubé, A. A. Madej, and B. Jian, “Sr+ single-ion clock,” J. Phys. Conf. Ser. 723, 012018 (2016).
- Schacht et al. [2015] M. Schacht, J. R. Danielson, S. Rahaman, J. R. Torgerson, J. Zhang, and M. M. Schauer, “171Yb+ 5D3/2 hyperfine state detection and F = 2 lifetime,” J. Phys. B: At. Mol. Opt. Phys. 48, 065003 (2015).
- Calamai and Johnson [1990] A. G. Calamai and C. E. Johnson, “Radiative lifetimes of metastable states of Hg+ and Hg2+,” Phys. Rev. A 42, 5425–5432 (1990).
- Muniz et al. [2021] J. A. Muniz, D. J. Young, J. R. K. Cline, and J. K. Thompson, “Cavity-QED measurements of the millihertz optical clock transition and determination of its natural linewidth,” Phys. Rev. Res. 3, 023152 (2021).
- Xu et al. [2014] C.-Y. Xu, J. Singh, J. C. Zappala, K. G. Bailey, M. R. Dietrich, J. P. Greene, W. Jiang, N. D. Lemke, Z.-T. Lu, P. Mueller, et al., “Measurement of the hyperfine quenching rate of the clock transition in ,” Phys. Rev. Lett. 113, 033003 (2014).
- Schäfer et al. [2017] F. Schäfer, H. Konishi, A. Bouscal, T. Yagami, and Y. Takahashi, “Spectroscopic determination of magnetic-field-dependent interactions in an ultracold Yb(3P2)-Li mixture,” Phys. Rev. A 96, 032711 (2017).
- Dzuba, Flambaum, and Schiller [2018] V. A. Dzuba, V. V. Flambaum, and S. Schiller, “Testing physics beyond the standard model through additional clock transitions in neutral ytterbium,” Phys. Rev. A 98, 022501 (2018).
- Safronova et al. [2018] M. S. Safronova, S. G. Porsev, C. Sanner, and J. Ye, “Two clock transitions in neutral Yb for the highest sensitivity to variations of the fine-structure constant,” Phys. Rev. Lett. 120, 173001 (2018).
- Kawasaki et al. [2023] A. Kawasaki, T. Kobayashi, A. Nishiyama, T. Tanabe, and M. Yasuda, “Observation of the clock transition at 431 nm in ,” Phys. Rev. A 107, L060801 (2023).
- Ishiyama et al. [2023] T. Ishiyama, K. Ono, T. Takano, A. Sunaga, and Y. Takahashi, “Observation of an inner-shell orbital clock transition in neutral ytterbium atoms,” Phys. Rev. Lett. 130, 153402 (2023).
- Qiao et al. [2024] H. Qiao, D. Ai, C.-Y. Sun, C.-Q. Peng, Q.-C. Qi, C.-C. Zhao, L.-M. Luo, T.-Y. Jin, T. Zhang, M. Zhou, et al., “Investigation of the () clock transition in atoms,” Phys. Rev. X 14, 011023 (2024).
- Kawasaki et al. [2024] A. Kawasaki, T. Kobayashi, A. Nishiyama, T. Tanabe, and M. Yasuda, “Isotope-shift analysis with the – transition in ytterbium,” Phys. Rev. A 109, 062806 (2024).
- Kulosa et al. [2015] A. P. Kulosa, D. Fim, K. H. Zipfel, S. Rühmann, S. Sauer, N. Jha, K. Gibble, W. Ertmer, E. M. Rasel, M. S. Safronova, et al., “Towards a Mg lattice clock: Observation of the transition and determination of the magic wavelength,” Phys. Rev. Lett. 115, 240801 (2015).
- Yamaguchi et al. [2019a] A. Yamaguchi, M. S. Safronova, K. Gibble, and H. Katori, “Narrow-line cooling and determination of the magic wavelength of Cd,” Phys. Rev. Lett. 123, 113201 (2019a).
- McGrew et al. [2018b] W. F. McGrew, X. Zhang, R. J. Fasano, S. A. Schäffer, K. Beloy, D. Nicolodi, R. C. Brown, N. Hinkley, G. Milani, M. Schioppo, et al., “Atomic clock performance enabling geodesy below the centimetre level,” Nature 564, 87–90 (2018b).
- McFerran et al. [2012] J. J. McFerran, L. Yi, S. Mejri, S. Di Manno, W. Zhang, J. Guéna, Y. Le Coq, and S. Bize, “Neutral atom frequency reference in the deep ultraviolet with ,” Phys. Rev. Lett. 108, 183004 (2012).
- Wu et al. [2023] L. Wu, X. Wang, T. Wang, J. Jiang, and C. Dong, “Be optical lattice clocks with the fractional stark shift up to the level of 10-19,” New J. Phys. 25, 043011 (2023).
- Abe et al. [2021] M. Abe, P. Adamson, M. Borcean, D. Bortoletto, K. Bridges, S. P. Carman, S. Chattopadhyay, J. Coleman, N. M. Curfman, K. DeRose, et al., “Matter-wave atomic gradiometer interferometric sensor (MAGIS-100),” Quantum Sci. Technol. 6, 044003 (2021).
- Norcia, Young, and Kaufman [2018] M. A. Norcia, A. W. Young, and A. M. Kaufman, “Microscopic control and detection of ultracold strontium in optical-tweezer arrays,” Phys. Rev. X 8, 041054 (2018).
- Cooper et al. [2018] A. Cooper, J. P. Covey, I. S. Madjarov, S. G. Porsev, M. S. Safronova, and M. Endres, “Alkaline-earth atoms in optical tweezers,” Phys. Rev. X 8, 041055 (2018).
- Zhang et al. [2023a] Z. Zhang, K. J. Arnold, R. Kaewuam, and M. D. Barrett, “176Lu+ clock comparison at the level via correlation spectroscopy,” Sci. Adv. 9, eadg1971 (2023a).
- Hinkley et al. [2013] N. Hinkley, J. A. Sherman, N. B. Phillips, M. Schioppo, N. D. Lemke, K. Beloy, M. Pizzocaro, C. W. Oates, and A. D. Ludlow, “An atomic clock with instability,” Science 341, 1215–1218 (2013).
- Zhang et al. [2017] W. Zhang, J. M. Robinson, L. Sonderhouse, E. Oelker, C. Benko, J. L. Hall, T. Legero, D. G. Matei, F. Riehle, U. Sterr, et al., “Ultrastable silicon cavity in a continuously operating closed-cycle cryostat at 4 K,” Phys. Rev. Lett. 119, 243601 (2017).
- Robinson et al. [2019] J. M. Robinson, E. Oelker, W. R. Milner, W. Zhang, T. Legero, D. G. Matei, F. Riehle, U. Sterr, and J. Ye, “Crystalline optical cavity at 4 K with thermal-noise-limited instability and ultralow drift,” Optica 6, 240–243 (2019).
- Dick [1987] G. J. Dick, “Local oscillator induced instabilities in trapped ion frequency standards,” Proceedings of the 19th Annual Precise Time and Time Interval Systems and Applications Meeting , 133–147 (1987).
- Santarelli et al. [1998] G. Santarelli, C. Audoin, A. Makdissi, P. Laurent, G. Dick, and A. Clairon, “Frequency stability degradation of an oscillator slaved to a periodically interrogated atomic resonator,” IEEE Trans. Ultrason. Ferroelect. Freq. Contr. 45, 887–894 (1998).
- Schioppo et al. [2017] M. Schioppo, R. C. Brown, W. F. McGrew, N. Hinkley, R. J. Fasano, K. Beloy, T. H. Yoon, G. Milani, D. Nicolodi, J. A. Sherman, et al., “Ultrastable optical clock with two cold-atom ensembles,” Nat. Photon. 11, 48–52 (2017).
- Zheng et al. [2022a] X. Zheng, J. Dolde, V. Lochab, B. N. Merriman, H. Li, and S. Kolkowitz, “Differential clock comparisons with a multiplexed optical lattice clock,” Nature 602, 425–430 (2022a).
- Bothwell et al. [2022] T. Bothwell, C. J. Kennedy, A. Aeppli, D. Kedar, J. M. Robinson, E. Oelker, A. Staron, and J. Ye, “Resolving the gravitational redshift across a millimetre-scale atomic sample,” Nature 602, 420–424 (2022).
- Lodewyck et al. [2012] J. Lodewyck, M. Zawada, L. Lorini, M. Gurov, and P. Lemonde, “Observation and cancellation of a perturbing dc stark shift in strontium optical lattice clocks,” IEEE Trans. Ultrason. Ferroelectr. Freq. Control 59, 411–415 (2012).
- Beloy et al. [2018] K. Beloy, X. Zhang, W. F. McGrew, N. Hinkley, T. H. Yoon, D. Nicolodi, R. J. Fasano, S. A. Schäffer, R. C. Brown, and A. D. Ludlow, “Faraday-shielded dc stark-shift-free optical lattice clock,” Phys. Rev. Lett. 120, 183201 (2018).
- Beloy et al. [2012] K. Beloy, J. A. Sherman, N. D. Lemke, N. Hinkley, C. W. Oates, and A. D. Ludlow, “Determination of the state lifetime and blackbody-radiation clock shift in Yb,” Phys. Rev. A 86, 051404 (2012).
- Beloy et al. [2014] K. Beloy, N. Hinkley, N. B. Phillips, J. A. Sherman, M. Schioppo, J. Lehman, A. Feldman, L. M. Hanssen, C. W. Oates, and A. D. Ludlow, “Atomic clock with room-temperature blackbody stark uncertainty,” Phys. Rev. Lett. 113, 260801 (2014).
- Takamoto, Tanaka, and Katori [2022] M. Takamoto, Y. Tanaka, and H. Katori, “A perspective on the future of transportable optical lattice clocks,” Appl. Phys. Lett. 120, 140502 (2022).
- Golovizin et al. [2024] A. Golovizin, D. Mishin, D. Provorchenko, D. Tregubov, and N. Kolachevsky, “Synchronous comparison of two thulium optical clocks,” JETP Lett. 119, 659–664 (2024).
- Westergaard et al. [2011] P. G. Westergaard, J. Lodewyck, L. Lorini, A. Lecallier, E. A. Burt, M. Zawada, J. Millo, and P. Lemonde, “Lattice-induced frequency shifts in Sr optical lattice clocks at the level,” Phys. Rev. Lett. 106, 210801 (2011).
- Ushijima, Takamoto, and Katori [2018] I. Ushijima, M. Takamoto, and H. Katori, “Operational magic intensity for Sr optical lattice clocks,” Phys. Rev. Lett. 121, 263202 (2018).
- Kim et al. [2023a] K. Kim, A. Aeppli, T. Bothwell, and J. Ye, “Evaluation of lattice light shift at low uncertainty for a shallow lattice Sr optical clock,” Phys. Rev. Lett. 130, 113203 (2023a).
- Bothwell et al. [2025a] T. Bothwell, B. D. Hunt, J. L. Siegel, Y. S. Hassan, T. Grogan, T. Kobayashi, K. Gibble, S. G. Porsev, M. S. Safronova, R. C. Brown, et al., “Lattice light shift evaluations in a dual-ensemble Yb optical lattice clock,” Phys. Rev. Lett. 134, 033201 (2025a).
- Ovsiannikov et al. [2016] V. D. Ovsiannikov, S. I. Marmo, V. G. Palchikov, and H. Katori, “Higher-order effects on the precision of clocks of neutral atoms in optical lattices,” Phys. Rev. A 93, 043420 (2016).
- Campbell et al. [2017] S. L. Campbell, R. B. Hutson, G. E. Marti, A. Goban, N. D. Oppong, R. L. McNally, L. Sonderhouse, J. M. Robinson, W. Zhang, B. J. Bloom, et al., “A Fermi-degenerate three-dimensional optical lattice clock,” Science 358, 90–94 (2017).
- Cidrim et al. [2021] A. Cidrim, A. Piñeiro Orioli, C. Sanner, R. B. Hutson, J. Ye, R. Bachelard, and A. M. Rey, “Dipole-dipole frequency shifts in multilevel atoms,” Phys. Rev. Lett. 127, 013401 (2021).
- Brown et al. [2017] R. C. Brown, N. B. Phillips, K. Beloy, W. F. McGrew, M. Schioppo, R. J. Fasano, G. Milani, X. Zhang, N. Hinkley, H. Leopardi, et al., “Hyperpolarizability and operational magic wavelength in an optical lattice clock,” Phys. Rev. Lett. 119, 253001 (2017).
- Siegel et al. [2024] J. L. Siegel, W. F. McGrew, Y. S. Hassan, C.-C. Chen, K. Beloy, T. Grogan, X. Zhang, and A. D. Ludlow, “Excited-band coherent delocalization for improved optical lattice clock performance,” Phys. Rev. Lett. 132, 133201 (2024).
- Koller et al. [2017] S. B. Koller, J. Grotti, S. Vogt, A. Al-Masoudi, S. Dörscher, S. Häfner, U. Sterr, and C. Lisdat, “Transportable optical lattice clock with uncertainty,” Phys. Rev. Lett. 118, 073601 (2017).
- Takamoto et al. [2020] M. Takamoto, I. Ushijima, N. Ohmae, T. Yahagi, K. Kokado, H. Shinkai, and H. Katori, “Test of general relativity by a pair of transportable optical lattice clocks,” Nat. Photon. 14, 411–415 (2020).
- Bothwell et al. [2025b] T. Bothwell, W. Brand, R. Fasano, T. Akin, J. Whalen, T. Grogan, Y.-J. Chen, M. Pomponio, T. Nakamura, B. Rauf, et al., “Deployment of a transportable Yb optical lattice clock,” Opt. Lett. 50, 646–649 (2025b).
- Amy-Klein et al. [2024] A. Amy-Klein, E. Benkler, P. Blondé, K. Bongs, E. Cantin, C. Chardonnet, H. Denker, S. Dörscher, C.-H. Feng, J.-O. Gaudron, et al. (International Clock and Oscillator Networking (ICON) Collaboration), “International comparison of optical frequencies with transportable optical lattice clocks,” (2024), arXiv:2410.22973 [physics.atom-ph] .
- Lisdat et al. [2016] C. Lisdat, G. Grosche, N. Quintin, C. Shi, S. M. F. Raupach, C. Grebing, D. Nicolodi, F. Stefani, A. Al-Masoudi, S. Dörscher, et al., “A clock network for geodesy and fundamental science,” Nat. Commun. 7, 12443 (2016).
- Clivati et al. [2022] C. Clivati, M. Pizzocaro, E. Bertacco, S. Condio, G. Costanzo, S. Donadello, I. Goti, M. Gozzelino, F. Levi, A. Mura, et al., “Coherent optical-fiber link across Italy and France,” Phys. Rev. Appl. 18, 054009 (2022).
- Beloy et al. [2021] K. Beloy, M. I. Bodine, T. Bothwell, S. M. Brewer, S. L. Bromley, J.-S. Chen, J.-D. Deschênes, S. A. Diddams, R. J. Fasano, T. M. Fortier, et al. (BACON Collaboration), “Frequency ratio measurements at 18-digit accuracy using an optical clock network,” Nature 591, 564–569 (2021).
- Takano et al. [2016] T. Takano, M. Takamoto, I. Ushijima, N. Ohmae, T. Akatsuka, A. Yamaguchi, Y. Kuroishi, H. Munekane, B. Miyahara, and H. Katori, “Geopotential measurements with synchronously linked optical lattice clocks,” Nat. Photon. 10, 662–666 (2016).
- Zheng et al. [2023] X. Zheng, J. Dolde, M. C. Cambria, H. M. Lim, and S. Kolkowitz, “A lab-based test of the gravitational redshift with a miniature clock network,” Nat. Commun. 14, 4886 (2023).
- Zheng, Dolde, and Kolkowitz [2024] X. Zheng, J. Dolde, and S. Kolkowitz, “Reducing the instability of an optical lattice clock using multiple atomic ensembles,” Phys. Rev. X 14, 011006 (2024).
- Hassan et al. [2024] Y. S. Hassan, T. Kobayashi, T. Bothwell, J. L. Seigel, B. D. Hunt, K. Beloy, K. Gibble, T. Grogan, and A. D. Ludlow, “Ratchet loading and multi-ensemble operation in an optical lattice clock,” Quantum Science and Technology 9, 045023 (2024).
- Okaba et al. [2024] S. Okaba, R. Takeuchi, S. Tsuji, and H. Katori, “Continuous generation of an ultracold atomic beam using crossed moving optical lattices,” Phys. Rev. Appl. 21, 034006 (2024).
- Norcia et al. [2019] M. A. Norcia, A. W. Young, W. J. Eckner, E. Oelker, J. Ye, and A. M. Kaufman, “Seconds-scale coherence on an optical clock transition in a tweezer array,” Science 366, 93–97 (2019).
- Covey et al. [2019] J. P. Covey, I. S. Madjarov, A. Cooper, and M. Endres, “2000-times repeated imaging of strontium atoms in clock-magic tweezer arrays,” Phys. Rev. Lett. 122, 173201 (2019).
- Leroux, Schleier-Smith, and Vuletić [2010] I. D. Leroux, M. H. Schleier-Smith, and V. Vuletić, “Orientation-dependent entanglement lifetime in a squeezed atomic clock,” Phys. Rev. Lett. 104, 250801 (2010).
- Hosten et al. [2016] O. Hosten, N. J. Engelsen, R. Krishnakumar, and M. A. Kasevich, “Measurement noise 100 times lower than the quantum-projection limit using entangled atoms,” Nature 529, 505–508 (2016).
- Braverman et al. [2019] B. Braverman, A. Kawasaki, E. Pedrozo-Peñafiel, S. Colombo, C. Shu, Z. Li, E. Mendez, M. Yamoah, L. Salvi, D. Akamatsu, et al., “Near-unitary spin squeezing in ,” Phys. Rev. Lett. 122, 223203 (2019).
- Robinson et al. [2024] J. M. Robinson, M. Miklos, Y. M. Tso, C. J. Kennedy, T. Bothwell, D. Kedar, J. K. Thompson, and J. Ye, “Direct comparison of two spin-squeezed optical clock ensembles at the level,” Nat. Phys. 20, 208–213 (2024).
- Cao et al. [2024] A. Cao, W. J. Eckner, T. Lukin Yelin, A. W. Young, S. Jandura, L. Yan, K. Kim, G. Pupillo, J. Ye, N. Darkwah Oppong, et al., “Multi-qubit gates and schrödinger cat states in an optical clock,” Nature 634, 315–320 (2024).
- Paul, Osberghaus, and Fischer [1958] W. Paul, O. Osberghaus, and E. Fischer, Berichte Des Wirtschaftsministeriums Nordrhein-Westfalen Nr. 415 (1958).
- Fischer [1959] E. Fischer, “Die dreidimensionale stabilisierung von ladungsträgern in einem vierpolfeld,” Z. Phys. 156, 1–26 (1959).
- Paul [1990] W. Paul, “Electromagnetic traps for charged and neutral particles,” Rev. Mod. Phys. 62, 531–540 (1990).
- Schmidt et al. [2005] P. O. Schmidt, T. Rosenband, C. Langer, W. M. Itano, J. C. Bergquist, and D. J. Wineland, “Spectroscopy using quantum logic,” Science 309, 749–752 (2005).
- Arnold et al. [2018] K. J. Arnold, R. Kaewuam, A. Roy, T. R. Tan, and M. D. Barrett, “Blackbody radiation shift assessment for a lutetium ion clock,” Nat. Commun. 9, 1650 (2018).
- Zhang et al. [2016] P. Zhang, J. Cao, H. lin Shu, J. bo Yuan, J. juan Shang, K. feng Cui, S. jia Chao, S. mao Wang, D. xin Liu, and X. ren Huang, “Evaluation of blackbody radiation shift with temperature-associated fractional uncertainty at level for 40Ca+ ion optical clock,” J. Phys. B: At. Mol. Opt. Phys. 50, 015002 (2016).
- Yudin et al. [2010] V. I. Yudin, A. V. Taichenachev, C. W. Oates, Z. W. Barber, N. D. Lemke, A. D. Ludlow, U. Sterr, C. Lisdat, and F. Riehle, “Hyper-ramsey spectroscopy of optical clock transitions,” Phys. Rev. A 82, 011804 (2010).
- Huntemann et al. [2012] N. Huntemann, B. Lipphardt, M. Okhapkin, C. Tamm, E. Peik, A. V. Taichenachev, and V. I. Yudin, “Generalized Ramsey excitation scheme with suppressed light shift,” Phys. Rev. Lett. 109, 213002 (2012).
- Brewer et al. [2019b] S. M. Brewer, J.-S. Chen, K. Beloy, A. M. Hankin, E. R. Clements, C. W. Chou, W. F. McGrew, X. Zhang, R. J. Fasano, D. Nicolodi, et al., “Measurements of and magnetic constants for improved ion-clock accuracy,” Phys. Rev. A 100, 013409 (2019b).
- Herschbach et al. [2012] N. Herschbach, K. Pyka, J. Keller, and T. E. Mehlstäubler, “Linear Paul trap design for an optical clock with Coulomb crystals,” Appl. Phys. B 107, 891–906 (2012).
- Arnold et al. [2015] K. Arnold, E. Hajiyev, E. Paez, C. H. Lee, M. D. Barrett, and J. Bollinger, “Prospects for atomic clocks based on large ion crystals,” Phys. Rev. A 92, 032108 (2015).
- Jiang, Arora, and Safronova [2008] D. Jiang, B. Arora, and M. S. Safronova, “Electric quadrupole moments of metastable states of , , and ,” Phys. Rev. A 78, 022514 (2008).
- Beloy, Leibrandt, and Itano [2017] K. Beloy, D. R. Leibrandt, and W. M. Itano, “Hyperfine-mediated electric quadrupole shifts in and ion clocks,” Phys. Rev. A 95, 043405 (2017).
- Shaniv et al. [2019] R. Shaniv, N. Akerman, T. Manovitz, Y. Shapira, and R. Ozeri, “Quadrupole shift cancellation using dynamic decoupling,” Phys. Rev. Lett. 122, 223204 (2019).
- Cui et al. [2022] K. Cui, J. Valencia, K. T. Boyce, E. R. Clements, D. R. Leibrandt, and D. B. Hume, “Scalable quantum logic spectroscopy,” Phys. Rev. Lett. 129, 193603 (2022).
- Kim et al. [2023b] M. E. Kim, W. F. McGrew, N. V. Nardelli, E. R. Clements, Y. S. Hassan, X. Zhang, J. L. Valencia, H. Leopardi, D. B. Hume, T. M. Fortier, et al., “Improved interspecies optical clock comparisons through differential spectroscopy,” Nat. Phys. 19, 25–29 (2023b).
- Tino et al. [2020] G. Tino, L. Cacciapuoti, S. Capozziello, G. Lambiase, and F. Sorrentino, “Precision gravity tests and the Einstein equivalence principle,” Prog. Part. Nucl. Phys. 112, 103772 (2020).
- Will [2006] C. M. Will, “The confrontation between general relativity and experiment,” Living Rev. Relativity 9, 3 (2006).
- DIRAC [1937] P. A. M. DIRAC, “The cosmological constants,” Nature 139, 323–323 (1937).
- Chodos and Detweiler [1980] A. Chodos and S. Detweiler, “Where has the fifth dimension gone?” Phys. Rev. D 21, 2167–2170 (1980).
- Wu and Wang [1986] Y.-S. Wu and Z. Wang, “Time variation of Newton’s gravitational constant in superstring theories,” Phys. Rev. Lett. 57, 1978–1981 (1986).
- Kiritsis [1999] E. Kiritsis, “Supergravity, D-brane probes and thermal super Yang-Mills: a comparison,” J. High Energy Phys. 1999, 010 (1999).
- Martins [2017] C. J. A. P. Martins, “The status of varying constants: a review of the physics, searches and implications,” Rep. Prog. Phys. 80, 126902 (2017).
- Bahcall and Schmidt [1967] J. N. Bahcall and M. Schmidt, “Does the fine-structure constant vary with cosmic time?” Phys. Rev. Lett. 19, 1294–1295 (1967).
- SHLYAKHTER [1976] A. I. SHLYAKHTER, “Direct test of the constancy of fundamental nuclear constants,” Nature 264, 340–340 (1976).
- Webb et al. [2001] J. K. Webb, M. T. Murphy, V. V. Flambaum, V. A. Dzuba, J. D. Barrow, C. W. Churchill, J. X. Prochaska, and A. M. Wolfe, “Further evidence for cosmological evolution of the fine structure constant,” Phys. Rev. Lett. 87, 091301 (2001).
- Wilczynska et al. [2020] M. R. Wilczynska, J. K. Webb, M. Bainbridge, J. D. Barrow, S. E. I. Bosman, R. F. Carswell, M. P. Dąbrowski, V. Dumont, C.-C. Lee, A. C. Leite, et al., “Four direct measurements of the fine-structure constant 13 billion years ago,” Sci. Adv. 6, eaay9672 (2020).
- Murphy et al. [2022] M. T. Murphy, P. Molaro, A. C. O. Leite, G. Cupani, S. Cristiani, V. D’Odorico, R. Génova Santos, C. J. A. P. Martins, D. Milaković, N. J. Nunes, et al., “Fundamental physics with ESPRESSO: Precise limit on variations in the fine-structure constant towards the bright quasar he 0515-4414,” Astron. Astrophys. 658, A123 (2022).
- Wang et al. [2024] Z.-F. Wang, L. Lei, L. Feng, and Y.-Z. Fan, “JWST observations constrain the time evolution of fine structure constants and dark energy-electromagnetic coupling,” Res. Astron. Astrophys. 24, 125012 (2024).
- Berengut and Flambaum [2012] J. C. Berengut and V. V. Flambaum, “Manifestations of a spatial variation of fundamental constants in atomic and nuclear clocks, Oklo, meteorites, and cosmological phenomena,” Europhys. Lett. 97, 20006 (2012).
- Beeks et al. [2024] K. Beeks, G. A. Kazakov, F. Schaden, I. Morawetz, L. T. de Col, T. Riebner, M. Bartokos, T. Sikorsky, T. Schumm, C. Zhang, et al., “Fine-structure constant sensitivity of the Th-229 nuclear clock transition,” (2024), arXiv:2407.17300 [nucl-th] .
- Flambaum and Tedesco [2006] V. V. Flambaum and A. F. Tedesco, “Dependence of nuclear magnetic moments on quark masses and limits on temporal variation of fundamental constants from atomic clock experiments,” Phys. Rev. C 73, 055501 (2006).
- Angstmann, Dzuba, and Flambaum [2004] E. J. Angstmann, V. A. Dzuba, and V. V. Flambaum, “Relativistic effects in two valence-electron atoms and ions and the search for variation of the fine-structure constant,” Phys. Rev. A 70, 014102 (2004).
- Dzuba, Flambaum, and Webb [1999] V. A. Dzuba, V. V. Flambaum, and J. K. Webb, “Calculations of the relativistic effects in many-electron atoms and space-time variation of fundamental constants,” Phys. Rev. A 59, 230–237 (1999).
- Tang et al. [2023] Z.-M. Tang, Y.-m. Yu, B. K. Sahoo, C.-Z. Dong, Y. Yang, and Y. Zou, “Simultaneous magic trapping conditions for three additional clock transitions in Yb to search for variation of the fine-structure constant,” Phys. Rev. A 107, 053111 (2023).
- Flambaum and Dzuba [2009] V. V. Flambaum and V. A. Dzuba, “Search for variation of the fundamental constants in atomic, molecular, and nuclear spectra,” Can. J. Phys. 87, 25–33 (2009).
- Dzuba and Flambaum [2008] V. A. Dzuba and V. V. Flambaum, “Relativistic corrections to transition frequencies of Ag I, Dy I, Ho I, Yb II, Yb III, Au I, and Hg II and search for variation of the fine-structure constant,” Phys. Rev. A 77, 012515 (2008).
- Safronova et al. [2014a] M. S. Safronova, V. A. Dzuba, V. V. Flambaum, U. I. Safronova, S. G. Porsev, and M. G. Kozlov, “Highly charged ions for atomic clocks, quantum information, and search for variation,” Phys. Rev. Lett. 113, 030801 (2014a).
- Berengut et al. [2011] J. C. Berengut, V. A. Dzuba, V. V. Flambaum, and A. Ong, “Electron-hole transitions in multiply charged ions for precision laser spectroscopy and searching for variations in ,” Phys. Rev. Lett. 106, 210802 (2011).
- Dzuba et al. [2015] V. A. Dzuba, M. S. Safronova, U. I. Safronova, and V. V. Flambaum, “Actinide ions for testing the spatial -variation hypothesis,” Phys. Rev. A 92, 060502 (2015).
- Berengut et al. [2012b] J. C. Berengut, V. A. Dzuba, V. V. Flambaum, and A. Ong, “Optical transitions in highly charged californium ions with high sensitivity to variation of the fine-structure constant,” Phys. Rev. Lett. 109, 070802 (2012b).
- Turneaure and Stein [1976] J. P. Turneaure and S. R. Stein, “An experimental limit on the time variation of the fine structure constant,” in Atomic Masses and Fundamental Constants 5, edited by J. H. Sanders and A. H. Wapstra (Springer US, Boston, MA, 1976) pp. 636–642.
- Godone et al. [1993] A. Godone, C. Novero, P. Tavella, and K. Rahimullah, “New experimental limits to the time variations of (/) and ,” Phys. Rev. Lett. 71, 2364–2366 (1993).
- Prestage, Tjoelker, and Maleki [1995] J. D. Prestage, R. L. Tjoelker, and L. Maleki, “Atomic clocks and variations of the fine structure constant,” Phys. Rev. Lett. 74, 3511–3514 (1995).
- Bize et al. [2003] S. Bize, S. A. Diddams, U. Tanaka, C. E. Tanner, W. H. Oskay, R. E. Drullinger, T. E. Parker, T. P. Heavner, S. R. Jefferts, L. Hollberg, et al., “Testing the stability of fundamental constants with the single-ion optical clock,” Phys. Rev. Lett. 90, 150802 (2003).
- Fortier et al. [2007] T. M. Fortier, N. Ashby, J. C. Bergquist, M. J. Delaney, S. A. Diddams, T. P. Heavner, L. Hollberg, W. M. Itano, S. R. Jefferts, K. Kim, et al., “Precision atomic spectroscopy for improved limits on variation of the fine structure constant and local position invariance,” Phys. Rev. Lett. 98, 070801 (2007).
- Godun et al. [2014] R. M. Godun, P. B. R. Nisbet-Jones, J. M. Jones, S. A. King, L. A. M. Johnson, H. S. Margolis, K. Szymaniec, S. N. Lea, K. Bongs, and P. Gill, “Frequency ratio of two optical clock transitions in 171Yb+ and constraints on the time variation of fundamental constants,” Phys. Rev. Lett. 113, 210801 (2014).
- Huntemann et al. [2014] N. Huntemann, B. Lipphardt, C. Tamm, V. Gerginov, S. Weyers, and E. Peik, “Improved limit on a temporal variation of from comparisons of Yb+ and Cs atomic clocks,” Phys. Rev. Lett. 113, 210802 (2014).
- Lange et al. [2021b] R. Lange, N. Huntemann, J. M. Rahm, C. Sanner, H. Shao, B. Lipphardt, C. Tamm, S. Weyers, and E. Peik, “Improved limits for violations of local position invariance from atomic clock comparisons,” Phys. Rev. Lett. 126, 011102 (2021b).
- Filzinger et al. [2023] M. Filzinger, S. Dörscher, R. Lange, J. Klose, M. Steinel, E. Benkler, E. Peik, C. Lisdat, and N. Huntemann, “Improved limits on the coupling of ultralight bosonic dark matter to photons from optical atomic clock comparisons,” Phys. Rev. Lett. 130, 253001 (2023).
- Marion et al. [2003] H. Marion, F. Pereira Dos Santos, M. Abgrall, S. Zhang, Y. Sortais, S. Bize, I. Maksimovic, D. Calonico, J. Grünert, C. Mandache, et al., “Search for variations of fundamental constants using atomic fountain clocks,” Phys. Rev. Lett. 90, 150801 (2003).
- Cingöz et al. [2007] A. Cingöz, A. Lapierre, A.-T. Nguyen, N. Leefer, D. Budker, S. K. Lamoreaux, and J. R. Torgerson, “Limit on the temporal variation of the fine-structure constant using atomic dysprosium,” Phys. Rev. Lett. 98, 040801 (2007).
- Le Targat et al. [2013] R. Le Targat, L. Lorini, Y. Le Coq, M. Zawada, J. Guéna, M. Abgrall, M. Gurov, P. Rosenbusch, D. G. Rovera, B. Nagórny, et al., “Experimental realization of an optical second with strontium lattice clocks,” Nat. Commun. 4, 2109 (2013).
- Leefer et al. [2013] N. Leefer, C. T. M. Weber, A. Cingöz, J. R. Torgerson, and D. Budker, “New limits on variation of the fine-structure constant using atomic dysprosium,” Phys. Rev. Lett. 111, 060801 (2013).
- McGrew et al. [2019] W. F. McGrew, X. Zhang, H. Leopardi, R. J. Fasano, D. Nicolodi, K. Beloy, J. Yao, J. A. Sherman, S. A. Schäffer, J. Savory, et al., “Towards the optical second: verifying optical clocks at the SI limit,” Optica 6, 448–454 (2019).
- Chin, Flambaum, and Kozlov [2009] C. Chin, V. V. Flambaum, and M. G. Kozlov, “Ultracold molecules: new probes on the variation of fundamental constants,” New Journal of Physics 11, 055048 (2009).
- Shelkovnikov et al. [2008] A. Shelkovnikov, R. J. Butcher, C. Chardonnet, and A. Amy-Klein, “Stability of the proton-to-electron mass ratio,” Phys. Rev. Lett. 100, 150801 (2008).
- Kobayashi, Ogino, and Inouye [2019] J. Kobayashi, A. Ogino, and S. Inouye, “Measurement of the variation of electron-to-proton mass ratio using ultracold molecules produced from laser-cooled atoms,” Nat. Commun. 10, 3771 (2019).
- Kozlov et al. [2018] M. G. Kozlov, M. S. Safronova, J. R. Crespo López-Urrutia, and P. O. Schmidt, “Highly charged ions: Optical clocks and applications in fundamental physics,” Rev. Mod. Phys. 90, 045005 (2018).
- Johnson and Soff [1985] W. Johnson and G. Soff, “The Lamb shift in hydrogen-like atoms, ,” At. Data Nucl. Data Tables 33, 405–446 (1985).
- Loetzsch et al. [2024] R. Loetzsch, H. F. Beyer, L. Duval, U. Spillmann, D. Banaś, P. Dergham, F. M. Kröger, J. Glorius, R. E. Grisenti, M. Guerra, et al., “Testing quantum electrodynamics in extreme fields using helium-like uranium,” Nature 625, 673–678 (2024).
- Berengut, Dzuba, and Flambaum [2010] J. C. Berengut, V. A. Dzuba, and V. V. Flambaum, “Enhanced laboratory sensitivity to variation of the fine-structure constant using highly charged ions,” Phys. Rev. Lett. 105, 120801 (2010).
- Rehbehn et al. [2021] N.-H. Rehbehn, M. K. Rosner, H. Bekker, J. C. Berengut, P. O. Schmidt, S. A. King, P. Micke, M. F. Gu, R. Müller, A. Surzhykov, et al., “Sensitivity to new physics of isotope-shift studies using the coronal lines of highly charged calcium ions,” Phys. Rev. A 103, L040801 (2021).
- Yudin, Taichenachev, and Derevianko [2014] V. I. Yudin, A. V. Taichenachev, and A. Derevianko, “Magnetic-dipole transitions in highly charged ions as a basis of ultraprecise optical clocks,” Phys. Rev. Lett. 113, 233003 (2014).
- Porsev et al. [2024] S. G. Porsev, C. Cheung, M. S. Safronova, H. Bekker, N.-H. Rehbehn, J. R. C. López-Urrutia, and S. M. Brewer, “Pr10+ as a candidate for a high-accuracy optical clock for tests of fundamental physics,” Phys. Rev. A 110, 042823 (2024).
- Safronova et al. [2014b] M. S. Safronova, V. A. Dzuba, V. V. Flambaum, U. I. Safronova, S. G. Porsev, and M. G. Kozlov, “Highly charged Ag-like and In-like ions for the development of atomic clocks and the search for variation,” Phys. Rev. A 90, 042513 (2014b).
- Dzuba, Flambaum, and Katori [2015] V. A. Dzuba, V. V. Flambaum, and H. Katori, “Optical clock sensitive to variations of the fine-structure constant based on the Ho14+ ion,” Phys. Rev. A 91, 022119 (2015).
- Nandy and Sahoo [2016] D. K. Nandy and B. K. Sahoo, “Highly charged , and ions as promising optical clock candidates for probing variations of the fine-structure constant,” Phys. Rev. A 94, 032504 (2016).
- Dzuba and Flambaum [2024] V. A. Dzuba and V. V. Flambaum, “Highly charged ions of heavy actinides as sensitive probes for time variation of the fine-structure constant,” Phys. Rev. A 110, 012801 (2024).
- Fawcett [1981] B. C. Fawcett, “Classification of the spectra of highly ionised atoms during the last seven years,” Phys. Scr. 24, 663 (1981).
- Geller [1970] R. Geller, “New high intensity ion source with very low extraction voltage,” Appl. Phys. Lett. 16, 401–404 (1970), https://pubs.aip.org/aip/apl/article-pdf/16/10/401/18423299/401_1_online.pdf .
- Arianer et al. [1975] J. Arianer, E. Baron, M. Brient, A. Cabrespine, A. Liebe, A. Sérafini, and T. Ton That, “Multiply charged ion source,” Nucl. Instrum. Methods 124, 157–160 (1975).
- Prior [1987] M. H. Prior, “Forbidden lines from highly charged, metastable ion beams,” J. Opt. Soc. Am. B 4, 144–147 (1987).
- Klaft et al. [1994] I. Klaft, S. Borneis, T. Engel, B. Fricke, R. Grieser, G. Huber, T. Kühl, D. Marx, R. Neumann, S. Schröder, P. Seelig, and L. Völker, “Precision laser spectroscopy of the ground state hyperfine splitting of hydrogenlike ,” Phys. Rev. Lett. 73, 2425–2427 (1994).
- Levine et al. [1988] M. A. Levine, R. E. Marrs, J. R. Henderson, D. A. Knapp, and M. B. Schneider, “The electron beam ion trap: A new instrument for atomic physics measurements,” Phys. Scr. 1988, 157 (1988).
- Schmöger et al. [2015] L. Schmöger, M. Schwarz, T. M. Baumann, O. O. Versolato, B. Piest, T. Pfeifer, J. Ullrich, P. O. Schmidt, and J. R. Crespo López-Urrutia, “Deceleration, precooling, and multi-pass stopping of highly charged ions in Be+ coulomb crystals,” Rev. Sci. Instrum. 86, 103111 (2015).
- Leopold et al. [2019] T. Leopold, S. A. King, P. Micke, A. Bautista-Salvador, J. C. Heip, C. Ospelkaus, J. R. Crespo López-Urrutia, and P. O. Schmidt, “A cryogenic radio-frequency ion trap for quantum logic spectroscopy of highly charged ions,” Rev. Sci. Instrum. 90, 073201 (2019).
- King et al. [2022] S. A. King, L. J. Spieß, P. Micke, A. Wilzewski, T. Leopold, E. Benkler, R. Lange, N. Huntemann, A. Surzhykov, V. A. Yerokhin, et al., “An optical atomic clock based on a highly charged ion,” Nature 611, 43–47 (2022).
- Bekker et al. [2019] H. Bekker, A. Borschevsky, Z. Harman, C. H. Keitel, T. Pfeifer, P. O. Schmidt, J. R. Crespo López-Urrutia, and J. C. Berengut, “Detection of the 5p – 4f orbital crossing and its optical clock transition in Pr9+,” Nat. Commun. 10, 5651 (2019).
- Chen et al. [2024b] S. Chen, Z. Zhou, J. Li, T. Zhang, C. Li, T. Shi, Y. Huang, K. Gao, and H. Guan, “Precision measurement of M1 optical clock transition in Ni12+,” Phys. Rev. Res. 6, 013030 (2024b).
- Rehbehn et al. [2023] N.-H. Rehbehn, M. K. Rosner, J. C. Berengut, P. O. Schmidt, T. Pfeifer, M. F. Gu, and J. R. C. López-Urrutia, “Narrow and ultranarrow transitions in highly charged Xe ions as probes of fifth forces,” Phys. Rev. Lett. 131, 161803 (2023).
- Wilzewski et al. [2024] A. Wilzewski, L. I. Huber, M. Door, J. Richter, A. Mariotti, L. J. Spieß, M. Wehrheim, S. Chen, S. A. King, P. Micke, et al., “Nonlinear calcium king plot constrains new bosons and nuclear properties,” (2024), arXiv:2412.10277 [physics.atom-ph] .
- Kroger and Reich [1976] L. Kroger and C. Reich, “Features of the low-energy level scheme of 229Th as observed in the -decay of 233U,” Nucl. Phys. A 259, 29–60 (1976).
- Reich and Helmer [1990] C. W. Reich and R. G. Helmer, “Energy separation of the doublet of intrinsic states at the ground state of ,” Phys. Rev. Lett. 64, 271–273 (1990).
- Helmer and Reich [1994] R. G. Helmer and C. W. Reich, “An excited state of at 3.5 eV,” Phys. Rev. C 49, 1845–1858 (1994).
- Yamaguchi et al. [2019b] A. Yamaguchi, H. Muramatsu, T. Hayashi, N. Yuasa, K. Nakamura, M. Takimoto, H. Haba, K. Konashi, M. Watanabe, H. Kikunaga, et al., “Energy of the nuclear clock isomer determined by absolute -ray energy difference,” Phys. Rev. Lett. 123, 222501 (2019b).
- Sikorsky et al. [2020] T. Sikorsky, J. Geist, D. Hengstler, S. Kempf, L. Gastaldo, C. Enss, C. Mokry, J. Runke, C. E. Düllmann, P. Wobrauschek, et al., “Measurement of the isomer energy with a magnetic microcalorimeter,” Phys. Rev. Lett. 125, 142503 (2020).
- Seiferle et al. [2019] B. Seiferle, L. von der Wense, P. V. Bilous, I. Amersdorffer, C. Lemell, F. Libisch, S. Stellmer, T. Schumm, C. E. Düllmann, A. Pálffy, et al., “Energy of the 229Th nuclear clock transition,” Nature 573, 243–246 (2019).
- Kraemer et al. [2023] S. Kraemer, J. Moens, M. Athanasakis-Kaklamanakis, S. Bara, K. Beeks, P. Chhetri, K. Chrysalidis, A. Claessens, T. E. Cocolios, J. G. M. Correia, et al., “Observation of the radiative decay of the 229Th nuclear clock isomer,” Nature 617, 706–710 (2023).
- Tiedau et al. [2024] J. Tiedau, M. V. Okhapkin, K. Zhang, J. Thielking, G. Zitzer, E. Peik, F. Schaden, T. Pronebner, I. Morawetz, L. T. De Col, et al., “Laser excitation of the Th-229 nucleus,” Phys. Rev. Lett. 132, 182501 (2024).
- Elwell et al. [2024] R. Elwell, C. Schneider, J. Jeet, J. E. S. Terhune, H. W. T. Morgan, A. N. Alexandrova, H. B. Tran Tan, A. Derevianko, and E. R. Hudson, “Laser excitation of the nuclear isomeric transition in a solid-state host,” Phys. Rev. Lett. 133, 013201 (2024).
- Zhang et al. [2024a] C. Zhang, T. Ooi, J. S. Higgins, J. F. Doyle, L. von der Wense, K. Beeks, A. Leitner, G. A. Kazakov, P. Li, P. G. Thirolf, et al., “Frequency ratio of the 229mTh nuclear isomeric transition and the 87Sr atomic clock,” Nature 633, 63–70 (2024a).
- Schaden et al. [2024] F. Schaden, T. Riebner, I. Morawetz, L. T. D. Col, G. A. Kazakov, K. Beeks, T. Sikorsky, T. Schumm, K. Zhang, V. Lal, et al., “Laser-induced quenching of the Th-229 nuclear clock isomer in calcium fluoride,” (2024), arXiv:2412.12339 [physics.atom-ph] .
- Zhang et al. [2024b] C. Zhang, L. von der Wense, J. F. Doyle, J. S. Higgins, T. Ooi, H. U. Friebel, J. Ye, R. Elwell, J. E. S. Terhune, H. W. T. Morgan, et al., “229ThF4 thin films for solid-state nuclear clocks,” Nature 636, 603–608 (2024b).
- Higgins et al. [2024] J. S. Higgins, T. Ooi, J. F. Doyle, C. Zhang, J. Ye, K. Beeks, T. Sikorsky, and T. Schumm, “Temperature sensitivity of a thorium-229 solid-state nuclear clock,” (2024), arXiv:2409.11590 [physics.atom-ph] .
- Yamaguchi et al. [2024] A. Yamaguchi, Y. Shigekawa, H. Haba, H. Kikunaga, K. Shirasaki, M. Wada, and H. Katori, “Laser spectroscopy of triply charged 229Th isomer for a nuclear clock,” Nature 629, 62–66 (2024).
- Burke et al. [1990] D. G. Burke, P. E. Garrett, T. Qu, and R. A. Naumann, “Additional evidence for the proposed excited state at 5 eV in ,” Phys. Rev. C 42, R499–R501 (1990).
- Guimarães Filho and Helene [2005] Z. O. Guimarães Filho and O. Helene, “Energy of the state of reexamined,” Phys. Rev. C 71, 044303 (2005).
- Beck et al. [2007] B. R. Beck, J. A. Becker, P. Beiersdorfer, G. V. Brown, K. J. Moody, J. B. Wilhelmy, F. S. Porter, C. A. Kilbourne, and R. L. Kelley, “Energy splitting of the ground-state doublet in the nucleus ,” Phys. Rev. Lett. 98, 142501 (2007).
- Hiraki et al. [2024] T. Hiraki, K. Okai, M. Bartokos, K. Beeks, H. Fujimoto, Y. Fukunaga, H. Haba, Y. Kasamatsu, S. Kitao, A. Leitner, et al., “Controlling 229Th isomeric state population in a VUV transparent crystal,” Nat. Commun. 15, 5536 (2024).
- Flambaum [2006] V. V. Flambaum, “Enhanced effect of temporal variation of the fine structure constant and the strong interaction in ,” Phys. Rev. Lett. 97, 092502 (2006).
- Morgan et al. [2024] H. W. T. Morgan, R. Elwell, J. E. S. Terhune, H. B. T. Tan, U. C. Perera, A. Derevianko, A. N. Alexandrova, and E. R. Hudson, “229Th-doped nonlinear optical crystals for compact solid-state clocks,” (2024), arXiv:2410.23364 [physics.optics] .
- Corbelli and Salucci [2000] E. Corbelli and P. Salucci, “The extended rotation curve and the dark matter halo of M33,” Mon. Not. Roy. Astron. Soc. 311, 441–447 (2000).
- Massey, Kitching, and Richard [2010] R. Massey, T. Kitching, and J. Richard, “The dark matter of gravitational lensing,” Rep. Prog. Phys. 73, 086901 (2010).
- Planck Collaboration [2020] Planck Collaboration, “Planck 2018 results - I. overview and the cosmological legacy of planck,” Astron. Astrophys. 641, A1 (2020).
- Brandt [2016] T. D. Brandt, “Constraints on macho dark matter from compact stellar systems in ultra-faint dwarf galaxies,” Astrophys. J. Lett. 824, L31 (2016).
- Navas et al. [2024] S. Navas, C. Amsler, T. Gutsche, C. Hanhart, J. J. Hernández-Rey, C. Lourenço, A. Masoni, M. Mikhasenko, R. E. Mitchell, C. Patrignani, et al. (Particle Data Group Collaboration), “Review of particle physics,” Phys. Rev. D 110, 030001 (2024).
- Hu, Barkana, and Gruzinov [2000] W. Hu, R. Barkana, and A. Gruzinov, “Fuzzy cold dark matter: The wave properties of ultralight particles,” Phys. Rev. Lett. 85, 1158–1161 (2000).
- Ferreira [2021] E. G. M. Ferreira, “Ultra-light dark matter,” Astron. and Astrophys. Rev. 29, 7 (2021).
- Aghaie et al. [2024] M. Aghaie, G. Armando, A. Dondarini, and P. Panci, “Bounds on ultralight dark matter from NANOGrav,” Phys. Rev. D 109, 103030 (2024).
- Blum et al. [2014] K. Blum, R. T. D’Agnolo, M. Lisanti, and B. R. Safdi, “Constraining axion dark matter with big bang nucleosynthesis,” Phys. Lett. B 737, 30–33 (2014).
- Raffelt [1990] G. G. Raffelt, “Astrophysical methods to constrain axions and other novel particle phenomena,” Phys. Rep. 198, 1–113 (1990).
- Carenza et al. [2019] P. Carenza, T. Fischer, M. Giannotti, G. Guo, G. Martínez-Pinedo, and A. Mirizzi, “Improved axion emissivity from a supernova via nucleon-nucleon bremsstrahlung,” J. Cosmol. Astropart. Phys. 2019, 016 (2019).
- Buschmann et al. [2022] M. Buschmann, C. Dessert, J. W. Foster, A. J. Long, and B. R. Safdi, “Upper limit on the QCD axion mass from isolated neutron star cooling,” Phys. Rev. Lett. 128, 091102 (2022).
- Du et al. [2018] N. Du, N. Force, R. Khatiwada, E. Lentz, R. Ottens, L. J. Rosenberg, G. Rybka, G. Carosi, N. Woollett, D. Bowring, et al. (ADMX Collaboration), “Search for invisible axion dark matter with the axion dark matter experiment,” Phys. Rev. Lett. 120, 151301 (2018).
- Kwon et al. [2021] O. Kwon, D. Lee, W. Chung, D. Ahn, H. Byun, F. Caspers, H. Choi, J. Choi, Y. Chong, H. Jeong, et al., “First results from an axion haloscope at capp around ,” Phys. Rev. Lett. 126, 191802 (2021).
- Chaudhuri et al. [2015] S. Chaudhuri, P. W. Graham, K. Irwin, J. Mardon, S. Rajendran, and Y. Zhao, “Radio for hidden-photon dark matter detection,” Phys. Rev. D 92, 075012 (2015).
- Salemi et al. [2021] C. P. Salemi, J. W. Foster, J. L. Ouellet, A. Gavin, K. M. W. Pappas, S. Cheng, K. A. Richardson, R. Henning, Y. Kahn, R. Nguyen, et al., “Search for low-mass axion dark matter with ABRACADABRA-10 cm,” Phys. Rev. Lett. 127, 081801 (2021).
- Della Valle et al. [2016] F. Della Valle, A. Ejlli, U. Gastaldi, G. Messineo, E. Milotti, R. Pengo, G. Ruoso, and G. Zavattini, “The PVLAS experiment: measuring vacuum magnetic birefringence and dichroism with a birefringent fabry–perot cavity,” Eur. Phys. J. C 76, 24 (2016).
- Ballou et al. [2015] R. Ballou, G. Deferne, M. Finger, M. Finger, L. Flekova, J. Hosek, S. Kunc, K. Macuchova, K. A. Meissner, P. Pugnat, et al. (OSQAR Collaboration), “New exclusion limits on scalar and pseudoscalar axionlike particles from light shining through a wall,” Phys. Rev. D 92, 092002 (2015).
- Graham et al. [2016] P. W. Graham, D. E. Kaplan, J. Mardon, S. Rajendran, and W. A. Terrano, “Dark matter direct detection with accelerometers,” Phys. Rev. D 93, 075029 (2016).
- Bergé et al. [2018] J. Bergé, P. Brax, G. Métris, M. Pernot-Borràs, P. Touboul, and J.-P. Uzan, “MICROSCOPE Mission: First constraints on the violation of the weak equivalence principle by a light scalar dilaton,” Phys. Rev. Lett. 120, 141101 (2018).
- Shaw et al. [2022] E. A. Shaw, M. P. Ross, C. A. Hagedorn, E. G. Adelberger, and J. H. Gundlach, “Torsion-balance search for ultralow-mass bosonic dark matter,” Phys. Rev. D 105, 042007 (2022).
- Abel et al. [2017] C. Abel, N. J. Ayres, G. Ban, G. Bison, K. Bodek, V. Bondar, M. Daum, M. Fairbairn, V. V. Flambaum, P. Geltenbort, et al., “Search for axionlike dark matter through nuclear spin precession in electric and magnetic fields,” Phys. Rev. X 7, 041034 (2017).
- Arvanitaki, Huang, and Van Tilburg [2015] A. Arvanitaki, J. Huang, and K. Van Tilburg, “Searching for dilaton dark matter with atomic clocks,” Phys. Rev. D 91, 015015 (2015).
- Derevianko and Pospelov [2014] A. Derevianko and M. Pospelov, “Hunting for topological dark matter with atomic clocks,” Nat. Phys. 10, 933–936 (2014).
- Roberts et al. [2017] B. M. Roberts, G. Blewitt, C. Dailey, M. Murphy, M. Pospelov, A. Rollings, J. Sherman, W. Williams, and A. Derevianko, “Search for domain wall dark matter with atomic clocks on board global positioning system satellites,” Nat. Commun. 8, 1195 (2017).
- Wcisło et al. [2018] P. Wcisło, P. Ablewski, K. Beloy, S. Bilicki, M. Bober, R. Brown, R. Fasano, R. Ciuryło, H. Hachisu, T. Ido, et al., “New bounds on dark matter coupling from a global network of optical atomic clocks,” Sci. Adv. 4, eaau4869 (2018).
- Van Tilburg et al. [2015] K. Van Tilburg, N. Leefer, L. Bougas, and D. Budker, “Search for ultralight scalar dark matter with atomic spectroscopy,” Phys. Rev. Lett. 115, 011802 (2015).
- Stadnik and Flambaum [2016a] Y. V. Stadnik and V. V. Flambaum, “Improved limits on interactions of low-mass spin-0 dark matter from atomic clock spectroscopy,” Phys. Rev. A 94, 022111 (2016a).
- Stadnik and Flambaum [2015] Y. V. Stadnik and V. V. Flambaum, “Can dark matter induce cosmological evolution of the fundamental constants of nature?” Phys. Rev. Lett. 115, 201301 (2015).
- Kim and Perez [2024] H. Kim and G. Perez, “Oscillations of atomic energy levels induced by QCD axion dark matter,” Phys. Rev. D 109, 015005 (2024).
- Wcisło et al. [2016] P. Wcisło, P. Morzyński, M. Bober, A. Cygan, D. Lisak, R. Ciuryło, and M. Zawada, “Experimental constraint on dark matter detection with optical atomic clocks,” Nat. Astron. 1, 0009 (2016).
- Stadnik and Flambaum [2016b] Y. V. Stadnik and V. V. Flambaum, “Enhanced effects of variation of the fundamental constants in laser interferometers and application to dark-matter detection,” Phys. Rev. A 93, 063630 (2016b).
- Kobayashi et al. [2022] T. Kobayashi, A. Takamizawa, D. Akamatsu, A. Kawasaki, A. Nishiyama, K. Hosaka, Y. Hisai, M. Wada, H. Inaba, T. Tanabe, et al., “Search for ultralight dark matter from long-term frequency comparisons of optical and microwave atomic clocks,” Phys. Rev. Lett. 129, 241301 (2022).
- Kennedy et al. [2020] C. J. Kennedy, E. Oelker, J. M. Robinson, T. Bothwell, D. Kedar, W. R. Milner, G. E. Marti, A. Derevianko, and J. Ye, “Precision metrology meets cosmology: Improved constraints on ultralight dark matter from atom-cavity frequency comparisons,” Phys. Rev. Lett. 125, 201302 (2020).
- Hees et al. [2016] A. Hees, J. Guéna, M. Abgrall, S. Bize, and P. Wolf, “Searching for an oscillating massive scalar field as a dark matter candidate using atomic hyperfine frequency comparisons,” Phys. Rev. Lett. 117, 061301 (2016).
- Sherrill et al. [2023] N. Sherrill, A. O. Parsons, C. F. A. Baynham, W. Bowden, E. A. Curtis, R. Hendricks, I. R. Hill, R. Hobson, H. S. Margolis, B. I. Robertson, et al., “Analysis of atomic-clock data to constrain variations of fundamental constants,” New J. Phys. 25, 093012 (2023).
- Banerjee et al. [2023] A. Banerjee, D. Budker, M. Filzinger, N. Huntemann, G. Paz, G. Perez, S. Porsev, and M. Safronova, “Oscillating nuclear charge radii as sensors for ultralight dark matter,” (2023), arXiv:2301.10784 [hep-ph] .
- Hees et al. [2018] A. Hees, O. Minazzoli, E. Savalle, Y. V. Stadnik, and P. Wolf, “Violation of the equivalence principle from light scalar dark matter,” Phys. Rev. D 98, 064051 (2018).
- Touboul et al. [2022] P. Touboul, G. Métris, M. Rodrigues, J. Bergé, A. Robert, Q. Baghi, Y. André, J. Bedouet, D. Boulanger, S. Bremer, et al. (MICROSCOPE Collaboration), “ mission: Final results of the test of the equivalence principle,” Phys. Rev. Lett. 129, 121102 (2022).
- Touboul et al. [2017] P. Touboul, G. Métris, M. Rodrigues, Y. André, Q. Baghi, J. Bergé, D. Boulanger, S. Bremer, P. Carle, R. Chhun, B. Christophe, V. Cipolla, T. Damour, P. Danto, H. Dittus, P. Fayet, B. Foulon, C. Gageant, P.-Y. Guidotti, D. Hagedorn, E. Hardy, P.-A. Huynh, H. Inchauspe, P. Kayser, S. Lala, C. Lämmerzahl, V. Lebat, P. Leseur, F. m. c. Liorzou, M. List, F. Löffler, I. Panet, B. Pouilloux, P. Prieur, A. Rebray, S. Reynaud, B. Rievers, A. Robert, H. Selig, L. Serron, T. Sumner, N. Tanguy, and P. Visser, “Microscope mission: First results of a space test of the equivalence principle,” Phys. Rev. Lett. 119, 231101 (2017).
- Schlamminger et al. [2008] S. Schlamminger, K.-Y. Choi, T. A. Wagner, J. H. Gundlach, and E. G. Adelberger, “Test of the equivalence principle using a rotating torsion balance,” Phys. Rev. Lett. 100, 041101 (2008).
- Stadnik [2020] Y. V. Stadnik, “New bounds on macroscopic scalar-field topological defects from nontransient signatures due to environmental dependence and spatial variations of the fundamental constants,” Phys. Rev. D 102, 115016 (2020).
- Kozyryev, Lasner, and Doyle [2021] I. Kozyryev, Z. Lasner, and J. M. Doyle, “Enhanced sensitivity to ultralight bosonic dark matter in the spectra of the linear radical SrOH,” Phys. Rev. A 103, 043313 (2021).
- Madge, Perez, and Meir [2024] E. Madge, G. Perez, and Z. Meir, “Prospects of nuclear-coupled-dark-matter detection via correlation spectroscopy of I and Ca+,” Phys. Rev. D 110, 015008 (2024).
- Tsai, Eby, and Safronova [2023] Y.-D. Tsai, J. Eby, and M. S. Safronova, “Direct detection of ultralight dark matter bound to the sun with space quantum sensors,” Nat. Astron. 7, 113–121 (2023).
- Oswald et al. [2022] R. Oswald, A. Nevsky, V. Vogt, S. Schiller, N. L. Figueroa, K. Zhang, O. Tretiak, D. Antypas, D. Budker, A. Banerjee, et al., “Search for dark-matter-induced oscillations of fundamental constants using molecular spectroscopy,” Phys. Rev. Lett. 129, 031302 (2022).
- Zhang et al. [2023b] X. Zhang, A. Banerjee, M. Leyser, G. Perez, S. Schiller, D. Budker, and D. Antypas, “Search for ultralight dark matter with spectroscopy of radio-frequency atomic transitions,” Phys. Rev. Lett. 130, 251002 (2023b).
- Zhou et al. [2024] Y. Zhou, R. Ranson, M. Panagiotou, and C. Overstreet, “Ytterbium atom interferometry for dark matter searches,” Phys. Rev. A 110, 033313 (2024).
- Fuchs et al. [2024] E. Fuchs, F. Kirk, E. Madge, C. Paranjape, E. Peik, G. Perez, W. Ratzinger, and J. Tiedau, “Implications of the laser excitation of the Th-229 nucleus for dark matter searches,” (2024), arXiv:2407.15924 [hep-ph] .
- Antypas et al. [2022] D. Antypas, A. Banerjee, C. Bartram, M. Baryakhtar, J. Betz, J. J. Bollinger, C. Boutan, D. Bowring, D. Budker, D. Carney, et al., “New horizons: Scalar and vector ultralight dark matter,” (2022), arXiv:2203.14915 [hep-ex] .
- Berengut et al. [2018] J. C. Berengut, D. Budker, C. Delaunay, V. V. Flambaum, C. Frugiuele, E. Fuchs, C. Grojean, R. Harnik, R. Ozeri, G. Perez, et al., “Probing new long-range interactions by isotope shift spectroscopy,” Phys. Rev. Lett. 120, 091801 (2018).
- Mikami, Tanaka, and Yamamoto [2017] K. Mikami, M. Tanaka, and Y. Yamamoto, “Probing new intra-atomic force with isotope shifts,” The European Physical Journal C 77, 896 (2017).
- Delaunay et al. [2017] C. Delaunay, R. Ozeri, G. Perez, and Y. Soreq, “Probing atomic higgs-like forces at the precision frontier,” Phys. Rev. D 96, 093001 (2017).
- Frugiuele et al. [2017] C. Frugiuele, E. Fuchs, G. Perez, and M. Schlaffer, “Constraining new physics models with isotope shift spectroscopy,” Phys. Rev. D 96, 015011 (2017).
- Merton [1919] T. R. Merton, “Societies and academies,” Nature 104, 406–408 (1919).
- Breit [1958] G. Breit, “Theory of isotope shift,” Rev. Mod. Phys. 30, 507–516 (1958).
- King [1963] W. H. King, “Comments on the article “peculiarities of the isotope shift in the samarium spectrum”,” J. Opt. Soc. Am. 53, 638–639 (1963).
- King [1984] W. H. King, Isotope Shifts in Atomic Spectra (Springer US, New York, NY, 1984).
- Fricke and Heilig [2004] G. Fricke and K. Heilig, Nuclear Charge Radii (Springer, 2004).
- Angeli and Marinova [2013] I. Angeli and K. Marinova, “Table of experimental nuclear ground state charge radii: An update,” Atomic Data and Nuclear Data Tables 99, 69–95 (2013).
- Griffith et al. [1981] J. A. R. Griffith, G. R. Isaak, R. New, and M. P. Ralls, “Anomalies in the optical isotope shifts of samarium,” J. Phys. B: At. Mol. Phys. 14, 2769 (1981).
- Palmer and Stacey [1982] C. W. P. Palmer and D. N. Stacey, “Theory of anomalous isotope shifts in samarium,” J. Phys. B: At. Mol. Phys. 15, 997 (1982).
- Miyake et al. [2019] H. Miyake, N. C. Pisenti, P. K. Elgee, A. Sitaram, and G. K. Campbell, “Isotope-shift spectroscopy of the 1SP1 and 1SP0 transitions in strontium,” Phys. Rev. Res. 1, 033113 (2019).
- Berengut and Oreshkina [2024] J. C. Berengut and N. S. Oreshkina, “Second-order hyperfine structure in strontium and impact on new physics searches using isotope shift spectroscopy,” (2024), arXiv:2409.01530 [physics.atom-ph] .
- Berengut et al. [2020] J. C. Berengut, C. Delaunay, A. Geddes, and Y. Soreq, “Generalized King linearity and new physics searches with isotope shifts,” Phys. Rev. Res. 2, 043444 (2020).
- Yamamoto [2023] Y. Yamamoto, “The dual King relation,” Phys. Lett. B 838, 137682 (2023).
- Firak, D. S. et al. [2020] Firak, D. S., Krasznahorkay, A. J., Csatlós, M., Csige, L., Gulyás, J., Koszta, M., Szihalmi, B., Timár, J., Nagy, Á., Sas, N. J., et al., “Confirmation of the existence of the X17 particle,” EPJ Web Conf. 232, 04005 (2020).
- Berengut and Delaunay [2025] J. C. Berengut and C. Delaunay, “Precision isotope-shift spectroscopy for new physics searches and nuclear insights,” Nat. Rev. Phys. 7, 119–125 (2025).
- Solaro et al. [2020] C. Solaro, S. Meyer, K. Fisher, J. C. Berengut, E. Fuchs, and M. Drewsen, “Improved isotope-shift-based bounds on bosons beyond the standard model through measurements of the interval in Ca+,” Phys. Rev. Lett. 125, 123003 (2020).
- Chang et al. [2024] T. T. Chang, B. B. Awazi, J. C. Berengut, E. Fuchs, and S. C. Doret, “Systematic-free limit on new light scalar bosons via isotope-shift spectroscopy in Ca+,” Phys. Rev. A 110, L030801 (2024).
- Hur et al. [2022] J. Hur, D. P. L. Aude Craik, I. Counts, E. Knyazev, L. Caldwell, C. Leung, S. Pandey, J. C. Berengut, A. Geddes, W. Nazarewicz, et al., “Evidence of two-source King plot nonlinearity in spectroscopic search for new boson,” Phys. Rev. Lett. 128, 163201 (2022).
- Ono et al. [2022] K. Ono, Y. Saito, T. Ishiyama, T. Higomoto, T. Takano, Y. Takasu, Y. Yamamoto, M. Tanaka, and Y. Takahashi, “Observation of nonlinearity of generalized King plot in the search for new boson,” Phys. Rev. X 12, 021033 (2022).
- Counts et al. [2020] I. Counts, J. Hur, D. P. L. Aude Craik, H. Jeon, C. Leung, J. C. Berengut, A. Geddes, A. Kawasaki, W. Jhe, and V. Vuletić, “Evidence for nonlinear isotope shift in Yb+ search for new boson,” Phys. Rev. Lett. 125, 123002 (2020).
- Figueroa et al. [2022] N. L. Figueroa, J. C. Berengut, V. A. Dzuba, V. V. Flambaum, D. Budker, and D. Antypas, “Precision determination of isotope shifts in ytterbium and implications for new physics,” Phys. Rev. Lett. 128, 073001 (2022).
- Door et al. [2025] M. Door, C.-H. Yeh, M. Heinz, F. Kirk, C. Lyu, T. Miyagi, J. C. Berengut, J. Bieroń, K. Blaum, L. S. Dreissen, et al., “Probing new bosons and nuclear structure with ytterbium isotope shifts,” Phys. Rev. Lett. 134, 063002 (2025).
- Krasznahorkay et al. [2016] A. J. Krasznahorkay, M. Csatlós, L. Csige, Z. Gácsi, J. Gulyás, M. Hunyadi, I. Kuti, B. M. Nyakó, L. Stuhl, J. Timár, T. G. Tornyi, Z. Vajta, T. J. Ketel, and A. Krasznahorkay, “Observation of anomalous internal pair creation in : A possible indication of a light, neutral boson,” Phys. Rev. Lett. 116, 042501 (2016).
- Raffelt [2012] G. Raffelt, “Limits on a -violating scalar axion-nucleon interaction,” Phys. Rev. D 86, 015001 (2012).
- Grifols and Massó [1986] J. Grifols and E. Massó, “Constraints on finite-range baryonic and leptonic forces from stellar evolution,” Physics Letters B 173, 237–240 (1986).
- Fan et al. [2023] X. Fan, T. G. Myers, B. A. D. Sukra, and G. Gabrielse, “Measurement of the electron magnetic moment,” Phys. Rev. Lett. 130, 071801 (2023).
- Nesvizhevsky, Pignol, and Protasov [2008] V. V. Nesvizhevsky, G. Pignol, and K. V. Protasov, “Neutron scattering and extra-short-range interactions,” Phys. Rev. D 77, 034020 (2008).
- Harris, Chen, and Mohideen [2000] B. W. Harris, F. Chen, and U. Mohideen, “Precision measurement of the casimir force using gold surfaces,” Phys. Rev. A 62, 052109 (2000).
- Potvliege et al. [2023] R. M. Potvliege, A. Nicolson, M. P. A. Jones, and M. Spannowsky, “Deuterium spectroscopy for enhanced bounds on physics beyond the standard model,” Phys. Rev. A 108, 052825 (2023).
- Ohayon et al. [2022] B. Ohayon, S. Hofsäss, J. E. Padilla-Castillo, S. C. Wright, G. Meijer, S. Truppe, K. Gibble, and B. K. Sahoo, “Isotope shifts in cadmium as a sensitive probe for physics beyond the standard model,” New J. Phys. 24, 123040 (2022).
- Bhatt, Kato, and Vutha [2020] N. Bhatt, K. Kato, and A. C. Vutha, “Nd+ isotope shift measurements in a cryogenically cooled neutral plasma,” Phys. Rev. A 101, 052505 (2020).
- Huang et al. [2021] W. Huang, M. Wang, F. Kondev, G. Audi, and S. Naimi, “The AME 2020 atomic mass evaluation (I). Evaluation of input data, and adjustment procedures,” Chin. Phys. C 45, 030002 (2021).
- Wang et al. [2021] M. Wang, W. Huang, F. Kondev, G. Audi, and S. Naimi, “The AME 2020 atomic mass evaluation (II). tables, graphs and references*,” Chin. Phys. C 45, 030003 (2021).
- Nesterenko et al. [2020] D. Nesterenko, R. de Groote, T. Eronen, Z. Ge, M. Hukkanen, A. Jokinen, and A. Kankainen, “High-precision mass measurement of 168Yb for verification of nonlinear isotope shift,” Int. J. Mass Spectrom. 458, 116435 (2020).
- Lee and Yang [1956] T. D. Lee and C. N. Yang, “Question of parity conservation in weak interactions,” Phys. Rev. 104, 254–258 (1956).
- Wu et al. [1957] C. S. Wu, E. Ambler, R. W. Hayward, D. D. Hoppes, and R. P. Hudson, “Experimental test of parity conservation in beta decay,” Phys. Rev. 105, 1413–1415 (1957).
- Bouchiat and Piketty [1983] C. Bouchiat and C. Piketty, “Parity violation in atomic cesium and alternatives to the standard model of electroweak interactions,” Phys. Lett. B 128, 73–78 (1983).
- Arcadi et al. [2020] G. Arcadi, M. Lindner, J. Martins, and F. S. Queiroz, “New physics probes: Atomic parity violation, polarized electron scattering and neutrino-nucleus coherent scattering,” Nucl. Phys. B 959, 115158 (2020).
- Marciano and Rosner [1990] W. J. Marciano and J. L. Rosner, “Atomic parity violation as a probe of new physics,” Phys. Rev. Lett. 65, 2963–2966 (1990).
- Dzuba, Flambaum, and Stadnik [2017] V. A. Dzuba, V. V. Flambaum, and Y. V. Stadnik, “Probing low-mass vector bosons with parity nonconservation and nuclear anapole moment measurements in atoms and molecules,” Phys. Rev. Lett. 119, 223201 (2017).
- Bouchiat and Bouchiat [1974] M. A. Bouchiat and C. Bouchiat, “I. parity violation induced by weak neutral currents in atomic physics,” J. Phys. France 35, 899 (1974).
- Bouchiat and Bouchiat [1975] M. A. Bouchiat and C. Bouchiat, “Parity violation induced by weak neutral currents in atomic physics. part II,” J. Phys. France 36, 493 (1975).
- Gilbert and Wieman [1986] S. L. Gilbert and C. E. Wieman, “Atomic-beam measurement of parity nonconservation in cesium,” Phys. Rev. A 34, 792–803 (1986).
- Wood et al. [1997] C. S. Wood, S. C. Bennett, D. Cho, B. P. Masterson, J. L. Roberts, C. E. Tanner, and C. E. Wieman, “Measurement of parity nonconservation and an anapole moment in cesium,” Science 275, 1759–1763 (1997).
- Tsigutkin et al. [2009] K. Tsigutkin, D. Dounas-Frazer, A. Family, J. E. Stalnaker, V. V. Yashchuk, and D. Budker, “Observation of a large atomic parity violation effect in ytterbium,” Phys. Rev. Lett. 103, 071601 (2009).
- Antypas et al. [2019] D. Antypas, A. Fabricant, J. E. Stalnaker, K. Tsigutkin, V. V. Flambaum, and D. Budker, “Isotopic variation of parity violation in atomic ytterbium,” Nat. Phys. 15, 120–123 (2019).
- Simsarian et al. [1996] J. E. Simsarian, A. Ghosh, G. Gwinner, L. A. Orozco, G. D. Sprouse, and P. A. Voytas, “Magneto-optic trapping of 210Fr,” Phys. Rev. Lett. 76, 3522–3525 (1996).
- Collister et al. [2014] R. Collister, G. Gwinner, M. Tandecki, J. A. Behr, M. R. Pearson, J. Zhang, L. A. Orozco, S. Aubin, and E. Gomez (FrPNC Collaboration), “Isotope shifts in francium isotopes and ,” Phys. Rev. A 90, 052502 (2014).
- Kalita et al. [2018] M. R. Kalita, J. A. Behr, A. Gorelov, M. R. Pearson, A. C. DeHart, G. Gwinner, M. J. Kossin, L. A. Orozco, S. Aubin, E. Gomez, et al., “Isotope shifts in the transition of francium: Measurements and comparison to ab initio theory,” Phys. Rev. A 97, 042507 (2018).
- Quirk et al. [2024] J. A. Quirk, A. Jacobsen, A. Damitz, C. E. Tanner, and D. S. Elliott, “Measurement of the static stark shift of the level in atomic cesium,” Phys. Rev. Lett. 132, 233201 (2024).
- Flambaum and Khriplovich [1985] V. Flambaum and I. Khriplovich, “On the enhancement of parity nonconserving effects in diatomic molecules,” Phys. Lett. A 110, 121–125 (1985).
- Leefer et al. [2014] N. Leefer, L. Bougas, D. Antypas, and D. Budker, “Towards a new measurement of parity violation in dysprosium,” (2014), arXiv:1412.1245 [physics.atom-ph] .
- Altuntaş et al. [2018] E. Altuntaş, J. Ammon, S. B. Cahn, and D. DeMille, “Demonstration of a sensitive method to measure nuclear-spin-dependent parity violation,” Phys. Rev. Lett. 120, 142501 (2018).
- Zeng et al. [2024] Z. Zeng, S. Deng, S. Yang, and B. Yan, “Three-dimensional magneto-optical trapping of barium monofluoride,” Phys. Rev. Lett. 133, 143404 (2024).
- Bouchiat et al. [1982] M. Bouchiat, J. Guena, L. Hunter, and L. Pottier, “Observation of a parity violation in cesium,” Phys. Lett. B 117, 358–364 (1982).
- Bouchiat, Guena, and Pottier [1986] M. Bouchiat, J. Guena, and L. Pottier, “Atomic parity violation measurements in the highly forbidden 6S1/2-7S1/2 caesium transition. II. analysis and control of systematic effects,” J. Phys. France 47, 1175 (1986).
- Noecker, Masterson, and Wieman [1988] M. C. Noecker, B. P. Masterson, and C. E. Wieman, “Precision measurement of parity nonconservation in atomic cesium: A low-energy test of the electroweak theory,” Phys. Rev. Lett. 61, 310–313 (1988).
- Luders [1954] G. Luders, “On the equivalence of invariance under time reersal and under particle-antiparticle conjugation for relativistic field theories,” Kgl. Danske Videnskab. Selskab. Mat.-Fys. Medd. 28, 1–17 (1954).
- Lüders [1957] G. Lüders, “Proof of the TCP theorem,” Annals of Physics 2, 1–15 (1957).
- Pospelov and Ritz [2014] M. Pospelov and A. Ritz, “CKM benchmarks for electron electric dipole moment experiments,” Phys. Rev. D 89, 056006 (2014).
- Ema, Gao, and Pospelov [2022] Y. Ema, T. Gao, and M. Pospelov, “Standard model prediction for paramagnetic electric dipole moments,” Phys. Rev. Lett. 129, 231801 (2022).
- Dzuba et al. [2018] V. A. Dzuba, V. V. Flambaum, I. B. Samsonov, and Y. V. Stadnik, “New constraints on axion-mediated ,-violating interaction from electric dipole moments of diamagnetic atoms,” Phys. Rev. D 98, 035048 (2018).
- Stadnik, Dzuba, and Flambaum [2018] Y. V. Stadnik, V. A. Dzuba, and V. V. Flambaum, “Improved limits on axionlike-particle-mediated , -violating interactions between electrons and nucleons from electric dipole moments of atoms and molecules,” Phys. Rev. Lett. 120, 013202 (2018).
- Purcell and Ramsey [1950] E. M. Purcell and N. F. Ramsey, “On the possibility of electric dipole moments for elementary particles and nuclei,” Phys. Rev. 78, 807–807 (1950).
- Chupp et al. [2019] T. E. Chupp, P. Fierlinger, M. J. Ramsey-Musolf, and J. T. Singh, “Electric dipole moments of atoms, molecules, nuclei, and particles,” Rev. Mod. Phys. 91, 015001 (2019).
- Schiff [1963] L. I. Schiff, “Measurability of nuclear electric dipole moments,” Phys. Rev. 132, 2194–2200 (1963).
- Sandars [1965] P. Sandars, “The electric dipole moment of an atom,” Physics Letters 14, 194–196 (1965).
- Cesarotti et al. [2019] C. Cesarotti, Q. Lu, Y. Nakai, A. Parikh, and M. Reece, “Interpreting the electron EDM constraint,” J. High Energy Phys. 2019, 59 (2019).
- Feinberg [1958] G. Feinberg, “Effects of an electric dipole moment of the electron on the hydrogen energy levels,” Phys. Rev. 112, 1637–1642 (1958).
- Nelson et al. [1959] D. F. Nelson, A. A. Schupp, R. W. Pidd, and H. R. Crane, “Search for an electric dipole moment of the electron,” Phys. Rev. Lett. 2, 492–495 (1959).
- Salpeter [1958] E. E. Salpeter, “Some atomic effects of an electronic electric dipole moment,” Phys. Rev. 112, 1642–1648 (1958).
- Burleson and Kendall [1960] G. R. Burleson and H. Kendall, “An experimental search for dipole structure of the electron,” Nuclear Physics 19, 68–78 (1960).
- Goldemberg and Torizuka [1963] J. Goldemberg and Y. Torizuka, “Upper limit of the electric dipole moment of the electron,” Phys. Rev. 129, 2580–2581 (1963).
- Wilkinson and Crane [1963] D. T. Wilkinson and H. R. Crane, “Precision measurement of the factor of the free electron,” Phys. Rev. 130, 852–863 (1963).
- Sandars and Lipworth [1964] P. G. H. Sandars and E. Lipworth, “Electric dipole moment of the cesium atom. a new upper limit to the electric dipole moment of the free electron,” Phys. Rev. Lett. 13, 718–720 (1964).
- Stein et al. [1967] T. S. Stein, J. P. Carrico, E. Lipworth, and M. C. Weisskopf, “Electric dipole moment of the cesium atom. a new upper limit to the electric dipole moment of the free electron,” Phys. Rev. Lett. 19, 741–743 (1967).
- Angel, Sandars, and Tinker [1967] J. Angel, P. Sandars, and M. Tinker, “Observation of a v × e effect in an electric dipole moment experiment using a reversible atomic beam machine,” Phys. Lett. A 25, 160–161 (1967).
- Weisskopf et al. [1968] M. C. Weisskopf, J. P. Carrico, H. Gould, E. Lipworth, and T. S. Stein, “Electric dipole moment of the cesium atom. a new upper limit to the electric dipole moment of the electron,” Phys. Rev. Lett. 21, 1645–1648 (1968).
- Carrico et al. [1968] J. P. Carrico, E. Lipworth, P. G. H. Sandars, T. S. Stein, and M. C. Weisskopf, “Electric dipole moments of alkali atoms. a limit to the electric dipole moment of the free electron,” Phys. Rev. 174, 125–138 (1968).
- Player and Sandars [1970] M. A. Player and P. G. H. Sandars, “An experiment to search for an electric dipole moment in the 3P2 metastable state of xenon,” J. Phys. B: At. Mol. Phys. 3, 1620 (1970).
- Gould [1970] H. Gould, “Search for an electric dipole moment in thallium,” Phys. Rev. Lett. 24, 1091–1093 (1970).
- Lamoreaux et al. [1987] S. K. Lamoreaux, J. P. Jacobs, B. R. Heckel, F. J. Raab, and N. Fortson, “New constraints on time-reversal asymmetry from a search for a permanent electric dipole moment of ,” Phys. Rev. Lett. 59, 2275–2278 (1987).
- Murthy et al. [1989] S. A. Murthy, D. Krause, Z. L. Li, and L. R. Hunter, “New limits on the electron electric dipole moment from cesium,” Phys. Rev. Lett. 63, 965–968 (1989).
- Abdullah et al. [1990] K. Abdullah, C. Carlberg, E. D. Commins, H. Gould, and S. B. Ross, “New experimental limit on the electron electric dipole moment,” Phys. Rev. Lett. 65, 2347–2350 (1990).
- Commins et al. [1994] E. D. Commins, S. B. Ross, D. DeMille, and B. C. Regan, “Improved experimental limit on the electric dipole moment of the electron,” Phys. Rev. A 50, 2960–2977 (1994).
- Regan et al. [2002] B. C. Regan, E. D. Commins, C. J. Schmidt, and D. DeMille, “New limit on the electron electric dipole moment,” Phys. Rev. Lett. 88, 071805 (2002).
- Cho, Sangster, and Hinds [1989] D. Cho, K. Sangster, and E. A. Hinds, “Tenfold improvement of limits on T violation in thallium fluoride,” Phys. Rev. Lett. 63, 2559–2562 (1989).
- Hudson et al. [2002] J. J. Hudson, B. E. Sauer, M. R. Tarbutt, and E. A. Hinds, “Measurement of the electron electric dipole moment using YbF molecules,” Phys. Rev. Lett. 89, 023003 (2002).
- Hudson et al. [2011] J. J. Hudson, D. M. Kara, I. J. Smallman, B. E. Sauer, M. R. Tarbutt, and E. A. Hinds, “Improved measurement of the shape of the electron,” Nature 473, 493–496 (2011).
- Eckel et al. [2013] S. Eckel, P. Hamilton, E. Kirilov, H. W. Smith, and D. DeMille, “Search for the electron electric dipole moment using -doublet levels in PbO,” Phys. Rev. A 87, 052130 (2013).
- Baron et al. [2014] J. Baron, W. C. Campbell, D. DeMille, J. M. Doyle, G. Gabrielse, Y. V. Gurevich, P. W. Hess, N. R. Hutzler, E. Kirilov, I. Kozyryev, et al. (The ACME Collaboration), “Order of magnitude smaller limit on the electric dipole moment of the electron,” Science 343, 269–272 (2014).
- Cairncross et al. [2017] W. B. Cairncross, D. N. Gresh, M. Grau, K. C. Cossel, T. S. Roussy, Y. Ni, Y. Zhou, J. Ye, and E. A. Cornell, “Precision measurement of the electron’s electric dipole moment using trapped molecular ions,” Phys. Rev. Lett. 119, 153001 (2017).
- Andreev et al. [2018] V. Andreev, D. G. Ang, D. DeMille, J. M. Doyle, G. Gabrielse, J. Haefner, N. R. Hutzler, Z. Lasner, C. Meisenhelder, B. R. O’Leary, et al., “Improved limit on the electric dipole moment of the electron,” Nature 562, 355–360 (2018).
- Roussy et al. [2023] T. S. Roussy, L. Caldwell, T. Wright, W. B. Cairncross, Y. Shagam, K. B. Ng, N. Schlossberger, S. Y. Park, A. Wang, J. Ye, et al., “An improved bound on the electron’s electric dipole moment,” Science 381, 46–50 (2023).
- Commins, Jackson, and DeMille [2007] E. D. Commins, J. D. Jackson, and D. P. DeMille, “The electric dipole moment of the electron: An intuitive explanation for the evasion of Schiff’s theorem,” Am. J. Phys. 75, 532–536 (2007).
- Skripnikov [2016] L. V. Skripnikov, “Combined 4-component and relativistic pseudopotential study of tho for the electron electric dipole moment search,” The Journal of Chemical Physics 145, 214301 (2016), https://pubs.aip.org/aip/jcp/article-pdf/doi/10.1063/1.4968229/15523128/214301_1_online.pdf .
- Abe et al. [2014] M. Abe, G. Gopakumar, M. Hada, B. P. Das, H. Tatewaki, and D. Mukherjee, “Application of relativistic coupled-cluster theory to the effective electric field in ybf,” Phys. Rev. A 90, 022501 (2014).
- Fleig [2017] T. Fleig, “-odd and magnetic hyperfine-interaction constants and excited-state lifetime for ,” Phys. Rev. A 96, 040502 (2017).
- Kawall et al. [2004] D. Kawall, F. Bay, S. Bickman, Y. Jiang, and D. DeMille, “Precision zeeman-stark spectroscopy of the metastable state of PbO,” Phys. Rev. Lett. 92, 133007 (2004).
- Baron et al. [2017] J. Baron, W. C. Campbell, D. DeMille, J. M. Doyle, G. Gabrielse, Y. V. Gurevich, P. W. Hess, N. R. Hutzler, E. Kirilov, I. Kozyryev, B. R. O’Leary, C. D. Panda, M. F. Parsons, B. Spaun, A. C. Vutha, A. D. West, E. P. West, and A. Collaboration, “Methods, analysis, and the treatment of systematic errors for the electron electric dipole moment search in thorium monoxide,” New Journal of Physics 19, 073029 (2017).
- Ang et al. [2022] D. G. Ang, C. Meisenhelder, C. D. Panda, X. Wu, D. DeMille, J. M. Doyle, and G. Gabrielse, “Measurement of the radiative lifetime in ThO,” Phys. Rev. A 106, 022808 (2022).
- Hutzler et al. [2011] N. R. Hutzler, M. F. Parsons, Y. V. Gurevich, P. W. Hess, E. Petrik, B. Spaun, A. C. Vutha, D. DeMille, G. Gabrielse, and J. M. Doyle, “A cryogenic beam of refractory, chemically reactive molecules with expansion cooling,” Phys. Chem. Chem. Phys. 13, 18976–18985 (2011).
- Wu et al. [2022] X. Wu, P. Hu, Z. Han, D. G. Ang, C. Meisenhelder, G. Gabrielse, J. M. Doyle, and D. DeMille, “Electrostatic focusing of cold and heavy molecules for the ACME electron EDM search,” New J. Phys. 24, 073043 (2022).
- Ng et al. [2022] K. B. Ng, Y. Zhou, L. Cheng, N. Schlossberger, S. Y. Park, T. S. Roussy, L. Caldwell, Y. Shagam, A. J. Vigil, E. A. Cornell, et al., “Spectroscopy on the electron-electric-dipole-moment–sensitive states of ThF+,” Phys. Rev. A 105, 022823 (2022).
- Fitch et al. [2020] N. J. Fitch, J. Lim, E. A. Hinds, B. E. Sauer, and M. R. Tarbutt, “Methods for measuring the electron’s electric dipole moment using ultracold YbF molecules,” Quantum Sci. Technol. 6, 014006 (2020).
- Aoki et al. [2021] T. Aoki, R. Sreekantham, B. K. Sahoo, B. Arora, A. Kastberg, T. Sato, H. Ikeda, N. Okamoto, Y. Torii, T. Hayamizu, et al., “Quantum sensing of the electron electric dipole moment using ultracold entangled Fr atoms,” Quantum Sci. Technol. 6, 044008 (2021).
- Kudashov et al. [2014] A. D. Kudashov, A. N. Petrov, L. V. Skripnikov, N. S. Mosyagin, T. A. Isaev, R. Berger, and A. V. Titov, “Ab initio study of radium monofluoride (RaF) as a candidate to search for parity- and time-and-parity–violation effects,” Phys. Rev. A 90, 052513 (2014).
- Garcia Ruiz et al. [2020] R. F. Garcia Ruiz, R. Berger, J. Billowes, C. L. Binnersley, M. L. Bissell, A. A. Breier, A. J. Brinson, K. Chrysalidis, T. E. Cocolios, B. S. Cooper, et al., “Spectroscopy of short-lived radioactive molecules,” Nature 581, 396–400 (2020).
- Yu and Hutzler [2021] P. Yu and N. R. Hutzler, “Probing fundamental symmetries of deformed nuclei in symmetric top molecules,” Phys. Rev. Lett. 126, 023003 (2021).
- Fleig and DeMille [2021] T. Fleig and D. DeMille, “Theoretical aspects of radium-containing molecules amenable to assembly from laser-cooled atoms for new physics searches,” New J. Phys. 23, 113039 (2021).
- Fan et al. [2021] M. Fan, C. A. Holliman, X. Shi, H. Zhang, M. W. Straus, X. Li, S. W. Buechele, and A. M. Jayich, “Optical mass spectrometry of cold RaOH+ and ,” Phys. Rev. Lett. 126, 023002 (2021).
- Kozyryev and Hutzler [2017] I. Kozyryev and N. R. Hutzler, “Precision measurement of time-reversal symmetry violation with laser-cooled polyatomic molecules,” Phys. Rev. Lett. 119, 133002 (2017).
- Denis et al. [2019] M. Denis, P. A. B. Haase, R. G. E. Timmermans, E. Eliav, N. R. Hutzler, and A. Borschevsky, “Enhancement factor for the electric dipole moment of the electron in the BaOH and YbOH molecules,” Phys. Rev. A 99, 042512 (2019).
- Augenbraun et al. [2021] B. L. Augenbraun, Z. D. Lasner, A. Frenett, H. Sawaoka, A. T. Le, J. M. Doyle, and T. C. Steimle, “Observation and laser spectroscopy of ytterbium monomethoxide, YbOCH3,” Phys. Rev. A 103, 022814 (2021).
- Lasner et al. [2024] Z. D. Lasner, A. Frenett, H. Sawaoka, L. Anderegg, B. Augenbraun, H. Lampson, M. Li, A. Lunstad, J. Mango, A. Nasir, et al., “Magneto-optical trapping of a heavy polyatomic molecule for precision measurement,” (2024), arXiv:2409.04948 [physics.atom-ph] .
- Zhang et al. [2023c] C. Zhang, P. Yu, A. Jadbabaie, and N. R. Hutzler, “Quantum-enhanced metrology for molecular symmetry violation using decoherence-free subspaces,” Phys. Rev. Lett. 131, 193602 (2023c).
- Graner et al. [2016] B. Graner, Y. Chen, E. G. Lindahl, and B. R. Heckel, “Reduced limit on the permanent electric dipole moment of ,” Phys. Rev. Lett. 116, 161601 (2016).
- Zheng et al. [2022b] T. A. Zheng, Y. A. Yang, S.-Z. Wang, J. T. Singh, Z.-X. Xiong, T. Xia, and Z.-T. Lu, “Measurement of the electric dipole moment of atoms in an optical dipole trap,” Phys. Rev. Lett. 129, 083001 (2022b).
- Bishof et al. [2016] M. Bishof, R. H. Parker, K. G. Bailey, J. P. Greene, R. J. Holt, M. R. Kalita, W. Korsch, N. D. Lemke, Z.-T. Lu, P. Mueller, et al., “Improved limit on the electric dipole moment,” Phys. Rev. C 94, 025501 (2016).